Lie groups, Liegruppen

Vorlesungsmanuskripte von Eberhard Freitag, Heidelberg (allgemeiner Teil, Kapitel I-III) Reinhardt Kiehl, Mannheim (Strukturtheorie der Liealgebren)

Contents

Chapter I. Elementary properties of Lie groups 1

1. Calculus 1 2. Analytic manifolds 3 3. The tangent space 7 4. Sub-manifolds 12 5. Vector fields 14

Chapter II. Basic properties of Lie groups and Lie algebras 17

1. Basic properties of Lie groups 17 2. The Lie algebra 19 3. The exponential map 21 4. The Taylor expansion of the exponential map 26

Chapter III. Relation between Lie groups and Lie algebras 30

1. Simple topological facts about Lie groups 30 2. Closed subgroups of Lie groups 30 3. Continuous homomorphisms of Lie group 32 4. The theorem of Poincaré Birkhoff Witt 33 5. Lie sub algebras and local subgroups 40 6. Local subgroups 42 7. Abstract Lie algebras as Lie algebras of Lie groups 44

Chapter IV. Lie-Algebren 47

1. The Campbell-Hausdorff series 47 2. Das Casimir-Element 55 3. Auflösbare und nilpotente Liealgebren 58 5. Halbeinfache Liealgebren 70 6. Der Satz von Ado 75

Contents II

Chapter V. Klassifikation halbeinfacher Liealgebren 79

1. Wurzeln und Wurzelräume 79 2. Cartanalgebren 87 3. Darstellungen der Liealgebra sl_{2} 91 4. Das Wurzelsystem einer halbeinfachen Liealgebra 92 5. Der Isomorphiesatz 98 6. Anhang 102

Chapter I. Elementary properties of Lie groups

1. Calculus

When we talk about differentiable functions we always mean functions which admit continuous derivatives of arbitrary order. We denote by

C^{∞}(U, V) = { f : U ⟶ V; f differentiable }

the set of all differentiable functions from an open subset U ⊂ ℝ^{n} to an open subset V ⊂ ℝ^{m}. In the case V = ℝ one obtains the ring of differentiable functions

C^{∞}(U) := C^{∞}(U, ℝ).

The operators

∂_{ν} := \frac{∂}{∂x_{ν}} : C^{∞}(U) ⟶ C^{∞}(U)

commute. For any multiindex

ν = (ν_{1}, …, ν_{n}) ∈ ℕ_{0}^{n} (ℕ_{0} = { 0, 1, 2, … })

we define

∂_{ν} = ∂_{ν_{1}} ∘ … ∘ ∂_{ν_{n}}.

The Jacobian matrix of a map f ∈ C^{∞}(U, V) is

J(f) = ( \begin{matrix}∂_{1}f_{1} & … & ∂_{n}f_{1} \\ ⋮ & & ⋮ \\ ∂_{1}f_{m} & … & ∂_{n}f_{m} \end{matrix} ) .

We denote by J(f, a) the evaluation of J(f) at a point a ∈ U.

1.1 The chain rule. Let U ⊂ ℝ^{n}, V ⊂ ℝ^{m}, W ⊂ ℝ^{p} be open subsets and f : U → V, g : V → W be differentiable functions. Then g ∘ f is differentiable and we have

J(g ∘ f, a) = J(g, f(a))J(f, a).

By definition a diffeomorphism f : U → V between open subsets U ⊂ ℝ^{n} and V ⊂ ℝ^{m} is a bijective map such that f and f^{−1} both are differentiable. From the chain rule follows n = m (purity of the dimension).

In general we denote a map f : U → V, U, V ⊂ ℝ^{n}, a local diffeomorphism at a ∈ U if there exists an open neighbourhood a ∈ U_{0} ⊂ U such that V_{0} = f(U_{0}) is open too and such that f induces a diffeomorphism U_{0} → V_{0}.

1.2 The theorem of invertible functions. A map f : U → V is a local diffeomorphism at a ∈ U, if and only if the Jacobian matrix J(f, a) is invertible.

The theorem of local immersions is a generalization of 1.2 (actually a consequence). A differentiable map

f : U ⟶ V, U ⊂ ℝ^{n}, V ⊂ ℝ^{n} (both open)

is called a local immersion at a ⊂ U if the Jacobian matrix has the rank n (which implies n ≤ m). In the case n = m a local immersion is a local diffeomorphism (1.2).

1.3 The theorem of local immersions. A differentiable map f : U → V is a local immersion at a ∈ U if and only if there exist open neighbourhoods

a ∈ U_{0} ⊂ U, f(a) ∈ V_{0} ⊂ V

with the following properties:

1. f(U_{0}) ⊂ V_{0}.

2. There exists a local diffeomorphism φ : U_{0} → ℝ^{n} at a and a local diffeomorphism ψ : V_{0} → ℝ^{m} at f(a) such that the diagram

U_{0} ⟶^{f} V_{0}
φ ↓  ↓ ψ
ℝ^{n} ⟶^{ι} ℝ^{m}

commutes. Here ι denotes the natural embedding

i(x_{1}, …, x_{n}) = (x_{1}, …, x_{n}, 0, …, 0).

We recall the notion of a real analytic function: A function f ∈ C^{∞}(U) is called real analytic, if every point a ∈ U admits an open neighbourhood a ∈ U_{0} ⊂ U such there exists an absolutely convergent expansion as power series

f(x) = ∑_{ν∈ℕ_{0}^{n}} a_{ν}(x − a)^{ν} (x ∈ U_{0}).

Here we use the notation

(x − a)^{ν} := (x_{1} − a_{1})^{ν_{1}} ⋯ (x_{n} − a_{n})^{ν_{n}}.

The coefficients a_{ν} are unique and can be determined by the well-known Taylor formula

a_{ν} = \frac{∂^{ν}f(a)}{ν!} (ν! = ν_{1} ⋯ ν_{n}).

§2. Analytic manifolds 3

A fundamental property of analytic functions is 1.4 The principle of analytic continuation. Let U ⊂ ℝ^{n} be a domain, i.e. an open and connected subset. Let f, g : U → ℝ be two analytic functions and a ∈ U a point such that all derivatives of f and g in a agree,

(∂_{ν}f)(a) = (∂_{ν}g)(a) for all ν ∈ ℕ_{0}^{n}.

Then f = g. Special case: f and g agree in a small neighbourhood of a.

We recall that a topological X space is called connected if any locally function on X is constant and arcwise connected if for any two points a, b ∈ X there exists a continous map of a real intervall into X (an arc) whose image contains a and b. Usually arcwise connectedness is a stronger condition than connectedness. But in all situations which will occur to us both notions agree, for example for open subsets of ℝ^{n} or more generally for topological manifolds (topological spaces which locally are homeomorphic to open subsets of ℝ^{n}).

We denote by C^{ω}(U) ⊂ C^{∞}(U) the set of all analytic functions and more generally by

C^{ω}(U, V) ⊂ C^{∞}(U, V) (U ⊂ ℝ^{n}, V ⊂ ℝ^{m} both open).

the set of analytic mappings (which means that all m components are analytic functions). The following permanence properties are well-known:

1. Sum and product of analytic functions are analytic, constant functions are analytic.

2. Let f be an analytic function without zeros, then 1/f is analytic too.

3. The composition of two analytic maps is analytic.

4. Let f be a diffeomorphism such that f is analytic. Then f^{−1} is analytic too.

2. Analytic manifolds

It is not possible to introduce the notion of a differentiable or analytic function on a general topological space. This requires an additional structure, which we are going to introduce here as a geometric structure. In the following we denote by

C(X) := { f : X ⟶ ℝ; f continous }

the ring of continuous functions on X.

2.1 Definition. A geometric structure O on a topological space X is a collection

O = (O(U))_{U⊂X open}

of subsets

O(U) ⊂ C(U) (U ⊂ X open)

of so-called distinguished functions such that the following properties are satisfied.

1. Sums and products of distinguished functions are distinguished, constant functions are distinguished.

2. Restrictions of distinguished functions are distinguished*).

3. To be distinguished is a local property**).

We describe two very trivial processes for geometric structures:

1) Let U ⊂ X be an open subset and O a geometric structure on X. The restriction O|U is the geometric structure on U defined by

(O|U)(V) := O(V).

Of course U is equipped with the induced topology. A subset V ⊂ U is open in U if and only it is open in X. (This is only true because we assumed that U is open.)

2) Let O be a geometric structure on X and f : X → Y a topological mapping, i.e. a bijective map which is continuous in both directions. Then one can define the transported geometric structure O′ on Y in a trivial manner: A function g : V → ℝ (V ⊂ Y open) is called distinguished if the pull-back g ∘f : f^{−1}(V) → ℝ is distinguished (with respect to O).

Examples of geometric structures on ℝ^{n} are given by

O(U) = C(U), C^{∞}(U), C^{ω}(U).

We denote these structures by O = C_{ℝ}^{n}, C_{ℝ^{n}}^{∞}, C_{ℝ^{n}}^{ω}. Analytic manifolds are modelled by the geometric structure C_{ℝ^{n}}^{ω}:

2.2 Definition. An analytic manifold (X, O) is pair consisting of a topological space and a geometric structure O such that the following is true:

For every point a ∈ X there exists an open neighbourhood a ∈ U ⊂ X and a topological map φ : U → V of U onto some open subset V ⊂ ℝ^{n} for some suitable n ≥ 0 such that the transport of O|U is C_{ℝ^{n}}^{ω}|V.

*) Let V ⊂ U ⊂ X open subsets and f ∈ O(U). Then f|V ∈ O(V).

**) Let f : U → ℝ be a function on an open subset of X such that for every point a ∈ U there exists an open neighbourhood U(a) ⊂ U such that f|U(a) is distinguished. Then f is distinguished.

Trivial remarks are:

1) (ℝ^{n}, C_{ℝ^{n}}^{ω}) is an analytic manifold.

2) If U is an open subset of an analytic manifold (X, O) then (U, O|U) is an analytic manifold too.

A topological space can carry different analytic structures. In the following it will be usually clear from the context which analytic structure is considered. Hence we write usually X instead of (X, O). We also will use the notation C_{X}^{ω} instead of O.

Standard notation for analytic manifolds: X = (X, C_{X}^{ω}).

2.3 Definition. A map f : X → Y between analytic manifolds id called analytic if it is continuous and if for any open subset V ⊂ V and any g ∈ C_{Y}^{ω} (V) one has g ∘ f ∈ C_{X}^{ω}(f^{−1}(V)).

We mention some trivialities:

1) The identity map id_{X} of an analytic manifold is analytic.

2) The composition of two analytic maps are analytic.

2.4 Remark. Let U ⊂ ℝ^{n} and V ⊂ ℝ^{n} be open subsets. For a map f : U → V the following to conditions are equivalent:

1. f is analytic in the usual sense (s. sec. 1).

2. The map f : (U, C_{ℝ^{n}}^{ω}|U) → (V, C_{ℝ^{m}}^{ω}|V) is an analytic map of analytic manifolds.

The proof is trivial.

2.5 Remark. Let f : U → ℝ be a function on an open subset of an analytic manifold (X, C_{X}^{ω}). The following two conditions are equivalent:

1. f ∈ C_{X}^{ω}(U).

2. The map f : (X, C_{X}^{ω}) → (ℝ, C_{ℝ}^{ω}) is an analytic map of analytic manifolds.

We skip the simple proof and mention just that 2.5 justifies to call the elements of C_{X}^{ω}(U) the analytic functions on U.

2.6 Definition. A map f : X → Y between analytic manifolds is called an analytic diffeomorphism if f is bijective and if both f and f^{−1} are analytic.

A different way to express this is:

The map f is topological and C_{Y}^{ω} is the transported structure C_{X}^{ω}.

Charts

Charts can be considered as local coordinates. A chart φ on a topological space X is a tripel (U, φ, V) consisting of an open subset U ⊂ X (the domain of definition of φ), an open subset V ⊂ ℝ^{n} (the chart-sheet) and a topological map φ : U → V. This means that every point a ∈ U can be described by a system of numbers φ(a) ∈ ℝ^{n} which should be considered as “coordinates” of the point a.

2.7 Definition. A chart φ : U → V on an analytic manifold X is called an analytic chart if

φ : (U, C_{X}^{ω}|U) ⟶ (V, C_{ℝ^{n}}^{ω}|V)

is an analytic diffeomorphism.

Charts are important because the allow to transform data on an abstract manifold into a description by coordinates. Assume that there are two charts

φ : U → V, φ′ : U′ → V′.

given on X. We do not assume that U equals U′ but what follows is only of relevance if U and U′ have non-empty intersection. We notice: The set U ∩ U′ is mapped topologically to φ(U ∩ U′) ⊂ V by φ and to φ′(U ∩ U′) ⊂ V′ by φ′. Inverting φ we obtain a toplogical map

φ′ ∘ φ^{−1} : φ(U ∩ U′) → φ′(U ∩ U′).

(The notation φ′ ∘ φ^{−1} does not follow the strong rules of set theory. We will use it nevertheless as long as it is clear what is meant.) We call this map φ′ ∘ φ^{−1} the coordinate transformation from φ to φ′. We notice:

1) The coordinate transformation is a topological map.

2) The sets φ(U ∩ U′) ⊂ V and φ′(U ∩ U′) ⊂ V′ are open (hence open in some “coordinate space” ℝ^{n}, ℝ^{n′}).

Finally we mention the notion of an atlas. An atlas A on a topological space is a set of charts whose domains of definition cover X,

X = ⋃_{φ:U_{φ}→V_{φ}} U_{φ}

and a topological manifold is just a topological space such that the set of all charts performs an atlas.

2.8 Remark. The set of all analytic charts on an analytic manifold performs an atlas. The coordinate transformation between two analytic charts is an analytic diffeomorphism.

Another quite trivial remark is

2.9 Remark. Let f : X → ℝ be a function on an analytic manifold. The function is analytic if and only if for any analytic chart φ : U → V the function

f_{φ} = f ∘ φ^{−1} : V_{φ} ⟶ ℝ

is analytic in the usual sense (V_{φ} is open in some ℝ^{n}).

§3. The tangent space 7

This remark opens a new approach to the notion of an analytic manifold. An atlas A is called analytic if the transformations between any two charts of A are analytic dffeomorphisms. An analytic manifold is defined as a pair (X, A) consisting of a topological space X and an analytic atlas A. Then a function f : U → ℝ is called analytic if for every φ : U_{φ} → V_{φ} in A the transformed function f ∘ φ^{−1} (its domain of definition is the open set φ(U_{φ} ∩ U) ⊂ V_{φ}) is analytic. This gives a geometric structure which turns out to be an analytic manifold in in the sense we introduced. The elements of A are special analytic charts but not every analytic chart is contained in A. The set of all analytic charts is an analytic atlas which contains A. It is the unique maximal analytic atlas which contains A. This in mind one should call two atlases A, B equivalent if they belong to the same maximal one, which means nothing else but that A∪B is analytic two. An analytic manifold should be defined as a topological space equipped with a full equivalence class of analytic atlases ore equivalently a topological space equipped with a maximal analytic atlas. We feel that the trouble with “equivalence” or “maximality” makes the foundation by means of charts somewhat unpleasant and hence introduced the geometric structures. But the reader should be aware that the difference between the two approaches is inessential.

All what we need from these remarks is

2.10 Proposition. Let A be an atlas on a topological space such that the coordinate transformations between two charts of A are analytic diffeomorphisms. Then there exists a unique structure as analytic manifold on X such that the elements of A are analytic charts.

It is clear how to define this analytic structure: A function f : U → ℝ on an open subset U ⊂ X is called distinguished if the function f ∘ φ^{−1} (which is defined on φ(U_{φ} ∩ U)) is analytic in the usual sense for every φ ∈ A).

3. The tangent space

Let X be an analytic manifold and a ∈ X a point. We consider pairs (U, f) where a ∈ U ⊂ X is an open neighbourhood and f ∈ C^{ω}(U) is an analytic function on U. We denote by 𝔐(X, a) the set of all pairs. We introduce an equivalence relation: Two pairs (U, f) and (V, g) are called equivalent ((U, f) ≡ (V, f)) if there exists a small neighbourhood a ∈ W ⊂ U ∩ V such that f|W = g|W. We denote the equivalence class by

f_{a} = [U, f]_{a} := { (V, g); (V, g) ≡ (U, f) }.

The set of all equivalence classes is denoted by

C_{X,a}^{ω} := 𝔐(X, a)/ ≡ .

Its elements are called germs of analytic functions at a. This set carries a natural structure as ring. For example the sum (similarly the product) of two germs is defined by

[U, f]_{a} + [V, g]_{a} := [U ∩ V, f|(U ∩ V) + g|(U ∩ V)]_{a}.

It has to be mentioned that this definition is independent of the choice of the representatives (U, f) and (V, g). Clearly C_{X,a}^{ω} carries also a natural structure as ℝ-vector space. We notice two further obvious constructions:

1) Let a be a point in an open subset U ⊂ X. There is a natural homomorphism

C_{X}^{ω}(U) ⟶ C_{X,a}^{ω}, f ⟼ [U, f]_{a}.

2) There is a natural “evaluation map”

C_{X,a}^{ω} ⟶ ℝ, [U, f]_{a} ⟼ f(a).

It is trivial that this definition does not depend on the choice of the representative (U, f) but it should be mentioned that a is the only point where the value of a germ can be evaluated. (Germs have only one value. But they are more that this value!).

3) Let f : X → Y be an analytic map of analytic manifolds and a ∈ X. There is a natural pull-back homomorphism

C_{Y,f(a)}^{ω} ⟶ C_{X,a}^{ω}.

If V ⊂ Y is an open neighbourhood of b = f(a) and U = f^{−1}(V) then the diagram

C_{Y}^{ω} (V) ⟶ C_{X}^{ω}(U)
↓  ↓
C_{Y,f(a)}^{ω} ⟶ C_{X,a}^{ω}

commutes (where the first row means the natural pull-back). This property dictates the definition of C_{Y,f(a)}^{ω} → C_{X,a}^{ω}, hence we skip it.

We mention that the natural inclusion U → X of an open subset induces for a ∈ U an isomorphism

C_{U,a}^{ω} ⟶^{∼} C_{X,a}^{ω}.

We will use this isomorphism to identify both sides.

3.1 Definition. Let a ∈ X be a point on an analytic manifold X. A tangent vector A of X at a is map

A : C_{X,a}^{ω} ⟶ ℝ

with the following properties:

a) A is ℝ-linear.

b) A(fg) = f(a)A(g) + g(a)A(f).

We will see that this abstract definition is consistent with the geometric meaning of tangents. Property b) is similar to the product rule of calculus. Therefore tangent vectors sometimes are called derivations of C_{X,a}^{ω} with values in ℝ.

It is clear how to add tangent vectors and how to multiply them by constants. The set of all tangent vectors is a vector space which we call the tangent space of X at a and which we denote by

T_{a}(X) = tangent space of X at a.

The tangent space is “functorial” by a naiv construction:

3.2 Definition. Let f : X → Y be an analytic map of analytic manifolds and a ∈ X a point. The natural map C_{Y,f(a)}^{ω} → C_{X,a}^{ω} induces a linear map

T_{a}(f) : T_{a}(X) ⟶ T_{f(a)}(Y).

A tangent vector B ∈ T_{f(a)}(Y) is mapped to the composition of B and the natural homomorphism C_{Y,f(a)}^{ω} → C_{X,a}^{ω}.

Notice: The map C_{Y,f(a)}^{ω} → C_{X,a}^{ω} is kontravariant (opposite direction as f) and T_{a}(X) ⟶ T_{f(a)}(Y) is kovariant (same direction as f). Quite trivial is the “chain rule”:

3.3 Remark. Let f : X → Y and g : Y → Z be analytic maps of analytic manifold and a ∈ X. Then

T_{a}(g ∘ f) = (T_{f(a)}(g)) ∘ (T_{a}(f)).

We mention two further trivialities:

1) Let U ⊂ X an open subset of an analytic manifold and a ∈ U. The natural isomorphism C_{X,a}^{ω} ≅ C_{U,a}^{ω} induces an isomorphism

T_{a}(X) ≅ T_{a}(U).

We usually will identify both sides.

2) If f : X → Y is an analytic diffeomorphism then

T_{a}(f) : T_{a}(X) ⟶ T_{f(a)}(Y)

is an isomorphism. Both remarks together show: Let φ : U → V be an analytic chart on the analytic manifold X and a ∈ U a point with image point b = f(a). Then φ gives an isomoprphism

T_{a}(X) ≅ T_{b}(ℝ^{n}).

If one wants (for example to save letters) one can always choose the analytic chart φ such that b = 0 is the origin. The above remark shows that to understand the notion of the tangent space one has to understand the tangent space of ℝ^{n}. We define special tangent vectors:

The operators

∂_{ν} : C_{ℝ^{n},a}^{ω} ⟶ ℝ, [U, f]_{a} ⟼ \frac{∂f}{∂x_{ν}} (a),

are tangent vectors. One only has to mention that this definition is independent of the choice of the representative (U, f) of the germ. We should use the notation [∂_{ν}]_{a} instead of ∂_{a} if different points a are coming into the game.

3.4 Proposition. The tangent vectors

[∂_{1}]_{a}, …, [∂_{n}]_{a}

perform a basis of T_{a}(ℝ^{n}). Alternatively. The map

ℝ^{n} ⟶ T_{a}(ℝ^{n}), α ⟼ ∑_{ν=1}^{n} α_{ν}[∂_{ν}]_{a},

is an isomorphism of vector spaces.

Proof. We start with two preliminary observations:

1) Let \underline{1} be the germ represented by the function “constant 1”. Then

A(\underline{1}) = A(\underline{1} · \underline{1}) = A(\underline{1}) + A(\underline{1}) ⟹ A(\underline{1}) = 0.

As a consequence all “constant germs” (germs which can be represented by constant functions) are annihilated by tangent vectors.

2) Let f, g be two germs whose evaluation at a is zero, “f(a) = g(a) = 0”, then

A(fg) = 0.

The proof follows immediately from the product rule. Now we come to: Proof of 3.4. We consider a germ represented by the function f ∈ C^{ω}(U) for an open neighbourhood a ∈ U ⊂ ℝ^{n}. Because we are interested only in the germ of f we may replace U by a smaller neighbourhood and hence assume that f admits an expansion into a power series in the whole U. This gives us

f(x) = f(a) + ∑_{ν=1}^{n} C_{ν}(x_{ν} − a_{ν}) + ∑_{μ,ν}(x_{μ} − a_{μ})(x_{ν} − a_{ν})φ_{μν}(x)

where the functions φ_{μν} are analytic in U. Let A be a tangent vector. We denote by p_{ν} the projection function p_{ν}(x) = x_{ν} (also its germ at a) and consider the numbers

α_{ν} = A(p_{ν}).

From the two preliminary remarks we obtain A(f) = ∑ C_{ν}α_{ν}. We also have ∂_{ν}f(a) = C_{ν} and obtain

A = ∑_{ν=1}^{n} α_{ν}[∂_{ν}]_{a}.

The uniqueness of this representation is also clear.  ⊔ ⊓

Let

f : U ⟶ V, U ⊂ ℝ^{n}, V ⊂ ℝ^{m} (both open),

be an analytic map. Using the isomorphism 3.4 we obtain a commutative diagram

\begin{matrix}ℝ^{n} & ⟶ & ℝ^{m} \\ ∥ & & ∥ \\ T_{a}(U) & ⟶ & T_{f(a)}(V) \end{matrix} .

We obtain two linear maps ℝ^{n} → ℝ^{m}. The first one is given by T_{a}(f) the other is the map given by the Jacobi matrix

J(f, a) : ℝ^{n} ⟶ ℝ^{m}.

Recall that this Jacobi map is the linear map which maps x to y where

y_{i} = ∑_{j=1}^{m} α_{ij}x_{i}, J(f, a) = (α_{ij}).

If one writes x and y as columns x⃗ and y⃗ this means

y⃗ = J(f, a)x⃗ (matrix multiplication).

3.5 Remark. Let

f : U ⟶ V, U ⊂ ℝ^{n}, V ⊂ ℝ^{m} (both open)

be an analytic map and a ∈ U a point. The two linear maps ℝ^{n} → ℝ^{m} induced by 1. the tangent map T_{a}(U) → T_{f(a)}(V),

2. the Jacobi matrix J(f, a) agree.

The proof is a direct application of the chain rule in the form

\frac{∂(g ∘ f)}{∂x_{μ}}(a) = ∑_{ν=1}^{m} \frac{∂f_{μ}}{∂y_{ν}} (f(a))\frac{∂g}{∂x_{ν}} (a)

for an analytic function f on V.

Tangents at curves

Let be

α : I ⟶ X (I ⊂ ℝ)

an analytic curve, i.e. an analytic map of an open intervall I ⊂ ℝ into the analytic manifold X. Let t_{0} ∈ I be a fixed parameter and a = α(t_{0}) the corresponding point on the curve. We consider the tangent map

ℝ ≅ T_{t_{0}}(I) ⟶ T_{a}(X).

We recall that 1 ∈ ℝ corresponds to the tangent vector [d/dt]_{t_{0}} . We consider the image in T_{a}(X). This image is called the tangent vector of α with respect to the parameter t_{0}.

In general one can identify the set of all linear mappings Hom(ℝ, V) from ℝ into a real vector space V,

Hom(ℝ, V) ⟶^{∼} V, l ⟼ l(1).

In this sense we can say:

3.6 Definition. Let α : I → X be an anlytic curve. The tangent vector with respect to the parameter t_{0} is the vector in T_{α(t_{0})}, which corresponds to the tangent map

ℝ ≅ T_{α}(I) ⟶ T_{α(t_{0})}.

The tangent vector is denoted by α̇(t_{0}).

What does this mean in the case X = ℝ^{n}? Because of 3.5 the tangent vector corresponds to (dα/dt)(t_{0}), if one uses the identification T_{a}(ℝ^{n}) ≅ ℝ^{n}. This is precisely the geometric tangent vector, which justifies also the notion α̇ in the abstract context.

We see that the abstract notion of a tangent vector has a geometric background. We will see this also in the next section.

4. Sub-manifolds

Let X be an analytic manifold. We know that open subsets also are analytic manifolds. We want to generalize this.

4.1 Definition. A subset Y ⊂ X of an analytic manifold is called (analytic) smooth if for every point a ∈ Y there exists an analytic chart

φ : U_{φ} ⟶ V_{φ}, a ∈ U_{φ} ⊂ X, V_{φ} ⊂ ℝ^{n},

such that φ(Y ∩ U_{φ}) is the intersection of V_{φ} with a sub-vector space of ℝ^{n}.

It is no loss of generality to assume that this sub-vector space is defined by

x_{m+1} = … = x_{n} = 0 (for suitable 0 ≤ m ≤ n).

Open subsets are smooth for trivial reason. Also single points are smooth. Any sub-vector space of ℝ^{n} is smooth.

We define the restriction of a geometric structure O on X to an arbitrary subset Y ⊂ X: First we equip Y with the induced topology. A subset V ⊂ Y is called open, if there exist an open subset U ⊂ X such that V = U ∩Y. A function f : V → ℝ on an open subset V ⊂ Y is called distinguished, if for every point a ∈ Y there exists an open neighbourhood U(a) ⊂ X and an analytic function f_{a} : U(a) → ℝ such that f|(U(a) ∩ Y) = f_{a}|(U(a) ∩ Y). The three properties of a geometric structure are obvious. It should be mentioned that for open Y we obtain back the trivial restriction process considered earlier.

4.2 Remark. Let Y be a smooth subset of an analytic manifold X. If one equips Y with the induced geometric structure, then Y gets an analytic manifold.

We always will use this structure for Y and denote it by C_{Y}^{ω} := C_{X}^{ω}|Y. We also call Y an embedded analytic manifold. Clearly a smooth subset of Y is smooth in X. The proof of 4.2 is easy. One has to reduce it to the case X = ℝ^{n} and the linear subspace defined by x_{m+1} = … = x_{n} = 0. But in this case 4.2 is trivial.

We will use the following convenient

§4. Sub-manifolds 13

4.3 Definition. An analytic map of analytic manifolds f : Y → X is called an embedding, if f(Y) is smooth and if the induced map Y → f(Y) is an analytic diffeomorphism.

The composition of embeddings is an embedding. Immersions have a universal property:

4.4 Remark. Let f : Y ↪ X be an immersion of analytic manifolds. A map g : Z → Y of a third analytic manifold into Y is analytic if and only if g ∘ f is analytic.

This can be reduced to the case of the standard embedding ℝ^{n} → ℝ^{m} (n ≤ m)

instead of f.  ⊔ ⊓

A subset Y ⊂ X of a topological space X is called locally closed, if one of the following three equivalent properties is satisfied:

1) For every point a ∈ Y there exists an open neighbourhood a ∈ U(a) ⊂ X such that Y ∩ U(a) is closed in Y.

2) There exists an open subset U ⊂ X such that Y is closed in U.

3) Y is the intersection of an open and a closed subset of X.

We leave the proof of the equivalence as an exercise. Clearly open and closed subsets are locally closed.

4.5 Remark. Smooth subsets of analytic manifolds are locally closed.

We introduce another important notion:

4.6 Definition. An analytic map f : Y → X is called a local immersion at a ∈ Y, if the tangent map T_{a}(Y) → T_{f(a)}(X) is injective.

We reformulate the theorem of local immersions 1.3 in a more abstract context:

4.7 Remark. Let f : Y → X be a local immersion at a. Then there exists an open neighbourhood a ∈ V ⊂ Y such that the restricted mapping V → X is an embedding in the sense of 4.3.

Special case. In the special case that T_{a}(Y) → T_{f(a)}(X) is an isomorphism, f is a local analytic diffeomorphism, i.e. it induces an analytic diffeomorphism of some open neighbourhood of a to an open neighbourhood of f(a).

A very important notion is that of a global immersion:

4.8 Definition. An analytic map of analytic manifolds is called a (global) immersion if f is injective and a local immersion everywhere.

We have the trivial 4.9 Remark. Embeddings are immersions. Warning. The converse is false!

We sketch an example. For an real α ∈ ℝ we define

ℝ ⟶ ℝ^{4}, t ⟼ (sin(t), cos(t), sin(αt), cos(αt)).

This is a local immersion everywhere. It is injective if α is irrational. But in this case the image of ℝ dense in S^{1} × S^{1}, where S^{1} denotes the circle x^{2} + y^{2} = 1 in ℝ^{2}. But then the image can not be smooth.

How are embeddings to be distinguished from immersions? Fortunately very easy.

4.10 Proposition. Let f : Y → X an immersion such that f(Y) is locally closed. (This is true for an embedding.) Then f is an embedding.

Proof . Let b = f(a), a ∈ Y be point in the image. We choose open neighbourhoods a ∈ V ⊂ Y and b ∈ U ⊂ X such that f induces an embedding V ↪ U. In the case f(V) = U ∩ f(Y) we are done. We show that this is the case if we replace U by a smaller open neighbourhood and V by the inverse image: For this we can assume that f(Y) is closed. Then the complement A of U ∩ f(Y) in f(Y) is closed

We replace U by the intersection of U with the complement of A.  ⊔ ⊓

The geometric tangent space of an embedded manifold

Let f : X → ℝ^{n} be a local immersion at a ∈ X. Then we have an injective linear map T_{a}(X) ↪ T_{f(a)}(ℝ^{n}). By means of the standard identification T_{b}(ℝ^{n}) ≅ ℝ we get a linear subspace T ⊂ ℝ. This is the geometric tangent space of f at a. Better to say it is the subvector space which corresponds to the geometric tangent space. The “visible” tangent space is the affine translate f(a) + T. We can apply this to a smooth subset X ⊂ ℝ^{n} (more precisely to the canonical injection X → ℝ^{n}).

5. Vector fields

A vector field on a subset Y of an analytic manifold is a collection

A = (A_{a})_{a∈Y} , A_{a} ∈ T_{a}(X),

of tangent vectors. In most cases Y will be an open subset which then can assumed to be the whole X. Let A be a vector field on X and f : U → ℝ an analytic function on an open subset U ⊂ X. For arbitrary a ∈ U we may consider consider the germ f_{a} of f in a and the value A_{a}(f_{a}) ∈ ℝ. The assignment a ⟼ A_{a}(f_{a}) is a new function on U which sometimes we simply denote by

Af : U ⟶ ℝ, (Af)(a) := A_{a}(f_{a}).

We see that vector fields induce operators. A more careful notation of this operator is A_{U}f instead of Af.

§5. Vector fields 15

5.1 Definition. A vector field A on the analytic manifold is called analytic, if Af is analytic for every function f which is defined and analytic on some open subset of X.

Analytic vector fields hence induce operators in the following sense.

5.2 Definition. An operator A on an analytic manifold is a family of mappings A_{U} : C^{ω}(U) → C^{ω}(U) (U ⊂ X open) with the following properties:

1. A_{U} is ℝ-linear for all U.

2. A_{U} is compatible with restriction, i.e. A_{U}(f)|V = A_{V} (f|V) for open V ⊂ U ⊂ X and f ∈ C^{ω}(U).

The set of all operators is a vector space. Moreover it is clear how o compose operators. This means that the set of all operators on X is a ring.

The operators which come from a vector field satisfy the product rule A(fg) = fA(g) + gA(f). This gives rise to

5.3 Definition. Let X be an analytic manifold. A derivation A on X is an operator with the following property.

A_{U}(fg) = fA_{U}(g) + gA_{U}(f) (f, g ∈ C_{X}^{ω}(U)).

We have seen that every analytic vector field induces a derivation. The converse is also true.

5.4 Remark. Every derivation is induced by a unique analytic vector field.

For the proof we have to describe the converse map. Let A be a derivation. We reconstruct the vector field. Let a be a point and f_{a} a germ of an analytic function f ∈ C_{X}^{ω}(U) in some open neighourhood of a ∈ U ⊂ X. The number A_{U}(f)(a) does not depend on the choice of the representative of the germ (because of 5.2 , 2). We

obtain a tangent vector and even more a vector field.  ⊔ ⊓

The Lie bracket

Let A, B be two derivations. We can consider the operator A ∘ B, which means the family (A∘B)_{U} := A_{U} ∘B_{U}. Usually this family will not be a derivation, the product rule may be violated. But the operator [A, B] := A ∘ B − B ∘ A is a derivation.

5.5 Remark. There is a natural pairing between analytic vector fields, so so-called Lie bracket. On the level of derivations it is given by

[A, B] := A ∘ B − B ∘ A.

The set of all analytic vector fields on X is not only a vector space but a module over the ring C_{X}^{ω}(X). The product of an element f ∈ C_{X}^{ω}(X) and an analytic vector field is defined by (fA)_{a} := f(a)A_{a}. On the level of derivations the formula is (fA)_{U} = (f|U)A_{U}. For any open subset U ⊂ ℝ^{n} the operator

∂_{ν} : C_{ℝ^{n}}^{ω}(U) ⟶ C_{ℝ^{n}}^{ω}(U) (1 ≤ ν ≤ n

has been defined. This is actually a derivation on ℝ^{n} and also on any open subset of ℝ^{n}.

5.6 Remark. Let U ⊂ ℝ^{n} be an open subset. The set of all analytic vector fields on U is a free module over C_{ℝ^{n}}^{ω}(U) with basis ∂_{1}, …, ∂_{n}. This means that analytic vector fields can be identified with elements of C^{ω}(U)^{n} by means of f ↦ ∑ f_{i}∂_{i}.

Let ∑_{i=1}^{n} f_{i}∂_{i}, ∑_{j=1}^{n} g_{j}∂_{j} be two analytic vector fields. The Lie bracket is ∑_{k=1}^{n} h_{k}∂_{h} where

h_{k} = ∑_{i=1}^{n} f_{i}∂_{i}g_{k} − g_{i}∂_{i}f_{k}.

This is the so-called Poisson bracket. A special case is [∂_{μ}, ∂_{ν}] = 0 which we already know because the partial derivatives commute.

Chapter II. Basic properties of Lie groups and Lie algebras

1. Basic properties of Lie groups

We need the notion of a cartesian product of analytic manifolds. Let X, X′ be two analytic manifolds. We equip X × X′ with the product topology. The open sets of X × Y are arbitrary unions of open squares. By an open square we understand a set of the form U × U′, where U ⊂ X and U′ ⊂ Y are open. Assume that U and U′ are domains of definitions of analytic charts

φ : U → V, φ′ : U′ → V′, (V ⊂ ℝ^{n}, V′ ⊂ ℝ^{m}).

Then

φ × φ′ : U × U′ ⟶ V × V′ ⊂ ℝ^{n+m}.

is a chart on X × X′. The set of all these charts is an analytic atlas on X × Y, which by I.2.10 induces an analytic structure C_{X×Y}^{ω}. The analytic space X × Y has an obvious universal property, which characterizes the structure on X × Y. Let Z be a third analytic manifold. A map Z → X × Y is analytic if and only of the compositions with the two projections X × Y \overrightarrow{⟶}X, Y are analytic. Especially the two projections are analytic. We come to the definition of Lie group:

1.1 Definition. A Lie group G is a group equipped with a structure as analytic variety such that the maps

\begin{matrix}G × G ⟶ G, (x, y) & ⟼ xy, \\ G ⟶ G, x & ⟼ x^{−1}, \end{matrix}

are analytic.

Let a ∈ G a fixed element. The translation maps

G ⟶ G, x ⟼ ax, x ⟼ xa,

are analytic diffeomorphisms.

1.2 Definition. A homomorphism of Lie groups f : G → H is a group homomorphism which is analytic as well.

Clearly an isomorphism of Lie groups means an isomorphism of groups which is also an analytic diffeomorphism.

Examples of Lie groups

1) The additive group ℝ^{n} is a Lie group.

2) The multiplicative group ℝ^{•} is a Lie group.

3) First we remark that every finite dimensional vector space V over ℝ carries a structure as an analytic manifold. This is defined by means of the choice of an isomorphism V → ℝ^{n} but the structure does not depend on this choice. The additive group of V is a Lie group. Every complex vector space can be considered also as real vector space. Hence also every finite dimensional complex vector space can be considered as an analytic manifold and its additive group is a Lie group.

In this connection we want to mention that most of the calculus can be generalized to abstract vector spaces. Let A, B be two finite dimensional vector spaces, U ⊂ A and V ⊂ B open subsets and f : U → V an analytic map. The Jacobi map J(f, a) : A → B can be defined for example using isomorphims A ≅ ℝ^{n} and B ≅ ℝ^{m}. The Jacobi map does not depend on the choice of the isomorphisms. The reason is the fact that the Jacobi map can be understood as the best linear approximations f near a. Also the isomorphism V ≅ T_{a}(V) inherited from the natural isomorphism T_{a}(ℝ^{n}) ≅ ℝ^{n} is intrinsic. The analogue of I.3.5 is true in the abstract context.

4) The set GL(n, ℝ) of all invertible n × n-matrices is an open subset of the vector space of all n × n-matrices and hence an analytic manifold. It is a Lie group. Also GL(n, ℂ) is a Lie group.

5) Let G be a subgroup and H ⊂ G an abstract subgroup which is also smooth. Then H is a Lie group (with the induced analytic structure).

We will see later that every closed subgroup of a Lie group is smooth. This gives many examples of Lie groups, because there are many closed subgroups of GL(n, ℝ), GL(n, ℂ) (the classical groups as the orthogonal, unitary, symplectic groups).

6) Consider the equivalence relation

x ∼ y ⟺ x − y ∈ ℤ^{n}

on ℝ^{n}. The set X = ℝ^{n}/ℤ^{n} of all equivalence classes is equipped with the quotient topology. A set in X is called open if its inverse image in ℝ is open. A function on an open subset on X is called analytic, if its pull-back to the inverse image in ℝ^{n} is analytic. This equips X with a structure as analytic manifold. There is also a structure of a group on X such that the canonical projection ℝ^{n} → X is a group homomorphism. We see that we obtain a Lie group X, the so-called torus. The canonical map ℝ^{n} → X is a homomorphism of Lie groups.

The cartesian product G × H of two Lie groups is a group an an analytic manifold as we mentioned above. It is actually a Lie group. In a similar way the product G = G_{1} × ⋯ × G_{n} of finitely many Lie groups is defined as a Lie group. If we denote the torus constructed in 6), we see

X_{n} = X_{1} × ⋯ × X_{1}.

The group X_{1} can be constructed in a different way. The set S^{1} of all complex numbers of absolute value 1 is smooth in ℂ ≅ ℝ^{2}. It is also a group under usual

§2. The Lie algebra 19

multiplication of complex numbers. The homomorphism ℝ → S^{1} given by t ↦ exp(2πit) induces an isomorphism of Lie groups S^{1} ≅ X_{1}.

Elements of GL(n, ℂ) can be considered as isomorphisms of the complex vector space ℂ^{n}. This is also a real vector space. Using an isomorphism ℂ^{n} ≅ ℝ^{2n} one obtains an embedding GL(n, ℂ) ↪ GL(2n, ℝ).

Up to now we constructed only Lie groups which sit inside some matrix group GL(n, ℝ). For example the additive group ℝ can by identified with the group of real matrices \binom{1 t}{0 1}. The product of GL(n, ℝ)×GL(m, ℝ) can be considered as subgroup of GL(n + m, ℝ) via (A, B) ↦ \binom{A 0}{0 B}. It seems worth while to give an example of a Lie group which is not embeddable as matrix group:

7) This example needs a little complex calculus. We use The group G = SL(2, ℝ) of all real 2 × 2-matrices of determinant 1 and the upper half plane H consisting of all complex numbers with positive imaginary part. This is a smooth subset of GL(2, ℝ), hence also a Lie group. We use the following to facts:

a) The group SL(2, ℝ) acts on H: For M = \binom{a b}{c d} ∈ SL(2, ℝ) and z ∈ H the expression

Mz := \frac{az + b}{cz + d}

is also contained in Hand satisfies (MN)z = M(Nz).

b) For every M ∈ SL(2, ℝ) the function cz+d on H has a continuous logarithm A(z). This is a continuous function on H with the property

exp(A(z)) = cz + d.

We define G̃ to be the set of all pairs (M, A), where M ∈ SL(2, ℝ) and where A is a continuous logarithm of cz+d. We define multiplication by

(M_{1}, A_{1})(M_{2}, A_{2}) = (M_{1}M_{2}, A), A(z) = A_{1}(M_{1}z)A_{2}(z).

It takes some computation to show that G̃ is a group. Then one has to show that G̃ carries a structure as analytic manifold. This structure is characterized by the fact that the natural projection G̃ → G is analytic and moreovere a local immersion everywhere. It can be shown that G̃ is connected. Actually it is the so-called universal covering of G. It is a Lie group which admits no immersion into any GL(n, ℝ).

2. The Lie algebra

Let G be a Lie Group. We denote by e the unit element of G and by

𝔤 := T_{e}(G)

the tangent space at this distinguished point. We will use the operator “left translation”

L_{a} : G ⟶ G, L_{a}(x) = ax,

to move the distinguished point e to an arbitrary one. This operator induces an isomorphism

T_{e}(L_{a}) : 𝔤 = T_{e}(G) ⟶^{∼} T_{a}(G).

More generally we have for two a, b ∈ G an isomorphism

T_{a}(L_{ba^{−1}}) : T_{a}(G) ⟶^{∼} T_{b}(G).

2.1 Definition. A vector field A on a Lie group G is called left invariant id the following condition is satisfied.

Let a, b ∈ G be two group elements. Then A_{a} is mapped to A_{b} under the isomorphism T_{a}(G) → T_{b}(G), which is induced by multiplication with ba^{−1} from the left.

A first observation is:

2.2 Remark. Left invariant vector fields are analytic.

The proof follows from the fact that left translations are analytic: Consider an analytic function f : U → ℝ on some open subset U ⊂ G and an left invariant vector field A. We compute the function g := A_{U}(f) in some detail. By definition is given as follows: g(a) = A_{a}(f_{a}), where f_{a} is the germ of f at a. From the left invariance we get g(a) = ((T_{e}(L_{a}))(A_{e}))(f_{a})). Now we use the definition of the tangent map (I.3.2) and get g(a) = A_{e}(f_{e}^{a}), where we use the notation

f^{a}(x) = f(ax)

This function is defined and analytic on a^{−1}U, which is an open neighbourhood of e. The function (x, a) ↦ f^{a}(x) is analytic on some open subset of G × G. This follows from the definition of Lie groups. Now it is clear that g is analytic. One uses local coordinates.

2.3 Remark. A vector field on a Lie group is left invariant if and only if the following condition is satisfied: Let f_{U} → ℝ be an analytic function on an open subset U ⊂ G. Then

A_{U}(f)^{a} = A_{a^{−1}U}(f^{a}) for all a ∈ G.

Let L ∈ 𝔤 = T_{e}(G). We can extend A to a vector field

Ã : Ã_{a} := T_{e}(L_{e})(A).

This is a vector field which extends A in the sense Ã_{e} = A. The vector field Ã is left invariant this follows from the chain rule applied to L_{ba^{−1}} ∘ L_{a} = L_{b}. Thus we obtain:

§3. The exponential map 21

2.4 Proposition. The vector space of all left invariant vector fields on a Lie group can be identified with the tangent space 𝔤 at the unit element. To obtain the identification on has to map a vector field Ã to the tangent vector A = Ã_{e}.

Frequently we will use the same letter for the vector field and the the corresponding element in 𝔤, i.e. we omit tildes.

Let A, B be elements of 𝔤. We consider them as vector fields. Recall that we defined the Lie-bracket [A, B] = AB − BA. It is clear that this vector field is left invariant again. Hence it corresponds to an element of 𝔤. We defined a map

𝔤 × 𝔤 ⟶ 𝔤 (A, B) ⟼ [A, B].

This is a fundamental algebraic structure which leads to the notion of a Lie algebra.

2.5 Definition. A Lie algebra over a field K is a vector space V equipped with a symmetric bilinear form

𝔤 × 𝔤 ⟶ 𝔤 (A, B) ⟼ [A, B]

with the properties

1. [A, A] = 0.

2. [[A, B], C] + [[B, C], A] + [[C, A], B] = 0 (Jacobi identity).

The Lie bracket of vector fields satisfies 1) and 2). We see The tangent space 𝔤 of a Lie algebra at the unit element carries a structure as Lie algebra over the field of real numbers.

3. The exponential map

3.1 Definition. Let A be a vector field on an analytic manifold X. An integral curve of A is an analytic map

I ⟶ X, I ⊂ ℝ an open interval,

with the property

α̇_{A}(t) = A_{α(t)}.

The integral curve is called maximal if it cannot be extended to an integral curve on a bigger Intervall.

Integral curves satisfy a differential equation: Let X = U ⊂ ℝ^{n} be an open subset and

A = f_{1}∂_{1} + ⋯ + f_{n}∂_{n}.

Then α is an integral curve if

\frac{dα_{ν}(t)}{dt} = f_{ν}(α(t)) (1 ≤ ν ≤ n).

This is a system of ordinary differential equations. Using the well-known theorem from the existence and uniqueness of an ordinary system of differential equations one concludes:

3.2 Proposition. Let A be an analytic vector field on an analytic manifold X and a ∈ X a given point. There exists a unique maximal integral curve α : I → X with the property

0 ∈ I, α(0) = a.

The existence of an integral curve with not necessarily maximal I follows from the local existence theorem. Two integral curves I → X, J → X agree in the intersection. This follows from the local uniqueness theorem. The extension to a maximal I is now easy to prove. On uses that the union of intervals which contain 0 is an interval again.

We apply this in a special case. Let G be a Lie group and A be an element of the Lie algebra 𝔤 which we consider as left invariant vector field. We can consider the maximal integral curve with the initial value α(0) = e.

3.3 Theorem. Let G be a Lie group G and A ∈ 𝔤 an element of the Lie algebra considered as left invariant vector field. There exists a unique integral curve

α_{A} : ℝ ⟶ G, α(0) = e.

This is a homomorphism of Lie groups, α_{A}(t + s) = α_{A}(t)α_{A}(s). Additional Remark. The map

ℝ × 𝔤 ⟼ G, (t, A) ⟼ α_{A}(t)

is analytic.

Proof. We consider the maximal open interval 0 ∈ I ⊂ ℝ such that the integral curve α with α(0) = e and α̇(t) = A_{α(t)} can be defined on I. We take an r > 0 such that I_{r} ⊂ I, where I_{r} = (−2r, −2r). We claim α(t′ + t) = α(t′ + t) if |t|, |t|′ < r. For the proof we fix t′ and consider t as variable. We consider β(t) = α(t′)^{−1}α(t′ + t). We compute the tangent map of β applying the chain rule to

ℝ ⟶^{t↦t′+t} ℝ ⟶^{α} G ⟶^{L_{α(t′)^{−1}}} G.

We fix a t_{0} ∈ ℝ. The image of [d/dt]_{t_{0}} ∈ T_{t_{0}}(ℝ) in T_{t′+t_{0}}(ℝ) under the map induced by t ↦ t′ + t is [d/dt]_{t′+t_{0}}. This follows from the fact that the derivative of the translation map is 1. Next we consider the image of [d/dt]_{t+t′_{0}} in T_{α(t′+t_{0})}(G). This is α̇(t′ + t_{0}) by definition of α̇. The image in T_{α(t′)}(G)

(T_{α(t′+t_{0})}(L_{α(t′)^{−1}}))(α̇(t′ + t_{0})) = (T_{α(t′+t_{0})}(L_{α(t′)^{−1}}))(A_{α(t′+t_{0})}) = A_{β(t_{0})}.

The last equation comes from the left invariance of the vector field associated to A. Thus we proved β̇(t_{0}) = A_{β(t_{0})}. We also have β(0) = e. The uniqueness theorem gives α = β which was the claim. We also see α(t) = α(t/2)^{2}. The right hand side is defined on I_{2r} and it is easy to prove that α is an integral curve on this bigger interval. This shows that the integral curve can extended successively to the intervals I_{2r}, I_{4r}, . . .. We obtain shows I = ℝ.

The additional remark has to be proved locally and is then a consequence of standard results on the analytic dependence of the solutions of differential equations on parameters. We skip details. ⊔ ⊓

It easy to see that every analytic homomorphism α : ℝ → G is an integral curve. One considers for an fixed t′ the map t ↦ α(t′ + t) = L_{α(t′)}α(t) and computes its derivative at t = 0. The result is α̇(t′) = T_{e}(L_{α(t′)}α̇(0)). We obtain:

3.4 Proposition. There is a one-to-one correspondence between the elements A of the Lie algebra and the analytic homomorphisms α : ℝ → G, given by α(t) = α_{A}(t).

The map t ↦ α_{A}(t′t) is also a homomorphism. Hence 3.4 implies:

Corollary. We have

α_{t′A}(t) = α_{A}(t′t).

We evaluate the integral curve α at the special value 1:

3.5 Definition. The exponential map

exp : 𝔤 ⟶ G

of the Lie algebra of a Lie group into the Lie group is defined as exp(A) = α_{A}(1).

We know already that exp is an analytic map.

We collect some basic facts about the exponential map:

3.6 Lemma. Let G by a Lie group and exp : 𝔤 → G the exponential map. For A ∈ G the map

ℝ ⟶ G, t ⟼ exp(tA),

coincides with the unique integral curve α_{A} (s. 3.3) of the corresponding left invariant vector field. As a consequence this map is an analytic homomorphism of groups (and every analytic homomorphism is of this kind with a unique A).

Proof. We have exp(tA) = exp_{tA}(1) = exp_{A}(t) by the corollary of 3.5 .  ⊔ ⊓

3.7 Proposition. The exponential map g → G of a Lie group is a local immersion at 0.

Proof. Consider an element A ∈ 𝔤 and the associated linear map ℝ → 𝔤, which maps t to tA. The tangent map ℝ ≅ T_{0}(𝔤) defines an element of T_{0}(𝔤). This defines an isomorphism 𝔤 ≅ T_{0}(𝔤)*). We obtain a map

𝔤 ≅ T_{0}(𝔤) ⟶^{T_{0}(exp)} T_{e}(G) ≅ 𝔤.

*) We mentioned already in example 3 of section 1 that the tangent space of a finite dimensional vector space V can be identified canonically with the vector space. The isomorphism mentioned there and the isomorphism which we use here (coming from differentiating t ↦ ta for a ∈ V) agree.

We have to show that this map β : 𝔤 → 𝔤 is an isomorphism. We claim that it is in fact the identity map: We compose the map t ↦ tA with the exponential map

ℝ ⟶^{t↦tA} 𝔤 ⟶^{exp} G.

The image of the tangent vector [d/dt]_{0} in 𝔤 ≅ T_{e}(𝔤) is A. The image in T_{e}(G) is the derivative of the integral curve t ↦ exp(tA) at t = 0, which is also A. From the

chain rule follows that the image of A ∈ 𝔤 ≅ T_{0}(𝔤) in 𝔤 = T_{e}(G) is A.  ⊔ ⊓

3.8 Proposition. Let f : G → H be a homomorphism of Lie groups. The induced tangent map φ : 𝔤 → 𝔥 is a homomorphism of Lie algebras. The diagram

\begin{matrix}G & ⟶ & H \\ ↑ & & ↑ \\ 𝔤 & ⟶ & 𝔥 \end{matrix}

commutes. (The vertical arrows are the exponential maps.)

Let A ∈ 𝔤, B = φ(A) and Ã, B̃ the corresponding vector fields. We claim

(T_{a}(f))(Ã) = B̃_{f(a)} (a ∈ G).

This is a consequence of the chain rule applied to the equation f ∘ L_{a}^{G} = L_{f(a)}^{H} ∘ f. Here L^{G}, L^{H} denotes left translation on G, H. For the corresponding derivations this means

Ã(h ∘ f) = B̃(h) ∘ f,

where h is an analytic function on some open subset of H. Let A′ ∈ 𝔤 be a second element and use B′, Ã′, B̃′ in the obvious sense. Then we get

[Ã, Ã′](h ∘ f) = [B̃, B̃′](h) ∘ f

or

(T_{a}(f))([Ã, Ã′]) = [B̃, B̃′]_{f(a)}.

Evaluating at a = e gives φ([A, A′]) = [B, B′]. Hence φ is a homomorphism of Lie algebras.

We also see that f(exp(tA)) is the integral curve corresponding to B, which shows f(exp(tA)) = exp(tB). Evaluating at t = 1 gives the claimed commutativity of the diagram.

Some examples of Lie groups and Lie algebras

The additive group G = ℝ is a Lie group. The Lie algebra 𝔤 can be identified with T_{0}(ℝ) ≅ ℝ. This algebra is abelian which means by definition [A, B] = 0 for all A, B. The exponential map exp : 𝔤 → G corresponds just to the identity map

id_{ℝ} : T_{0}(ℝ) ≅ ℝ ⟶ ℝ.

The one-parameter subgroups simply are the functions t ↦ at.

The (multiplicative) group G = ℝ_{>0} of positive is an open subset of ℝ. The tangent space 𝔤 at 1 can be identified with ℝ. It also an abelian Lie-algebra. The exponential map exp : 𝔤 → G now corresponds to the usual exponential function

T_{1}(ℝ_{>0}) ≅ ℝ ⟶ ℝ_{>0}, x ⟼ e^{x}.

By the way the groups (ℝ, +) and (ℝ_{>0}, ·) are isomorphic, the isomorphism is given is given by the exponential (logarithm).

A basic example

The group GL(n, ℝ) is an open subset of the vector space M_{n}(ℝ) of n×n matrices. Recall that we have natural isomorphisms

T_{e}(GL(n, ℝ)) ≅ T_{e}(M_{n}(ℝ)) ≅ M_{n}(ℝ).

We denote by 𝔤𝔩(n, ℝ) the Lie algebra of M(n, ℝ) We constructed a natural isomorphism of vector spaces

𝔤𝔩(n, ℝ) ≅ GL(n, ℝ).

We use the matrix exponential function

M_{n}(ℝ) ⟶ GL(n, ℝ), A ⟼ e^{A} := ∑_{m=0}^{∞} \frac{A^{m}}{m!} .

It is easy to prove that this series converges and defines an analytic function M_{n}(ℝ) → M_{n}(ℝ). It also easy to prove that

e^{A+B} = e^{A}e^{B} if AB = BA.

The standard proof of analysis (reduction to a the binomial identity) works when A and B commute. Hence t ↦ e^{tA} is an analytic homomorphism. Its derivative at t = 0 is A. Hence we obtain

e^{tA} = exp(tA)

if we identify T_{e}(GL(n, ℝ)) with M_{n}(ℝ).

We want to determine the Lie bracket. For this we need the action of the derivation Ã which corresponds to an element A ∈ M_{n}(ℝ). It is sufficient to compute Ã(L) for a linear function L. One needs the formula

(Ã(L))(X) = L(XA).

We leave this formula as an exercise to the reader. It also follows from the general formula 4.1 which we derive in the next section. The function (Ã(L))(X) is linear again and we get (ÃB̃(L))(X) = L(XAB). We obtain

(ÃB̃(L))(X) − (B̃Ã(L))(X) = L(X(AB − BA))

for linear functions. Hence ÃB̃ − B̃Ã and ˜ [A, B] have the same effect on linear functions. Because they are derivations they agree. We see:

3.9 Proposition. If one identifies the Lie algebra 𝔤𝔩(n, ℝ) of GL(n, ℝ) with M_{n}(ℝ) using the natural isomorphism above then

1. The Lie bracket of 𝔤𝔩(n, ℝ) corresponds to [A, B] = AB − BA on M_{n}(ℝ).

2. e^{A} = exp(A) .

In a similar way one obtaines 3.10 Proposition. The Lie algebra 𝔤𝔩(n, ℂ) of GL(n, ℂ) can be identified with M_{n}(ℂ) such that

1. The Lie bracket of 𝔤𝔩(n, ℂ) corresponds to [A, B] = AB − BA on M_{n}(ℂ).

2. e^{A} = exp(A) .

Many Lie groups can be obtained as closed subgroups of GL(n, ℝ), GL(n, ℂ). We will see a little later (III.2.1) that every closed subgroup H of a Lie group G is smooth and carries itself a structure as a Liegroup. The Lie algebra of H can be identified with the set of all A ∈ 𝔤 such that exp(tA) is contained in H for all t ∈ ℝ. We give an example. Let A be an invertible real (complex) n × n-Matrix. The set G(A) of all real (complex) n × n matrices g with the property g′Ag = A is a group, actually a closed subgroup of GL(n, ℝ) (GL(n, ℂ)), hence a Lie group and its Lie algebra consists of all real (complex) matrices X with the property e^{t}X′A = A^{t}X. It is easy to prove that this equation means

AX + X′A = 0.

The most interesting cases are that A is symmetric (A′ = A) or alternating (A′ = −A). In the first case G(A) is called an orthogonal group in the second one a symplectic group. In the orthogonal case a matrix X is in the Lie algebra if AX is alternating, in the symplectic case iif AX is symmetric.

We can use the theory of the Lie algebra to determine the dimension of the group G(A):

The dimension of an analytic manifold X at a point a is

dim_{a} X := dim T_{a}(X).

It is clear that that the function a ↦ dim_{a} X is locally constant. We call X pure dimensional, if this function is constant and call dim X := dim_{a} X the dimension of X in this case. We have

a) Connected manifolds are pure dimensional.

b) Lie groups are pure dimensional.

The latter statement follows from the fact that translations are diffeomorphisms.

3.11 Remark. Let A be an invertible real n×n-matrix. The group G(A) consisting of all real solutions of g′Ag = A is a Lie group of dimension n(n − 1)/2 in the symmetric and n(n + 1)/2 in the alternating case case. In the complex case the dimensions have to be doubled.

The reader is invited to prove this without Lie theory.

§4. The Taylor expansion of the exponential map 27

4. The Taylor expansion of the exponential map

Let G be a Lie group, exp : 𝔤 → G the exponential, f an analytic function on some open neighbourhood of e ∈ G and A ∈ 𝔤 an element of the Lie algebra. We denote the corresponding left invariant vector field by Ã. We apply the chain rule to the situation

ℝ ⟶^{exp(tA)} G ⟶^{f} ℝ.

More precisely we evaluate the tangent map at 0 ∈ ℝ. The tangent map of the composition is given by the number

\frac{df(exp(tA))}{dt} t = 0 .

The image of 1 ∈ ℝ ≅ T_{0}ℝ under the first map A. We have (T_{e}(f))(A) = (Af)(e). This is the definition of the vector field which corresponds to A. We obtain the simple formula

\frac{df(exp(tA))}{dt} t = 0 = A(f) = (Ã(f))(e).

A slight and obvious generalization (using the left invariance of Ã in the form 2.3) is

\frac{df(g exp(tA))}{dt} t = 0 = (Ã(f))(g).

We apply this formula inductively to Ãf, Ã(Ãf), … and obtain

4.1 Remark. Let f : U → ℝ be an analytic function on some open subset U ⊂ G and A ∈ 𝔤 be an element of the Lie algebra. Then the formula

\frac{d^{n}f(g exp(tA))}{dt^{n}} t = 0 = (Ã^{n}(f))(g)

holds. Corollary The Taylor expansion of f(g exp(tA)) around t = 0 (for fixed g and A) is of the form

f(g exp(tA)) = ∑ \frac{a_{m}}{m!} t^{m}, a_{m} = (Ã^{n}(f))(g).

(This expansion is valid in a small neighbourhood (−ε, ε) of t = 0.) Additional Remark. For every fixed g_{0} ∈ G and A_{0} ∈ 𝔤 there exist small open neighbourhoods such that ε can be choosen as the same for all A, g in these neighbourhoods.

The additional remark is a consequence that f(g exp(tA)) is real analytic on G × ℝ × 𝔤 by the general permanence properties of analycity. The Taylor series must be the specialization of an expansion in “all variables”. Another application of 4.1 gives 4.2 Lemma. Let A, B be elements in 𝔤 and f an analytic function in some open neighbourhood of e ∈ G. The function f(exp(tA) exp(sB)) is real analytic in some neighbourhood of (t, s) = (0, 0). Its Taylor expansion is of the form

f(exp(tA) exp(sB)) = ∑_{n,m}((Ã^{n}B̃^{m})(f))(e) \frac{t^{n}s^{m}}{n!m!} .

We want to use the Taylor expansion to obtain information about the product exp(tA) exp(tB). We recall that there exists small open neighbourhoods 0 ∈ V ⊂ 𝔤 and e ∈ U ⊂ G such that exp defines an analytic diffeomorphism V → U. We denote the inverse map by log : U → V. For sufficiently small t the function

C(t) := log(exp(tA) exp(tB))

is defined and analytic in t. The values of C are in 𝔤. We consider the Taylor expansion of C,

C(t) = C_{0} + tC_{1} + C_{2} \frac{t^{2}}{2} + · · · .

We want to compute the first coefficients, more precisely we claim

C_{0} = 0, C_{1} = A + B, C_{2} = [A, B].

The first one C_{0} = 0 is clear. The others have to be drawn from the relation

exp(C(t)) = exp(tA) exp(tB).

Instead of this we consider

f(exp(C(t))) = f(exp(tA) exp(tB))

for an arbitrary analytic function in some open neighbourhood of e. We want apply 4.1 and write C(t) = tD(t), where D is analytic in a small neighbourhood of 0. From 4.1 (including the additional remark) we get

\begin{aligned}f(exp(C(t)) & = ∑_{n=0}^{∞} \frac{D̃(t)^{n}(f)}{n!}(e) t^{n} \\ & = f(e) + (C̃_{1}(f))(e) t + (C̃_{2}(f))(e) t^{2} + \frac{(C̃_{1}^{2}(f))(e)}{2} t^{2} + · · · . \end{aligned}

The dots includes all higher powers of t. Comparing this expression with 4.2 (specialized to t = s) we obtain

\begin{aligned}(C̃_{1}(f))(e) & = (Ã(f))(e) + (B̃(f))(e), \\ (C̃_{2}(f))(e) + \frac{(C̃_{1}^{2}(f))(e)}{2} & = \frac{(Ã^{2}(f))(e)}{2} + \frac{(B̃^{2}(f))(e)}{2} + ((ÃB̃)(f))(e). \end{aligned}

The first equation means C_{1} = A + B and this equation implies C̃_{1} = Ã + B̃. Squaring yields C̃_{1}^{2} = Ã^{2} + B̃^{2} + ÃB̃ + B̃Ã. Now we see C_{1} = A + B, C_{2} = [A, B] as has been claimed. We formulate this and two similar results in

§4. The Taylor expansion of the exponential map 29

4.3 Proposition. Let A, B be two elements of the Lie algebra 𝔤 of a Lie group G. Then there exists ε > 0 such that

a) exp(tA) exp(tB) = exp\{t(A + B) + \frac{t^{2}}{2} [A, B] + t^{3}X(t)\}

b) exp(−tA) exp(−tB) exp(tA) exp(tB) = exp{t^{2}[A, B] + t^{3}Y (t)}

c) exp(tA) exp(tB) exp(−tA) = exp{tB + t^{2}[A, B] + t^{3}Z(t)} for |t| < ε. Here in each case X, Y, Z denotes an analytic function in (−ε, ε) with values in 𝔤. Let A, B be two elements of a Lie algebras 𝔤 with the property [A, B] = 0. The the operators Ã and B̃ commute. A standard argument using 4.2 gives 4.4 Remark. Assume [A, B] = 0. Then exp(A + B) = exp(A) exp(B).

Chapter III. Relation between Lie groups and Lie algebras

1. Simple topological facts about Lie groups

We collect some simpel topological facts about Lie groups:

1.1 Lemma. Let G_{0} ⊂ G be an open subgroup of topological group (for example a Lie group). Then G_{0} is also closed.

Proof. The complement G − G_{0} is the union of all open sets gG_{0}, g ∈ G − G_{0}, hence

open.  ⊔ ⊓

1.2 Lemma. Let G_{0} be a locally closed subgroup of a topological group G. Then G_{0} is closed.

Proof. It is easy to see that the closure H of G_{0} is a subgroup of G. A locally closed set always is open in its closure. This follows immediately from the definition of “locally closed”. Hence G_{0} is an open subgroup of H, hence closed in H, hence closed in G. ⊔ ⊓

We recall the notion of arc-component of a topological space. Two points of X are called equivalent if the are contained in an arc. The equivalence classes with respect to this equivalence relation are called the arc components. If X is a manifold then the arc components obviously are open. The are of course connected and we call them simply the connected components. They are not only open but closed because their complement is a union of connected components. (By the way this shows that every connected manifold is also arcwise connected as we mentioned earlier.)

1.3 Lemma. The connected component G^{0} of the identity element of a Lie group is a subgroup. It is generated as abstract group by any non-empty neighbourhood e ∈ U ⊂ G^{0} of the unity.

Proof. If let a(t), b(t) we arcs whose image contains e and a, b. We can assume a(0) = b(0) = e and a(1) = a, b(1) = b. Then a(t)b(t)^{−1} is an arc which contains ab^{−1}. This proves the first statement. To prove the second statement, we assume that U is closed under a → a^{−1}. This is no loss of generality, because we can replace U by U ∪ U^{−1}, which is connected too. We consider the image of the cartesian product U^{n} → G under the multiplication map. It is easy to see that this set is open. For example the image of U × U is the union of all gU. The union of all images of U^{n} is the group H generated by U. hence H is open in G^{0} and hence also closed. We obtain H = G^{0} because G^{0} is connected.

§2. Closed subgroups of Lie groups 31

2. Closed subgroups of Lie groups

Let G be Lie group with Lie algebra 𝔤 and H a sub group, which is also smooth as subset. Then H is also a Lie group. The canonical injection H → G induces a homomorphism of the Lie algebras, 𝔥 → 𝔤. This homomorphism is injective. Hence it is an isomorphism of 𝔥 to a Lie subalgebra of 𝔤. This subalgebra consists of all A ∈ 𝔤 sich that exp(tA) ∈ H for all t ∈ ℝ and the exponential map for H arises as restriction of the exponential map of G.

A first very remarkable application of the concept of Lie algebras is:

2.1 Theorem. A closed subgroup of a Lie group is smooth, hence also a Lie group.

The rest of this section is dedicated to the proof. We begin with a slight generalization of II.3.7. The proof is the same.

2.2 Lemma. Let

𝔤 = V ⊕ W

the decomposition of the Lie algebra of a Lie group G as direct sum of two sub vector spaces. The map

𝔤 = V ⊕ W ⟶ G, (A, B) ⟼ exp(A) exp(B)

is a local diffeomorphism at 0.

From now on H is a closed subgroup of a Lie group G.

2.3 Lemma. Let A_{n} ∈ 𝔤 be a sequence which converges to A ∈ 𝔤. Let also t_{n} be a sequence of real numbers all different from 0 which converges to 0. Assume

exp(t_{n}A_{n}) ∈ H for all n ∈ ℕ.

Then

exp(tA) ∈ H for all t ∈ ℝ.

Proof. Let be t ∈ ℝ. It is easy to construct a sequence of integers m_{n} such that m_{n}t_{n} → t. But then exp(m_{n}t_{n}A_{n}) = exp(t_{n}A_{n})^{m_{n}} is contained in H. The limit

exp(tA) is also contained in H because H is closed.  ⊔ ⊓

2.4 Lemma. The set

𝔥 := { A ∈ 𝔤; exp(tA) ∈ H for all t ∈ ℝ }

is a Lie subalgebra of 𝔤.

Proof. Let A, B ∈ 𝔥. From II.4.3 a) we know

exp(tA) exp(tB) = exp(t(A + B + R(t)))

with a remainder R(t) which converges to 0 as t goes to 0. Consider the sequence C_{n} = A + B + R(1/n). This sequence converges to C_{n} and exp(C_{n}/n) ∈ H. From 2.3 we see that all t(A + B) are contained in 𝔥. Thus 𝔥 is a vector space. A similar argument shows that [A, B] is contained in 𝔥. ⊔ ⊓ Proof of 2.1. We have to show that the closed subgroup H is smooth. It is sufficient to prove that H is smooth at e because e can be transported to any point by a translation, which is an analytic diffeomorphism. We know that exp induces an analytic diffeomorphism of an open neighbourhood 0 ∈ U ⊂ 𝔤 onto an open neighbourhood e ∈ V ⊂ G. The set exp(𝔥 ∩ U) is smooth by definition of smoothness. Hence it is sufficient to show

exp(𝔥 ∩ U) = H ∩ V

for sufficiently small U By definition of 𝔥 the left hand sind is contained in the right hand side. Our proof will be indirect, hence we assume that the converse is false for all U. This means that we find a sequence h_{n} ∈ H,of which tends to e but which is not of the form h_{n} = exp(H_{n}) with a sequence H_{n} ∈ 𝔥, which tends to 0. We will derive a contradiction. For this we choose a subvector space V ⊂ 𝔤 such that 𝔤 = 𝔥 ⊕ V. From 2.2 we obtain the existence of sequences H_{n} ∈ 𝔥 and A_{n} ∈ V which tend to 0 and such that exp(H_{n}) exp(A_{n}) = h_{n}. The elements A_{n} can be assumed to be different from 0. We choose any norm || · || on V, for example the maximum norm with respect to a basis of V. We can assume that A_{n}/||A_{n}|| is convergent and the limit B is different from 0. This limit is contained in V. From Lemma 2.3 we can conclude exp(ℝB) ⊂ H, hence B ∈ 𝔥. from 𝔥 ∩ V = 0 we obtain

B = 0 which gives the desired contradiction.  ⊔ ⊓

3. Continuous homomorphisms of Lie group

We begin with a quite trivial remark.

3.1 Remark. Let G → H be an analytic homomorphism of lie groups which is open (the images of open sets are open). Then the induced homomorphism of Lie algebras is surjective.

Let f : G → H be a continuous homomorphism of Lie groups. Assume that there exists a homomorphism of Lie algebras φ : 𝔥 → 𝔤 of the corresponding Lie algebras such that the diagram

\begin{matrix}G & ⟶^{f} & H \\ ↑ & & ↑ \\ 𝔤 & ⟶^{φ} & 𝔥 \end{matrix}

§4. The theorem of Poincaré Birkhoff Witt 33

commutes. Then f is analytic in a small open neighbourhood of the unit element. Using translations it follows that f is analytic everywhere.

We want to proof the same result without knowledge of the existence of φ. Hence we have to construct φ for an arbitrary continuous homomorphism f. The graph

G̃ = { (g, f(g)), g ∈ G }

is a closed subgroup. We know (2.1) that G̃ is smooth. We need the Lie algebra of G × H. The direct product 𝔤 × 𝔥 of two Lie algebras is defined in an obvious way. For example the Lie bracket is given by

[(A, A′), (B, B′)] := ([A, A′], [B, B′]).

The natural projections of G × H to G and H induce an isomorphism of the Lie algebra of G × H to 𝔤 × 𝔥. We identify the Lie algebra of G × H with 𝔤 × 𝔥. The exponential of G×H then simply is exp(A, B) = (exp(A), exp(B)). The Lie algebra 𝔤̃ of G̃ consists of all pairs (A, B) ∈ 𝔤 × 𝔥 such that exp(tB) = f(exp(tA) for all t in ℝ. Hence B is determined by A and we write B = A′. The first projection 𝔤̃ → 𝔤 induces an injective homomorphism. The natural homomorphism G̃ → G is a topological map, hence open. Therefore 𝔤̃ → 𝔤 is also surjective (3.1). Hence it is an isomorphism. We consider the converse map 𝔤 → 𝔤̃. We compose this map with the projection to the second factor to obtain a homomorphism α : 𝔤 → 𝔥. It is clear that the diagram

\begin{matrix}G & ⟶^{f} & H \\ ↑ & & ↑ \\ 𝔤 & ⟶^{φ} & 𝔥 \end{matrix}

commutes. We obtain:

3.2 Theorem. Any continuous homomorphism of Lie groups is analytic.

Assume that G is a topological group which admits two structures O_{1} and O_{2} as Lie groups. The identity map id_{G} is a topological map. By 3.2 it is an analytic diffeomorphism which means nothing else but O_{1} = O_{2}. This shows:

3.3 Theorem. When a topological group admits a structure as a Lie group this structure is unique.

Now a natural question arises which topological group admit a structure as a Lie group. This question is one of the famous Hilbert problems. It has been solved by:

Every topological group which is a topological manifold admits a structure as a Lie group.

4. The theorem of Poincaré Birkhoff Witt

We introduced the notion of a Lie algebra over an arbitrary field k. An associative algebra over k is an associative ring A with unit element 1_{A} together with a ring homomorphism ι : k → A with the property 1_{k} ↦ 1_{A}. Then A can be considered as a k vector space by means of αa := ι(α)a.

Any associative algebra A over k can be considered as Lie algebra Ã over k by means of [A, B] = AB − BA. A homomorphism of associative algebras A → B is also a homomorphism of the corresponding Lie algebras. Let 𝔤 be a Lie algebra. By definition a Lie homomorphism f : 𝔤 → A is a homomorphism of Lie algebras 𝔤 → Ã.

4.1 Definition. let 𝔤 be a Lie algebra. A universal enveloping algebra is an associative algebra U together with a Lie homomorphism

ι : 𝔤 ⟶ U,

such that the following universal property holds:

If α : 𝔤 → A is a Lie homomorphism of 𝔤 into any associative algebra A then there exists a unique homomorphism f : U ⟶ A of associative algebras such that ι ∘ f = α.

A universal enveloping algebra (U, ι) is uniquely determined in the following sense: Assume that (U′, ι′) is a second one then there exists a unique isomorphism of associative algebras σ : U → U′ such that σ ∘ ι = ι′. For this reason we will talk about the universal enveloping algebra, as soon as its existence has been proved.

Proof of existence of the universal enveloping algebra

We need some multilinear algebra:

Let V be a vector space. A basis of V (e_{ν})_{ν∈I} is a family of elements such that every element v ∈ V has a unique representation of the form

v = ∑_{i∈I} t_{i}e_{i},

where all but finitely many t_{i} ∈ k are different from 0 (which implies that the sum is well-defined). We use the well-known corollary from Zorn’s lemma that every vector space admits a basis.

Let n be a natural number and let V_{1}, … V_{n} be vector spaces. We consider a pair (V, τ_{n}) consisting of vector space V and a multilinear form

τ : V_{1} × ⋯ × V_{n} ⟶ V,

such that the following property holds:

Let (e_{ν}^{(i)})_{ν∈I_{i}} be a basis of V_{i} (1 ≤ i ≤ n). The family of elements

τ(e_{ν_{1}}^{(1)}, …, e_{ν_{n}}^{(n)})

parametrized by I_{1} × … × I_{n} is a basis of V. It is clear that (V, τ) will satisfy a universal property: Any multilinear map V_{1} × ⋯ × → W into any vector space W factors through a unique linear map V → W.

We indicate the existence proof. We fix a basis e_{ν}^{(n)} of V. One then considers symbols e_{i_{1}}^{(1)} ⊗ … ⊗ e_{i_{n}}^{(n)} and a vector space V, which has these symbols as basis. Then one defines τ by the condition

τ(e_{i_{1}}, …, e_{i_{n}}) = e_{i_{1}}^{(1)} ⊗ … ⊗ e_{i_{n}}^{(n)}.

After that one proves the universal property for (V, τ) and finally the defining property also for arbitrary basises f_{ν}^{(i)} instead of e_{ν}^{(i)}. We will use the notation

τ(x_{1}, …, x_{n}) = x_{1} ⊗ ⋯ ⊗ x_{n}.

The space V together with the multilinear map τ is called the tensor product of V_{1}, …, V_{n}. We write

V = V_{1} ⊗ ⋯ ⊗ V_{n}.

We apply this in the special case V_{1} = … = V_{n}. We write

V (n) = V ⊗ ⋯ ⊗ V n factors

in this case. The map τ_{1} : V → V (1) is an isomorphism. We also define V (0) = k. Next we consider the direct sum

T (V) := ⨁_{n=0}^{∞} V (n).

By definition this consists of all sequences

(v(0), v(1), …), v(n) ∈ V (n),

such that all but finitely many terms are different from 0. It is clear that T (V) is a vector space. We have natural embeddings

V (n) ⟶ T (V).

An element v ∈ V (n) is mapped to the sequence with v as n-the entry and all other entries 0.

We describe a natural bilinear map

τ_{n,m} : V (n) × V (m) ⟶ V (n + m).

If n = 0 or m = 0 this is the usual multiplication with scalars. In general it is defined by the prescription

(e_{i_{1}} ⊗ … ⊗ e_{i_{n}}, e_{j_{1}} ⊗ … ⊗ e_{j_{m}}) ⟼ e_{i_{1}} ⊗ … ⊗ e_{i_{n}} ⊗ e_{j_{1}} ⊗ … ⊗ e_{j_{m}}.

It is clear how to extend this to a bilinear map

τ : T (V) × T (V) ⟶ T (V).

One defines τ((v(n)), (v′(n))) = ((w(n)) where

w(n) = ∑_{ν+μ=n} τ_{νμ}(v(ν), v′(μ)).

and this is a ring structure on T (V). The algebra structure comes from the additional map k ≅ V (0) ↪ T (V).

The algebra T (V) is called the tensor algebra of V. On should consider it together with the natural linear map V ≅ V (1) ↪ T (V). It has an obvious universal property. Let V → A be a linear map to an associative algebra, This map extends to a unique homomorphism T (V) → A of associative algebras.

This universal property gives the hint how the universal enveloping algebra of a Lie algebra 𝔤 can be constructed. On starts with the tensor algebra T (𝔤). But the natural map 𝔤 → T (𝔤) needs not to be a Lie homomorphism. The idea now is to consider a factor algebra of T (V). Recall that that an ideal of an associative ring R is an additive subgroup A ⊂ R with the property that ra ∈ A and ar ∈ A for a ∈ A and r ∈ R. Let R/A denote the factor group. It is easy to show that this carries a unique ring structure such that the natural projection R → R/A is a ring homomorphism. If in addition R is an algebra and A a sub vector space then R/A will be an algebra. Back to our situation ι : 𝔤 → T (𝔤). We consider all ideals of T (𝔤) which contain all elements of the form ι(A)ι(B) − ι(B)ι(A). There exist ideals of this kind for example T (V) is such an ideal. The intersection of an arbitrary system of ideals is also an ideal. We see that there is a smallest ideal with this property. We define U(𝔤) to be the quotient of T (V) by this ideal. Composing the map 𝔤 → T (𝔤) with the natural projection T (𝔤) → U(𝔤) we get a linear map

𝔤 ⟶ U(𝔤)

which by construction is a homomorphism of algebras. It is easy to verify the universal property.

We want to prove that 𝔤 → U(𝔤) is injective. This is always true but we will prove it only when the dimension of V is finite or countable. In this case the index set I of a basis can to be taken as {1, …, n} for some integer n ≥ 0 or as the set of all natural numbers ℕ. In any case we have an ordering of the elements of I and we write

I = { 1, 2 … }

keeping it open whether this “sequence” is finite or not. We consider now ordered tuples

M = (i_{1} ≤ i_{2} … ≤ i_{m}), i_{ν} ∈ I,

of arbitrary length m > 0 and denote by e(M) the image of e_{i_{1}} ⊗ ⋯ ⊗ e_{i_{m}} in U(𝔤). We admit also the empty tuple M and define e_{M} = 1 ∈ k in this case. The famous Poincaré-Birkhoff-Witt theorem states:

4.2 Theorem. Let e_{1}, e_{2}, … be a vector space basis of a Lie algebra 𝔤. The elements

e(M) ∈ U(𝔤), M = (i_{1}, …, i_{m}), i_{1} ≤ i_{2} ≤ … ≤ i_{m}, 0 ≤ m,

(images of e_{i_{1}} ⊗ ⋯ ⊗ e_{i_{m}}) build a basis of the vector space U(𝔤). Corollary. The natural map 𝔤 → U(𝔤) is injective.

Proof. In a first step we prove that the ordered products generate U(𝔤) as vector space. At the moment we only know that the products e_{i_{1}} ⊗ ⋯ ⊗ e_{i_{m}} without the ordering condition generate U(𝔤). But we have a relation [e_{i}, e_{j}] = ∑_{ν} C_{ν}e_{ν} with constants C_{ν}. Assume that a product is given with possibly not ordered (i_{1}, …, i_{m}). Assume for example i_{1} > i_{2}. The we replace e_{i_{1}} ⊗ e_{i_{2}} by the sum of e_{i_{2}} ⊗ e_{i_{1}} and a linear combination of the e_{i}. This means that we replace the whole product by e_{i_{2}} ⊗ e_{i_{1}} ⊗ e_{i_{3}} ⋯ ⊗ e_{i_{m}} and a sum of products of less then m terms. Thus we can prove the claim by induction on m.

We have to prove the linear independence of the ordered products e(M). This is a little more involved. The idea is to construct a vector space S and a Lie homomorphism α : 𝔤 → End(S). such that the ordered products e′(M) = α(e_{i_{1}}) ⋯ α(e_{i_{m}}) are linearly independent. From the universal property we obtain a homomorphism U(𝔤) → End(S) such that e(M) is mapped to e′(M). Then the e(M) must be independent.

We will take for S the ring of polynomials k[X_{1}, X_{2}, …] where the variables X_{i} are indexed by I. A basis of S is given by the monomials

X^{λ} := ∏_{j∈I} X_{j}^{λ_{j}}, λ = (λ_{1}, λ_{2}, …),

where only a finite number of the λ_{j} is allowed to be different from 0. We define the degree of a monomial by deg(X^{λ}) = λ_{1} + ⋯ + λ_{n} and of an arbitrary non-zero element as the maximum of all degrees of occurring monomials. The multiplication law in S is defined in an obvious way, the essential property is that the X_{i} commute. A formal way to define such a polynomial ring is to take the factor algebra of the tensor algebra T (V) by the ideal which is generated by e_{i}e_{j} − e_{j}e_{i}. The “indeterminates” X_{i} then are the images of the e_{i}.

Again we consider ordered tuples

M = (i_{1} ≤ i_{2} … ≤ i_{m}), i_{ν} ∈ I

and we define

X(M) := X_{i_{1}} ⋯ X_{i_{m}}.

Theses elements are nothing else but the monomials X^{λ}. For example X_{1}^{2}X_{3} = X(1, 1, 3). Hence the elements X(M) build a basis of S.

We now want to define a linear map α : 𝔤 → End(S). This means that we have to define for any basis element e_{i} an endomorphism of vector spaces α(e_{i}) : S → S. We simply will write

AP := (α(A))(P) for A ∈ 𝔤, P ∈ S.

To get a linear map α the e_{i}X(M) could be described quite arbitrarily. But we want to have that α is a Lie homomorphism and this requires conditions. We demand the following conditions for α.

A) e_{i}X(M) = X_{i} X(M) for i ≤ i_{1} and e_{i}1 = X_{i}.

B) deg(e_{i}X(M)) ≤ deg X(M) + 1,

deg(e_{i}X(M)) − X_{i} X(M) ≤ deg(X(M)). C) e_{i}(e_{j}X(N)) = e_{j}(e_{i}X(N)) + [e_{i}, e_{j}]X(N).

Assume for a moment that α exists. From C) follows that α is a Lie homomorphism. From A) follows e(M)1 = X(M). This gives the proof of 4.2. Proof of existence of α. The following proof will not only show the existence but also the uniqueness of α. We assume for a moment that α has been constructed. The problem is to get hold of e_{i}X(M) for i > i_{1}. We will write

i > M instead of i > i_{1}.

We have X(M) = X_{i_{1}} X(N) where N = {i_{2}, …, i_{m}}. Using C) we obtain

e_{i}X(M) = e_{i}(X_{i_{1}} X(N)) =^{A} e_{i}(e_{i_{1}} X(N)) =^{C} [e_{i}, e_{i_{1}}]X(N) + e_{i_{1}}(e_{i} X(N))

= [e_{i}, e_{i_{1}}]X(N) + e_{i_{1}}(e_{i} X(N) − X_{i} X(N)) + e_{i_{1}}X_{i} X(N)

=^{A} [e_{i}, e_{i_{1}}]X(N) + e_{i_{1}}(e_{i} X(N) − X_{i} X(N)) + X_{i_{1}}X_{i} X(N)

= [e_{i}, e_{i_{1}}]X(N) + e_{i_{1}}(e_{i} X(N) − X_{i} X(N)) + X_{i} X(M) .

The point now is that the degree of X(N) and also of e_{i} X(N) − X_{i} X(N) (because of condition B) is smaller than the degree of X(M). This gives us the possibility for an inductive definition. To be precise we will prove the following statement by induction on n. Let S(n) the subspace of elements of degree ≤ n (including 0). There exists a (unique) bilinear map

α : 𝔤 × S_{n} ⟶ S_{n+1}, (A, P) ⟼ AP,

such that A, B are satisfied for deg X(M) ≤ n and such that C is satisfied for deg(X(N)) < n. Assume that α exists for a certain n. We then define

(∗)  e_{i}X(M) := [e_{i}, e_{i_{1}}]X(N) + e_{i_{1}}(e_{i} X(N) − X_{i} X(N)) + X_{i} X(M).

(The first two terms on the right hand side are defined by induction.) We have to verify the conditions A, B, C for deg X(M) ≤ n + 1, deg X(N) ≤ n. The conditions A, B are clear. It remains to verify C. We first notice that the definition (*) implies

e_{i}X(M) := [e_{i}, e_{i_{1}}]X(N) + e_{i_{1}}(e_{i}X(N))

But X(M) = X_{i_{1}}X_{N} = e_{i_{1}}X_{N} because of A. This implies condition C for j = i_{1}.

e_{i}(e_{j}X(N)) = e_{j}(e_{i}X(N)) + [e_{i}, e_{j}]X(N) j = i_{1}.

The restriction j = i_{1} means j ≤ M. In a similar way this condition is verified for i ≤ M. It remains to prove

e_{i}(e_{j}X(N)) = e_{j}(e_{i}X(N)) + [e_{i}, e_{j}]X(N) for i > N and j > N.

This is somewhat tedious. We write

N = (ν, Q), hence ν < i, ν < j.

We have X(N) = X_{ν} X(Q) = e_{ν}X(Q) and by induction

e_{j}X(N) = e_{j}(e_{ν}X(Q)) = e_{ν}(e_{j}X(Q)) + [e_{j}, e_{ν}]X(Q).

We want to get a new expression for e_{i}(e_{j}X(N)). By induction we get

e_{i}([e_{j}, e_{ν}]X(Q)) = [e_{j}, e_{ν}](e_{i}X(Q))) + [e_{i}, [e_{j}, e_{ν}]]X(Q).

Similarly we claim

e_{i}(e_{ν}(e_{j}X(Q))) = e_{ν}(e_{i}(e_{j}X(Q))) + [e_{i}, e_{ν}](e_{j}X(Q)).

The induction assumption cannot directly be applied, but we can argue as follows: We rewrite e_{j}X(Q) = X_{j}X(Q) + w, where deg(w) ≤ deg(X(Q)). By induction assumption it is sufficient to show

e_{i}(e_{ν}(X_{j}X(Q))) = e_{ν}(e_{i}(X_{j}X(Q))) + [e_{i}, e_{ν}](X_{j}X(Q)).

But this case we treated in advance! The last two formulae give

\begin{matrix}e_{i}(e_{j}X(N)) = \\ e_{ν}(e_{i}(e_{j}X(Q))) + [e_{i}, e_{ν}](e_{j}X(Q)) + [e_{j}, e_{ν}](e_{i}X(Q))) + [e_{i}, [e_{j}, e_{ν}]]X(Q). \end{matrix}

The same formula can be applied for (j, i) instead of (i, j). Taking the difference several terms cancel and we get

\begin{matrix}e_{i}(e_{j}X(N)) − e_{j}(e_{i}X(N)) = e_{ν}(e_{i}(e_{j}X(Q))) − e_{ν}(e_{j}(e_{i}X(Q))) \\ + [e_{i}, [e_{j}, e_{ν}]]X(Q) − [e_{j}, [e_{i}, e_{ν}]]X(Q). \end{matrix}

The induction assumption can be applied twice to obtain

e_{i}(e_{j}X(N)) − e_{j}(e_{i}X(N)) =

e_{ν}([e_{i}, e_{j}]X(Q)) + [e_{i}, [e_{j}, e_{ν}]]X(Q) − [e_{j}, [e_{i}, e_{ν}]]X(Q) =

[e_{i}, e_{j}](e_{ν}X(Q))) + [e_{ν}, [e_{i}, e_{j}]]X(Q)) + [e_{i}, [e_{j}, e_{ν}]]X(Q) + [e_{j}, [e_{i}, e_{ν}]]X(Q).

Now the Jacobi identity has to by applied!

The result is as desired:

e_{i}(e_{j}X(N)) − e_{j}(e_{i}X(N)) = [e_{i}, e_{j}](e_{ν}X(Q))) = [e_{i}, e_{j}]X(N).

This completes the proof of 4.2.

We finally mention that Poincaré-Birkhoff-Witt theorem is true also if the dimension is not finite or countable. One has to use a well-ordered basis of 𝔤.

The Poincaré-Birkhoff-Witt theorem sometimes is expressed in a different way. We assume for simplicity that 𝔤 is finite dimensional with basis e_{1}, …, e_{n} and that the characteristic of the ground field is different from 0. We consider an arbitrary linear combination X(t) := t_{1}e_{1} + ⋯ + t_{n}e_{n} with variable coefficients t = (t_{1}, …, t_{m}). Then we consider the power X^{m} in U(𝔤) for some integer m ≥ 0. This can be written in the form

X(t)^{m} = ∑_{λ} t^{λ}e^{[λ]}, t^{λ} = t_{1}^{λ_{1}} ⋯ t_{n}^{λ_{n}}

with certain e^{[λ]} ∈ U(𝔤). For example in the case n = 2 we have

e^{[1,1]} = e_{1}e_{2} + e_{2}e_{1} = 2e_{1}e_{2} + [e_{1}, e_{2}].

We see in this special example and then in general, that the basis elements e(M) can be expressed as linear combinations of the e^{[λ]}. This the Poincaré-Birkhoff-Witt theorem states:

4.3 Remark. Let the base field be of characteristic 0 and let 𝔤 be of finite dimension 0. The elements e^{[λ]} defined by

(t_{1}e_{1} + ⋯ + t_{n}e_{n})^{m} = ∑_{λ} t^{λ}e^{[λ]}, λ = (λ_{1}, …, λ_{n}), λ_{i} ≥ 0,

build a vector space basis of U(𝔤).

In this connection we mention:

4.4 Remark. Let 𝔤 be a Lie algebra. The universal envelopping algebra is generated as vector space by powers C^{k} of elements C ∈ 𝔤.

. This is clear because the e^{[λ]} can be reconstructed from the powers of the t_{1}e_{1} +

⋯ + t_{n}e_{n}.  ⊔ ⊓

§5. Lie sub algebras and local subgroups 41

5. Lie sub algebras and local subgroups

The difficult part of the Poincaré-Birkhoff-Witt theorem was the proof of the linear independence of the ordered products. When 𝔤 is the Lie algebra of a Lie group G the proof is far easier:

We recall that the elements A of 𝔤 can be considered as operators Ã on G and that the set of all operators is an algebra. Hence we can consider the algebra which is generated by the Ã.

5.1 Definition. The algebra D(G) is the algebra of all operators on a Lie group G which is generated by the left invariant operators.

It can be shown that D(G) is the algebra of all linear and left invariant differential operators on G, but we will not need this.

The natural map 𝔤 → D(G) is a Lie homomorphism. This follows from the definition of the Lie bracket. It extends to a homomorphism of algebras U(𝔤) → D(G), which is surjective.

5.2 Proposition. The natural homomorphism

U(𝔤) ⟶ D(G)

is an isomorphism.

Corollary. Let G, H be two Lie groups with Lie algebras 𝔤, 𝔥. Let 𝔤 → 𝔥 be a homomorphism of Lie algebras. Then there exists a unique homomorphism of associative algebras D(G) → D(H), such that the diagram

D(G) ⟶ D(H)
↑  ↑
𝔤 ⟶  𝔥

commutes. (The vertical arrows are natural inclusions.)

Proof. Let e_{1}, …, e_{n} be a basis of 𝔤 and E_{1}, …, E_{n} the corresponding elements in D(G). It is sufficient to show (4.3) that the

E^{[λ]} : (t_{1}E_{1} + ⋯ + t_{n}E_{n})^{m} = ∑ t^{λ}E^{[λ]}

are linearly independent. From the corollary of II.4.1 we obtain

f(exp(t_{1}E_{1} + ⋯ + t_{n}E_{n})) = ∑_{λ} t^{λ}(E^{[λ]}f)(e).

Hence the (E^{[λ]}f)(e) are Taylor coefficients of the function (t_{1}, …, t_{n}) ↦ f(exp(t_{1}E_{1}+ ⋯ + t_{n}E_{n})). We can take the functions f(exp(t_{1}E_{1} + ⋯ + t_{n}E_{n})) = t^{λ} to prove the linear independence of the E^{[λ]}. ⊔ ⊓

As we mentioned, this proof implies the linear independence of the ordered products of the e_{1}, …, e_{n} in U(𝔤), which was the difficult part of the Poincaré-Birkhoff-Witt theorem.

An important application of 5.2 is:

5.3 Proposition. Let G, H be two Lie groups with Lie algebras 𝔤, 𝔥 and σ : 𝔤 → 𝔥 a homomorphism of Lie algebras. Denote by F_{G}(X, Y), F_{H}(X, Y) the functions log(exp(X) exp(Y)) in both cases. (They are analytic functions in small open neighbourhoods of the origin in 𝔤 × 𝔤 resp. 𝔥 × 𝔥 with values in 𝔤 resp. 𝔥.) We have

F_{H}(σ(X), σ(Y)) = σ(F_{G}(X, Y)),

where X, Y varies in some small neighbourhood of 0 in 𝔤.

This result is of course trivial when σ is induced by a homomorphism of Lie groups G → H. The point is that it is true without this assumption.

Proof of 5.3. It is sufficient to show

L(F_{H}(σ(X), σ(Y))) = L(σ(F_{G}(X, Y)))

for an arbitrary linear function L : 𝔥 → ℝ. Because of II.4.2 we have to show that for A, B ∈ 𝔤

σ̃(A)^{n}σ̃(B)^{m}(L ∘ log_{H}))(e) = Ã^{n}B̃^{m}(L ∘ σ ∘ log_{G})(e).

Now we use the homomorphism σ̃ : D(G) → D(H) from the corollary of 5.2. Using this homomorphism the claim can be written as

σ̃(Ã^{n}B̃^{m})(L ∘ log_{H}))(e) = Ã^{n}B̃^{m}(L ∘ σ ∘ log_{G})(e).

This formula is true for all D ∈ C^{ω}(G) instead of Ã^{n}B̃^{m}. To prove this we use the fact that D(G) is generated as vector space by powers C̃^{k} of elements C ∈ 𝔤. Thus the claim reduces to

σ̃(C̃^{k})(L ∘ log_{H}))(e) = C̃^{k}(L ∘ σ ∘ log_{G})(e).

This formula follows from:

5.4 Remark. Let G be a Lie group with Lie algebra 𝔤 and L : 𝔤 → ℝ a linear form. Then

(C̃^{k}(L ∘ log))(e) = \{ \begin{matrix}L(C) & for n = 1, \\ 0 & for k > 1 \end{matrix} .

This and hence 5.3 follows from II.4.1.  ⊔ ⊓

This result states that the multiplication law of a Lie group in a small neighbourhood e ∈ G only depends on the structure of the Lie algebra. Later we will derive an explicit formula for the local multiplication law F in terms of the Lie algebra (Campbell-Hausdorff formula).

§6. Local subgroups 43

6. Local subgroups

6.1 Definition. Let G be a Lie group. A local subgroup of G is a subset U ⊂ G with the following property:

1. U contains e ∈ G.

2. U is smooth in G.

3. There exists an open neighbourhood e ∈ U′ ⊂ U (open in U) such that

(a, b) ∈ U′ × U′ ⟹ ab ∈ U.

We call dim_{e}(U) the dimension of the local subgroup.

Another important application of 5.2 is

6.2 Proposition. Let G be a Lie group with Lie algebra 𝔤. let 𝔥 ⊂ 𝔤 be a Lie sub algebra, i.e. a sub vector space which is closed under the Lie bracket. For a sufficiently small open neighbourhood e ∈ W ⊂ 𝔥 the subset exp(W) is a local subgroup of dimension dim_{ℝ}(𝔥).

Proof. For sufficiently small A, B the element exp(A) exp(B) is of the form exp(C). We have to show C ∈ ℌ if A, B ∈ 𝔥. Actually we will show that for every linear form L : 𝔤 → ℝ which vanishes on 𝔥 also L(C) = 0. This means

L(A) = L(B) = 0 ⟶ L(log(exp(A) exp(B))) = 0.

Using the Taylor expansion this follows from D(L ∘ log)(e) = 0 for D = Ã^{n}B̃^{m}. From the Birkhoff-Witt theorem follows that U(𝔥) is embedded as subalgebra in U(𝔤) which can be identified with D(G). Thus D can be considered as element of U(𝔤). Again from the Birkhoff-Witt theorem follows that U(𝔤) is generated by as vector space by powers C^{k}, C ∈ 𝔥. Recall that C^{k}(L ∘ log)(e) = 0 for k > 1 and L(C) for k = 1. But L(C) is 0 by choice of L. This proves 6.2. ⊔ ⊓

We want to identify two local subgroups U_{1}, U_{2}, if there is a small open neighbourhood of e in G such the intersections of U_{1} and U_{2} with this neighbourhood agree. In this connection we should mention that an open neighbourhood e ∈ U_{0} ⊂ U of a local group is also a local group. This follows from the continuity of the multiplication U′ ×U′ → U. To be more precisely we call two local subgroups with this property equivalent.

6.3 Definition. A sub group germ of a Lie group is a full equivalence class of local sub groups.

We have seen that a Lie sub algebra gives rise to a subgroup germ. Conversely let U ⊂ G be a local subgroup. We define 𝔥 to be the set of all A ∈ 𝔤 such that exp(tA) ∈ U for all t in some small neighbourhood (−ε, ε) of 0. The proof as of 2.4 shows thet 𝔥 is a Lie algebra and of course this Lie algebra depends only on the germ of U. This shows:

6.4 Theorem. There is a one-to-one correspondence between subgroup germs of Lie group G and sub Lie algebras of 𝔤.

There exists also the notion of a local homomorphism:

6.5 Definition. Let G, H be a Lie group. A local homomorphism of G into H is a pair (U, f) consisting of an open neighbourhood of e ∈ U ⊂ G and an analytic map f : U → H such that f(ab) = f(a)f(b) for all a, b in a sufficiently small neighbourhood of e.

A local homomorphism induces a tangent map 𝔤 → 𝔥. The same proof as of II.3.8 shows that this is a homomorphism of Lie algebras.

We call two local homomorphisms from G into H equivalent, if they agree in a small neighbourhood of e.

6.6 Definition. A germ of a homomorphism of G into H is a full equivalence class of local homomorphisms.

Clearly the induced homomorphism 𝔤 → 𝔥 of a local homomorphism only depends on the gern of the homomorphism.

6.7 Theorem. Let G, H be Lie groups. There is a natural bijection between homomorphisms 𝔤 → 𝔥 of the corresponding Lie albgebras and the germs of homomorphisms from G into H.

Proof. It remains to show that for a homomorphism φ : 𝔤 → 𝔥 the map g ↦ exp_{H}(φ(log_{G}(g) defines a local homomorphism. But this follows from 5.3.

7. Abstract Lie algebras as Lie algebras of Lie groups

In this section we want to prove that every subalgebra of 𝔤𝔩(n, ℝ) is isomorphic to the Lie algebra of a Lie group. Later we will see that every Lie algebra 𝔤 over ℝ of finite dimension is isomorphic to a sub algebra of 𝔤𝔩(n, ℝ). Then we obtain that every 𝔤 is isomorphic to the Lie algebra of a Lie group.

What we actually prove is:

7.1 Proposition. Let U ⊂ G be a local subgroup. Then there exists a connected Lie group H and an immersion of Lie groups ι : H → G, such that that the natural homomorphism 𝔥 → 𝔤 is an isomorphism of 𝔥 onto the Lie algebra of the germ of U. Corollary. Every subalgebra of 𝔤𝔩(n, ℝ) is isomorphic to the Lie algebra of a Lie group. Additional Remark. The pair (H, ι) is determined by the germ of U in the following sense. Assume that (H′, ι′) is a second pair. Then there exists a unique isomorphism σ : H → H′ of Lie groups with ι = ι′ ∘ σ.

§7. Abstract Lie algebras as Lie algebras of Lie groups 45

Proof . Let U ⊂ G be a local subgroup. We can assume that U is connected. We consider the abstract subgroup H ⊂ G generated by U. We define a topology of H which may be very different from the induced topology. A subset V ⊂ H is called open, if for every a ∈ V there exists a small neighbourhood e ⊂ W ⊂ U of the uni element in U ssuch that aW ⊂ V. It is easy to see that this system satisfies the axioms of a topology, morover U is open in H because of the assumption that U is a local subgroup. We define a geometric structure on H. A function f : V → ℝ is called analytic if in the situation above x ↦ f(a^{−1}x is analytic on W. recall that W is open in U; which is smooth in G, hence an analytic manifold. It is easy to see that H gets a Lie group and moreover that the canonical immersion H → G defines an immmersion. The rest of 7.1 is clear. ⊔ ⊓

We give some comments which shed some new light into the connection between Lie algebras and Lie groups. We use topological covering theory. A homomorphism of connected Lie groups G → H is called a covering, if the kernel is a discrete subgroup. If conversely Γ ⊂ G is a discrete an normal subgroup of G, then the quotient G/Γ carries a natural structure as Lie group such that G → G/Γ is a covering. Moreover if G → H is a covering of connected Lie groups with kernel Γ, then the canonical map G/Γ → H is an isomorphism of Lie groups. Let G be a connected Lie group and G̃ its universal covering. This is a simply connected Lie group together with a covering homomorphism π : G̃ → G. The universal covering exists and is unique in an obvious sense. Every homomorphism of connected Lie groups f : G → H lifts in a unique way to a homomorphism˜:G̃ → H̃ of the universal coverings. The covering homomorphism π : G̃ → G induces an isomorphism of the corresponding Lie algebras G̃ → G.

A relatively simple topological argument gives:

7.2 Proposition. Every germ of a homomorphism of a simply connected Lie group G into a Lie group H can be represented by a global homomorphism of Lie groups G → H.

We recall a Lie algebras realizable if it is isomorphic to a Lie algebra of a Lie group. We will see later that every finite dimensional Lie algebra over ℝ is realizable. From 7.2 it is easy to derive:

7.3 Theorem. The category of realizable Lie algbras is eqivalent to the category of connected and simply connected Lie groups.

This means:

1. Every realizable Lie algebra is isomorphic to the Lie algebra of a connected and simply connected Lie group and this Lie group is unique up to isomorphism.

2. Let G, H be connected and simply connected Lie groups. There is a natural bijection between the analytic homomorphisms G → H and the homomorphisms of Lie algebras.

Another way to express 1) is: Two connected Lie groups have isomorphic Lie algebras if any and only if their universal coverings are isomorphic as lie groups.

We give an interesting example: Let G be an abelian connected Lie group. It follows from the corollary of 5.2 that𝔤 → G is a homomorphism of the additive group of 𝔤 into G and more over that this is the universal covering of G. It is known that every discrete subgroup Γ ⊂ V of a finite dimensional real vector space is of the form

Γ = ℤe_{1} + ⋯ ℤe_{m}

with linearly independent vectors e_{1}, …, e_{m}. from this follows: Every connected abelian Lie group is isomorphic to the direct product of a ℝ^{n} and a torus S_{1}^{m}.

Chapter IV. Lie-Algebren

1. The Campbell-Hausdorff series

This section is not needed for the rest of the text. Nevertheless we include it because of its own interest.

We want to give an explicit formula for the function

F (X, Y) = log(exp(X) exp(Y),

which is defined in a small neighbourhood of (0, 0) ∈ 𝔤 × 𝔤, where 𝔤 is a Lie algebra of a Lie group G. We assume that 𝔤 is isomorphic to a subalgebra of 𝔤𝔩(n, ℝ). We will see later that this is no restriction (Ado’s theorem). We know that F_{G} is the restriction of the same function for GL(n, ℝ). Hence an explicit formula for 𝔤𝔩(n, ℝ) will imply an explicit formula for G.

This means that we can restrict to the case

G = GL(n, ℝ), 𝔤 = 𝔤𝔩(n, ℝ).

In this case it is easy to get a local inversion of the exponential function. For purposes of convergence we chose on M_{n}(ℝ) the norm

||A|| = \sqrt{∑ a_{ij}^{2}} = trace(A′A).

It has the propertes

||A + B|| ≤ ||A|| + ||B||, ||tA|| = |t|||A||, ||AB|| ≤ ||A|| ||B||

for A, B ∈ M_{n}(ℝ), t ∈ ℝ. Let E be the unit matrix. The series

log(E + A)A := ∑_{ν=1}^{∞}(−1)^{ν+1} \frac{A^{ν}}{ν}

converges for ||A|| < 1 and one inverts the exponential, which justifies the notion log for this series.

1.1 Proposition. We have

e^{log(E+A)} = E + A (||A|| < 1).

We indicate a proof. The formula can be formulated also for complex A with an obvious definition of the norm. It is sufficient to prove the formula for diagionalizable A by a density argument. Then it is easy to reduce the claim to diagonal matrices and then to the known case n = 1. ⊔ ⊓

A formal calculation shows for small X, Y

log(exp(X) exp(Y)) = ∑_{m=1}^{∞} ∑_{{p_{i}≥0, q_{j} ≥0 \atop p_{i}+q_{i}≥1}} \frac{(−1)^{m+1}}{mp_{1}!q_{1}! … p_{m}!q_{m}!}X^{p_{1}}Y^{q_{1}} ⋯ X^{p_{m}}Y^{q_{m}}.

A majorant for the series can be obtained by

\frac{1}{1 − (exp ||X|| exp ||Y || − 1).}

Thus all formal rearrangements are allowed. We can write this series in the from

F (X, Y) = log(exp(X) exp(Y)) = ∑_{p=1}^{∞} F_{p}(X, Y),

wherw F_{p} are polynomials, homogeneous of degree p, i.e.

F_{p}(tX, tY) = t^{p}F_{p}(X, Y).

We know already F_{1}(X, Y) = X + Y, F_{2}(X, Y) = [X, Y]/2. With some patience one can obtain more terms, for example

F_{3}(X, Y) = \frac{1}{12}([X, [X, Y]] + [Y, [Y, X]), F_{4}(X, Y) = \frac{1}{24}[X, [Y, [X, Y]]].

we observe that that the F_{p} have been expressed in terms of the structure of the Lie algebra 𝔤𝔩(n, ℝ)(n, ℝ). The point is that this is true for all p. This is the content of the Campbell Hausdorff formula: To formulate ist we introduce the multiple Lie bracket

ϕ(X_{1}, X_{2}, …, X_{m−1}, X_{m})) := [X_{1}, [X_{2}, X_{3} … [X_{m} − 1, X_{m}]] …].

We also use the notation

ϕ(X^{p_{1}}Y^{q_{1}} ⋯ X^{p_{m}}Y^{q_{m}}) := ϕ(\overbrace{X, …, X}^{p_{1}}, Y, …, Y, . . . . . . , X …, X, \overbrace{Y, … Y}^{q_{m}}).

This should be considered as a formal notation. A more careful notation is something like ϕ(p_{1}, q_{1} …, p_{m}, q_{m}, X, Y). We define polynomials G_{p}(X, Y) by the two properties

\begin{matrix}G_{p}(tX, tY) = t^{p}G_{p}(X, Y), \\ ∑_{p=1}^{∞} & G_{p}(X, Y) = ∑_{m=1}^{∞} ∑_{{p_{i}≥0, q_{j} ≥0 \atop p_{i}+q_{i}≥1}} \frac{(−1)^{m+1}}{mp_{1}!q_{1}! … p_{m}!q_{m}!}ϕ(X^{p_{1}}Y^{q_{1}} ⋯ X^{p_{m}}Y^{q_{m}}). \end{matrix}

§1. The Campbell-Hausdorff series 49

The Campbell-Hausdorff formula states 1.2 Theorem. The polynomial identity

F_{p}(X, Y) = G_{p}(X, Y)

holds for all p. Corollary Let 𝔤 the Lie algebra of a Lie group G. Assume that 𝔤 is embeddable (isomorphic to a Lie sub algebra of some 𝔤𝔩(n, ℝ)). Then the function F_{G}(X, Y) (s. III.5.3) admits the expansion

F_{G}(X, Y) = ∑_{p=1}^{∞} G_{p}(X, Y).

In the rest of this section we scetch a proof.

The ground field K can be an arbitrary field of characteristic 0. We introduce a very special Lie algebra, the so-called free Lie algebra of two generators. For this we consider the tensor algebra T = T (V) over the two dimensional vector space V = K^{2}. We denote the standard basis of K^{2} by

X = (1, 0), Y = (0, −1).

We recall that T is the direct sum of the spaces

T (n) = V^{⊗n} := V ⊗ ⋯ ⊗ V (T (0) = K).

The elements of T (n), n > 0, are linear combinations of monomials

X^{p_{1}}Y^{q_{1}}X^{p_{2}}Y^{q_{2}} ⋯ X^{p_{m}}Y^{q_{m}}, n = p_{1} + q_{1} + ⋯ p_{m} + q_{m}.

We call n = p_{1} + q_{1} + ⋯ p_{m}q_{m} also the degree of the monomial. We only assume p_{i}, q_{i} ≥=, hence trivial factors in the monomial can occur. A monomial is called reduced, if it is X^{n} or Y^{n} or if all p_{i} and q_{i} are positive. Clearly the reduced monomials build a basis of T (n). We call T also the polynomial ring over K im two non-commuting variables. We also consider the Lie bracket [x, y] = xy − yx which makes T to a Lie algebra. We also consider the Lie subalgebra L of T which is generated by X and Y. As vector space L is generated by all elements of the form

[a_{1}, [a_{2}, [… [a_{r−1}, a_{r}]] …], a_{i} ∈ {X, Y }.

The canonical embedding

L ⟶ T

is a Lie homomorphism. It is easy to see that it has the universal property. Hence T is the universal enveloping algebra of L. It is clear that the dimension of L is finite or countable. Let u_{1}, u_{2}, … be a basis of the vector space L. We will use the Birkhoff-Witt theorem which states that the ordered monomials

u_{1}^{ν_{1}}u_{2}^{ν_{2}} ⋯

(all but a finite number of the exponents ν_{j} is assumed to be zero) are linearly independent and moreover they define a basis of T as vector space.

The free Lie algebra L has a remarkable property 1.3 Remark. Let A, B be two elements of a Lie algebra K. Then there exist a unique homomorphism of Lie algebras

L ⟶ 𝔤, X ⟼ A, Y ⟼ B.

Proof . The linear map

XK + Y K ⟶ 𝔤, X ⟼ A, Y ⟼ B

extends to a homomorphism of associative algebras

T ⟶ U(𝔤)

by the universal property of the tensor algebra. Its restriction to L is a homorphism from L to 𝔤. ⊔ ⊓

We will prove a polynomial identity F_{p} = G_{p} in T. If 𝔤 → 𝔤𝔩(n, ℝ) is a homomorphism of Lie algebras this formula will specialize to 1.2. It will be easier to write the identity as identity of formal series ∑ F_{p} = ∑ G_{p}. Hence one has to introduce a bigger algebra T̂ ⊃ T.

Recall that the tensor algebra T (V) of a vector space V is the direct sum of all V^{⊗m}. One can consider also the direct product

T̂(V) = ∏_{n=0}^{∞} V^{⊗}.

We write its elements as formal sums

∑ v_{n}, v_{n} ∈ V^{⊗},

which now can contain infinitely many terms different from 0. The point is that in the formal product of two such expressions only finite occur! Hence T̂(V) is a algebra which contains T (V) as subalgebra.

We use this for V = K^{2}. We should consider T̂ as algebra of formal power series in two non-commuting variables X, Y. The advantage of this construction that we can define exp(X) and exp(Y) als elements of T̂ in an obvious way. This we can define also exp(X) exp(Y) ∈ T̂. There is another obvious definition, which we leave to the reader. If H ∈ T̂ is an element with a non-zero constant coefficient ∈ K = V^{⊗0} then log H ∈ T̂ can be defined in a formal way. Especially

F (X, Y) := log(exp(X) exp(Y)

is a well defined element of T̂. And also the series

G = ∑_{p=1}^{∞} G_{p}(X, Y) = ∑_{m=1}^{∞} ∑_{{p_{i}≥0, q_{j} ≥0 \atop p_{i}+q_{i}≥1}} \frac{(−1)^{m+1}}{mp_{1}!q_{1}! … p_{m}!q_{m}!}ϕ(X^{p_{1}}Y^{q_{1}} ⋯ X^{p_{m}}Y^{q_{m}}).

an be considered as element of T̂. The most general version of the Hausdorff Campbell formula states:

1.4 Theorem. The formula

log(exp(X) exp(Y)) = ∑_{m=1}^{∞} ∑_{{p_{i}≥0, q_{j} ≥0 \atop p_{i}+q_{i}≥1}} \frac{(−1)^{m+1}}{mp_{1}!q_{1}! … p_{m}!q_{m}!}ϕ(X^{p_{1}}Y^{q_{1}} ⋯ X^{p_{m}}Y^{q_{m}})

holds in the algebra of formal power series in to non-commuting indeterminates X, Y over an arbitrary field of characteristic zero.

The proof uses the concept of the tensor product of algebras A, B over a field. The point is that A ⊗_{K} B carries a structure as algebra which is defined as follows: Consider the natural embedding

A ⟶ A ⊗_{K} B, a ⟼ a ⊗ 1_{B}

(similarly B → A × B). The product in A ⊗ B is characterized by

((a ⊗ b)(a′ ⊗ b′) = (ab) ⊗ (a′b′).

It is easy to construct this product by means of basises. The maps A, B → A × B are injective. We identify a ∈ A with a ⊗ 1 and b ∈ B with 1 ⊗ b and consider A, B as subalgebras of A × B in this way. The formula for the product shows that

ab = ba for a ∈ a, b ∈ B.

It is convenient now to use the notation

T = ⟨X, Y ⟩, T̂ = K⟨⟨X, Y ⟩⟩.

We use a second copy of this algebra in two different variables X̄, Ȳ and then consider the tensor product

K⟨X, Y ⟩ ⊗_{K} K⟨X̄, Ȳ ⟩.

We also consider the natural isomorphism

K⟨X, Y ⟩ ⟶^{∼} K⟨X̄, Ȳ ⟩, P ⟼ P,

which extends X ↦ X̄, Y ↦ Ȳ. There is a natural homomorphism of algebras

Δ : K⟨X, Y ⟩ ⟶ K⟨X, Y ⟩ ⊗_{K} K⟨X̄, Ȳ ⟩, P (X, Y) ⟼ P (X + X̄, Y + Ȳ).

For P ∈ KX + KY we have ΔP = P + P̄. But also

Δ[X, Y] = (X + X̄)(Y + Ȳ) − (Y + Ȳ)(X + X̄) = [X, Y] + [X̄, Ȳ],

because the variables with bars commute with the variables without bar. From this one easily concludes

Δ(P) = P + P̄ for all P ∈ L.

Actually the Lie algebra L ⊂ K⟨X, Y ⟩ is characterized by this property:

1.5 Basic Lemma.

L = { v ∈ K⟨X, Y ⟩; Δ(v) = v + v̄. }

It is useful to express this lemma also on the level o formal power series. For this we have to use the formal tensor product

K⟨⟨X, Y ⟩⟩ ˆ⊗_{K}K⟨⟨X̂, Ŷ ⟩⟩

This can be defined as the quotient of the formal algebra of power series in 4 indeterminates X, Y, X̄, Ȳ by the ideal which is generated by all

XX̄ − X̄X, XȲ − Ȳ X Y Ȳ − Ȳ Y.

There is a natural homomorphism

K⟨X, Y ⟩ ⊗_{K} K⟨X̄, Ȳ ⟩ ⟶ K⟨⟨X, Y ⟩⟩ ˆ⊗_{K}K⟨⟨X̂, Ŷ ⟩⟩,

which can be shown to be injective and Δ extends to homomorphism

Δ : K⟨⟨X, Y ⟩⟩ ⟶ K⟨⟨X, Y ⟩⟩ ˆ⊗_{K}K⟨⟨X̄, Ȳ ⟩⟩.

An obvious variant of the basic lemma states:

1.6 Basic Lemma, variant. The set

{ P ∈ K⟨⟨X, Y ⟩⟩; Δ(P) = P + P̄ }

consist of all power series P whose homogeneous parts P_{m} are contained in L.

Beweis. Sei v ∈ k⟨X, Y ⟩ ein Element mit Δ(v) = v + \overline{v}. Es genügt zu zeigen daß v von der Gestalt

v = c + u ; c ∈ k, u ∈ L

ist. Denn dann gilt Δ(v) = c + u + \overline{u} = v + \overline{v} = 2c + u + \overline{u}. Es folgt c = 0. Wir werden dieses Resultat auf folgendes kommutative Analogon zurückführen

1.7 Lemma. Seien Z_{1}, …, Z_{r}, \overline{Z}_{1}, … \overline{Z}_{r} 2r Variable. Wir betrachten folgenden Einsetzungshomomorphismus kommutativer Polynomringe

\begin{matrix}D : k[Z_{1}, …, Z_{r}] & → k[Z_{1}, …, Z_{r}, \overline{Z}_{1}, … \overline{Z}_{r}] \\ P = P (Z_{1}, …, Z_{r}) & ↦ \overline{P} = P (Z_{1} + \overline{Z}_{1}, …, Z_{r} + \overline{Z}_{r}) \end{matrix}

Sei P = P (Z_{1}, …, Z_{r}) ein Polynom für das D(P) Summe eines Polynoms in k[Z_{1}, …, Z_{r}] und eines Polynoms in k[\overline{Z}_{1}, … \overline{Z}_{r}] ist. D(P) enthält also keine gemischten Potenzprodukte in Z_{ν} und \overline{Z}_{μ}. Dann ist der Grad von P kleiner oder gleich 1. Das Polynom hat also die Form

P = c + ∑_{ν=1}^{r} c_{ν}Z_{ν}

Beweis des Lemmas. Wir betrachten die Taylorentwicklung

P (Z + \overline{Z}) = P (Z) + \frac{∂P (Z)}{∂Z_{1}} \overline{Z}_{1} + … + \frac{∂P (Z)}{∂Z_{r}} \overline{Z}_{r} + ⋯

Da D(P) keine gemischten Potenzprodukte enthält müssen alle partiellen Ableitungen \frac{∂P}{∂Z_{1}} (Z), …, \frac{∂P}{∂Z_{r}} (Z) konstante Polynome sein. Daraus folgt die Behauptung.

⊔ ⊓ Wir kehren zurück zum Beweis des “Basic Lemma“ 1.5.

Sei u_{1}, u_{2}, u_{3}, … eine Basis des Vektorraums L. Wir benutzen das Theorem III.4.2 von Poincaré-Birkhoff-Witt. Jedes Element v ∈ k⟨X, Y ⟩ kann dargestellt werden für eine genügend große Zahl r als

v = P (u) = P (u_{1}, …, u_{r})

= ∑_{ν=0}^{m} P_{ν}(u_{1}, …, u_{r})

P_{s}(u) = ∑_{ν_{1}+...+ν_{r}=s} c_{ν_{1}ν_{2}...ν_{r}}u_{1}^{ν_{1}}u^{ν_{2}} … u_{r}^{ν_{r}}

Diese Darstellung ist eindeutig. Die Gesammtheit aller solcher Elemente mit m ≤ n bilden einen Untervektorraum F_{n} = F_{n}k⟨X, Y ⟩ von k⟨X, Y ⟩. Die aufsteigende Kette der F_{n} heißt eine Filtrierung von k⟨X, Y ⟩. Der Ring k⟨X, Y; \overline{X}, \overline{Y} ⟩ besitzt dann als Basis über k die Monome

u_{1}^{ν_{1}} … u_{s}^{ν_{s}} \overline{u}_{1}^{ν_{1}} … \overline{u}_{s}^{ν_{s}}

Analog wie in k⟨X, Y ⟩ erklärt man eine Filtrierung durch die Untervektorräume F_{n} = F_{n}k⟨X, Y; \overline{X}, \overline{Y} ⟩ des Ringes k⟨X, Y; \overline{X}, \overline{Y} ⟩. Ein Kommutator u_{ν}u_{μ} − u_{μ}u_{ν} liegt in L, läßt sich also darstellen als endliche Linearkombination der Elemente u_{ν}. Daraus folgt für die Filtrierungen beider Ringe

\begin{aligned}F_{m}F_{n} & ⊆ F_{m+n} \\ v ∈ F_{m}, w ∈ F_{m} & ⇒ vw = wv modulo F_{m+n−1} \end{aligned}

Sei nun

P = P (u) = P (u_{1}, …, u_{r})

= ∑_{ν=0}^{m} P_{ν}(u_{1}, …, u_{r})

P_{s}(u) = ∑_{ν_{1}+...+ν_{r}=s} c_{ν_{1}ν_{2}...ν_{r}}u_{1}^{ν_{1}}u^{ν_{2}} … u_{r}^{ν_{r}}

ein Element mit Δ(P) = P + \overline{P} und mit m ≥ 2. Es genügt zu zeigen, daß P_{m} = 0. Wir wollen Lemma 1.7 anwenden. Aus den Eigenschaften der Filtrierung, insbesondere der Kommutativität modulo dieser Filtrierung folgt, daß für das ”kommutative“ homogene Polynom Q = P_{m}(z_{1}, …, z_{r}) ∈ k[z_{1} … z_{r}] gilt

D(Q) = Q(z + \overline{z}) = Q + \overline{Q}.
Aus Lemma 1.7 folgt dann wie gewünscht Q = 0.  ⊔ ⊓

Wir betrachten nun die nichtkommutative Potenzreihe

exp(X)exp(Y) ∈ k⟨⟨X, Y ⟩⟩

Eine Potenzreihe ohne konstanten Term kann in eine andere eingesetzt werden. Daher ist

Q = Q(X, Y) = log(exp(X)exp(Y)) = log(1 + (exp(X)exp(Y) − 1)) ∈ k⟨⟨X, Y ⟩⟩ erklärt. Weil X, Y mit \overline{X}, \overline{Y} vertauschbar sind, gilt

exp(X + \overline{X}) = exp(X)exp(\overline{X}), ⋯

Eine Potenzreihe in X, Y ist mit einer in \overline{X}, \overline{Y} vertauschbar. log überführt ein Produkt vertauschbarer Elemente in eine Summe. Es folgt

Δ(Q) = Q(X + \overline{X}, Y + \overline{Y}) = log(exp(X + \overline{X})exp(Y + \overline{Y}))
= log(exp(X)exp(\overline{X})exp(Y)exp(\overline{Y}))
= log((exp(X)exp(Y))(exp(\overline{X})exp(\overline{Y})))
= log((exp(X)exp(Y)) + log(exp(\overline{X})exp(\overline{Y}))
= Q + \overline{Q}

Daraus folgt nach Lemma 1.5, daß alle homogenen Komponenten der Potenzreihe

log(exp(X)exp(Y))

in der Liealgebra L liegen. Das beweist den wichtigsten ersten Teil des Satzes von Campbell-Hausdorff. Der zweite Teil, die genaue Form der Campbell-Hausdorff-Reihe kann aus dieser Existenzaussage ziemlich leicht gefolgert werden. Skizze: Wir betrachten die lineare Abbildung

ϕ : k⟨X, Y ⟩ ⟶ L

die einem Produkt v_{1}v_{2} … v_{m} von Elementen v_{ν} ∈ kX + kY bzw. einem Element v ∈ kX + k das Lieprodukt

[v_{1}, [v_{2}, … [v_{m−1}, v_{m}] …] ∈ L

bzw. das Element v zuordnet und die Konstanten auf 0 abbildet. Man zeigt durch Induktion nach m:

ϕ([v_{1}, [v_{2}, … [v_{m−1}, v_{m}] …]) = m[v_{1}, [v_{2}, … [v_{m−1}, v_{m}] …]) ∀v_{ν} ∈ kX + kY

Die Elemente aus L sind also Eigenvektoren von ϕ. Man wendet ϕ auf die Reihe

F (X, Y) = log(exp(X)exp(Y))

= ∑_{m=1}^{∞} ∑_{p_{i}+q_{i}≥1;p_{i},q_{i}≥0} \frac{(−1)^{m+1}}{mp_{1}!q_{1}! … p_{m}!q_{m}!}X^{p_{1}}Y^{q_{1}} … X^{p_{m}}Y^{q_{m}}

= ∑_{m=1}^{∞} F_{m}(X, Y)

–gliedweise– an und erhält:

ϕ(F (X, Y)) = ∑_{n=1}^{∞} ϕ(F_{n}(X, Y))

§2. Das Casimir-Element 55

Dabei ist F_{n} die homogene Polynomkomponente in X, Y von F vom Grad n, also die Zusammenfassung aller Glieder der Reihe vom Grad n. Diese Reihe läßt sich andererseits, wie wir bewiesen haben, darstellen als

∑_{n=1}^{∞} P_{n}

Dabei ist P_{n} eine endliche Linearkombination von Elementen der Form

[v_{1}, [v_{2}, … [v_{n−1}, v_{n}] …] v_{ν} ∈ kX + kY,

also ein Eigenvektor zum Eigenwert n.

ϕ(P_{n}) = nP_{n}

Durch Vergleich der homogenen Polynomkomponenten in X, Y vom Grad n erhält man

nP_{n} = ϕ(F_{n})

Damit sind die Liepolynome P_{n} bestimmt.

2. Das Casimir-Element

In diesem Abschnitt sollen Elemente des Zentrums der universellen Hülle U(L) einer Liealgebra L konstruiert werden. Ein Element C ∈ U(L) liegt im Zentrum von U(L), wenn es mit allen Elementen von U(L) vertauschbar ist. Sei ρ : L → End(V) eine Darstellung in einem Vektorraum V. Dann läßt sich ρ zu einem Homomorphismus ρ : U(L) → End(V) fortsetzen. Die lineare Abbildung

ρ(C) : V ⟶ V

ist dann mit allen Endomorphismen ρ(g), g ∈ L vertauschbar, ist also ein Homomorphismus von Moduln über L. Kern und Kokern von ρ(C) sind dann ebenfalls Moduln über der Liealgebra L.

Ist der Grundkörper der Körper der reellen Zahlen und L die Liealgebra einer zusammenhängenden Liegruppe G, so kann man die Elemente von U(L) als gegen Linkstranslationen invariante lineare Differentialoperatoren auf der Liegruppe interpretieren. Die Elemente des Zentrums von U(L) sind dann diejenigen linearen Differentialoperatoren, die sowohl bei Links- wie auch bei Rechtstranslatione invariant sind. Für die Analysis auf einer Liegruppe spielen diese vollinvarianten Differentialoperatoren eine ausgezeichnete Rolle.

Im folgenden werden wir der Einfachheit halber immer voraussetzen, daß der Grundkörper k die Charakteristik Null hat. Aus der linearen Algebra.

Sei V ein endlichdimensionaler Vektorraum der Dimension n. Sei f_{1}, …, f_{n} eine Basis von V. Wir betrachten den nichtkommutativen Polynomring R = k⟨x_{1}, …, x_{n}⟩ in den Variablen x_{1}, …, x_{n} über dem Grundkörper k und die lineare Abbildung

\begin{matrix}j : V & → R \\ ∑ c_{ν}f_{ν} & ↦ ∑ c_{ν}x_{ν} \end{matrix}

R, genauer j, ist eindeutig bis auf eindeutige Isomorphie durch eine universelle Abbildungseigenschaft gekennzeichnet, damit unabhängig von der Wahl der Basis: Zu jeder linearen Abbildung f von V in einen Ring, der k enthält, gibt es genau einen Ringhomomorphismus f̃: R → S, der auf k die Identität ist mit f = f̃∘ j. Wegen dieser Unabhängigkeit von der Wahl der Basis von V schreiben wir auch k⟨V ⟩ = R; V ⟶ k⟨V ⟩ ist dann die kanonische Abbildung in R.

Sei R_{2} ⊆ R der Untervektorraum der homogenen Polynome vom Grad 2. Er hat die Dimension n^{2}; eine Basis bilden die n^{2} Monome x_{ν}x_{μ}. Sei β(u, v) eine symmetrische Bilinearform auf V. Es gilt dann

\begin{aligned}β(u, v) & = β(v, u) \\ β(u, V) = 0 & ⇒ u = 0 \end{aligned}

Wir konstruieren die lineare Abbildung

R_{2} ⟶ End(V)

die dem Basiselement x_{ν}x_{μ} die lineare Abbildung y ↦ β(f_{ν}, y)f_{μ} zuordnet. Dem Element j(u)j(v) wird dann die lineare Abbildung y ↦ β(u, y)v zuordnet. Die Abbildung R_{2} ⟶ End(V) ist also unabhängig von der Auswahl der Basis. Die Abbildung R_{2} ⟶ End(V) ist injektiv, damit ein Isomorphismus: Sei ∑_{ν} j(u_{ν})x_{ν} ∈ R_{2} und ϕ : V → V die zugeordnete lineare Abbildung. Sei ϕ = 0. Dann gilt für jedes Element y ∈ V: ∑_{ν} β(u_{ν}, y)f_{ν} = 0. Es folgt β(u_{ν}, y) = 0 , ∀y ∈ V , ν = 1, …, n. Da β nichtausgeartet ist folgt u_{ν} = 0 ∀ν. Also verschwindet ∑_{ν} j(u_{ν})x_{ν} ∈ R_{2}.

Da wir vorausgesetzt haben, daß k die Charakteristik 0 hat, gibt es zu β eine Orthonormalbasis e_{1}, …, e_{n} :

β(e_{ν}, e_{μ}) = \{ \begin{matrix}1 & ν = μ \\ 0 & μ ≠ μ \end{matrix}

Dann gilt für ein Element y ∈ V

y = ∑_{ν} β(e_{ν}, y)e_{ν}

Sei c = ∑_{ν} j(e_{ν})j(e_{ν}) ∈ R_{2} und ϕ die zugehörige lineare Abbildung. Dann gilt ϕ(y) = ∑_{ν} β(e_{ν}, y)e_{ν} = y. Die zugehörende lineare Abbildung ist also die Identität. Das Element c = ∑_{ν} j(e_{ν})j(e_{ν}) ist unabhängig von der Wahl de Orthonormalbasis. Sei f : V ⟶ V eine bzgl. β schiefsymmetrische lineare Abbildung:

β(f(u), v) = −β(u, f(v))

Dann gilt 2.1 Lemma. Es gilt

∑_{ν}(j(f(e_{ν}))j(e_{ν}) + j(e_{ν})j(f(e_{ν}))) = 0

Beweis. Sei ϕ die dem Element ∑_{ν}(j(f(e_{ν}))j(e_{ν}) + j(e_{ν})j(f(e_{ν}))) zugeordnete Abbildung und seien u, v ∈ V. Es gilt

β(ϕ(u), v) = ∑_{ν}(β(f(e_{ν}), u)β(e_{ν}, v) + β(e_{ν}, u)β(f(e_{ν}), v))

= β(f(∑_{ν} β(e_{ν}, v)e_{ν}), u)) + β(f(∑_{ν} β(e_{ν}, u)e_{ν}), v))

= β(f(v), u) + β(f(u), v)

= 0

Sei jetzt L eine endlichdimensionale Liealgebra. Eine symmetrische Bilinearform β auf dem Vektorraum L heißt invariant , wenn gilt:

β([x, y], z) + β(y, [x, z]) = 0 ∀x, y, z ∈ L

Äquivalent dazu ist: Für alle Elemente x ∈ L ist die lineare Abbildung

\begin{matrix}L & → L \\ y & ↦ [x, y] \end{matrix}

schiefsymmetrisch. Sei ferner H ⊆ L ein Ideal von L. Die Einschränkung von β auf H sei nicht ausgeartet. Die universelle Hülle U(H) von H wird kanonisch injektiv in die universelle Hülle U(L) von L abgebildet. Wir fassen L als Untervektorraum von U(L) auf. Wir betrachten das natürliche kommutative Diagramm

H = V →^{j} k⟨V ⟩ → U(H)
↓  ↓
L  ↪  U(L)

Sei e_{1}, …, e_{n} eine Orthonormalbasis von H ⊆ L bzgl. β. Sei C = ∑_{ν} e_{ν}e_{ν} ∈ U(L). Dann ist C das Bild von ∑_{ν} j(e_{ν})j(e_{ν}) ∈ k⟨V ⟩, damit unabhängig von der Wahl der Orthonormalbasis. Sei g ∈ L. Aus dem vorhergehenden Lemma folgt das Verschwinden des Elementes

∑_{ν}(j([g, e_{ν}])j(e_{ν}) + j(e_{ν})j([g, e_{ν}]) ∈ k⟨V ⟩.

Deshalb verschwindet auch sein Bildelement in U(L) :

∑_{ν}([g, e_{ν}]e_{ν} + e_{ν}[g, e_{ν}]) = 0

Es gilt also in U(L)

gC − Cg = ∑_{ν}(ge_{ν}e_{ν} − e_{ν}e_{ν}g) = ∑_{ν}([g, e_{ν}]e_{ν} + e_{ν}[g, e_{ν}] + e_{ν}e_{ν}g − e_{ν}e_{ν}g) = 0

Das Element C ∈ U(L) ist mit allen Elementen g ∈ L, damit mit allen Elementen von U(L) vertauschbar, liegt also im Zentrum von U(L). Sei

ρ : L ⟶ End(M)

eine Darstellung der Liealgebra L im endlichdimensionalen Vektorraum M. ρ läßt sich kanonisch fortsetzen zu einem Ringhomorphismus von U(L) in den Endomorphismenring End(M) von M, den wir auch mit ρ bezeichnen. Wir bezeichnen mit Sp(f) die Spur von f ∈ End(M). Dann ist

β_{ρ}(x, y) = Sp(ρ(x)ρ(y))

eine invariante symmetrische Bilinearform auf L, die zu ρ assozierte Bilinearform.

2.2 Satz und Definition. Sei ρ eine Darstellung der Liealgebra L im endlichdimensionalen Vektorraum M. Sei H ⊆ L ein von Null verschiedenes Ideal von L und sei die Einschränkung der zu ρ assozierten Bilinearform β_{ρ} = β auf H nicht ausgeartet. Wir wählen eine Orthonormalbasis e_{1}, …, e_{n} von H bzgl. β. Das Element C = ∑_{ν} e_{ν}e_{ν} ∈ U(L) hängt nicht von der Auswahl der Orthonormalbasis ab. Das Element C liegt im Zentrum von U(L). Es heißt das Casimirelement zu ρ und zum Ideal H. Es gilt

(1) Sp(ρ(C)) = dim(H) = n

(2) ρ(C) : M ⟶ M ist ein Homomorphismus von Moduln über L. Wenn M einfacher L-Modul ist, wenn also M keinen von M und 0 veschiedenen L-Untermodul besitzt, ist ρ(C) ein Isomorphismus.

Zum Beweis. ρ(C) ist mit allen Elementen von L vertauschbar, damit ein Homomorphismus von Moduln über L.

Sp(ρ(C)) = ∑_{ν=1}^{n} Sp(ρ(e_{ν})ρ(e_{ν})) = ∑_{ν=1}^{n} β(e_{ν}, e_{ν}) = ∑_{ν=1}^{n} 1 = n

Insbesondere ist ρ(C) ≠ 0. Ist M ein einfacher Modul, so ist ρ(M) = M also ρ(C) ein Isomorphismus. Bemerkung. Wenn ρ = ad die adjungierte Darstellung von L ist , so nennt man die assozierte Bilinearform auch die Killingform der Liealgebra L.

§4. Auflösbare und nilpotente Liealgebren 59

4. Auflösbare und nilpotente Liealgebren

Die Sätze von Engel, Lie und Cartan Das Radikal einer Liealgebra

Auflösbare Liealgebren sind die Liealgebren auflösbarer Liegruppen; nilpotente Liealgebren gehören zu nilpotenten Liegruppen, einer Unterklasse der auflösbaren Liegruppen. Charakteristisches Beispiel für nilpotente Liealgebren sind die Lieunteralgebren der Liealgebra der nilpotenten Dreiecksmatrizen; das sind diejenigen Dreiecksmatrizen mit 0 als Diagonalelementen. Hauptresultat über nilpotente Liealgebren ist der Satz von Engel. Hauptresultate über auflösbare Liealgebren sind die Sätze von Lie und Cartan. Lie’s Theorem sagt aus, daß sich eine auflösbare Unterliealgebra des Endomorphismenringes End(V) eines Vektorraumes V bei geeigneter Wahl einer Basis von V immer als Unterliealgebra der Liealgebra der Dreiecksmatrizen darstellen läßt, falls der Grungkörper algebraisch abgeschlossen ist. Cartans Kriterium charakterisiert auflösbare Liealgebren mit Hilfe gewisser invarianter symmetrischer Bilinearformen auf der Liealgebra.

Der Grundkörper k hat im folgenden immer die Charakteristik 0. Wir werden der Einfachheit halber häufig annehmen, daß k = ℂ, der Körper der komplexen Zahlen ist. Der allgemeine Fall läßt sich mit Hilfe von Basiswechsel und Technik der endlich bzw. abzählbar vielen Konstanten ohne Probleme auf diesen speziellen Fall zurückführen. Alle Liealgebren sind endlichdimensional; ebenfalls im allgemeinen die Darstellungsmoduln.

Aus der linearen Algebra

Sei A : V → V eine lineare Abbildung eines endlichdimensionalen Vektorraumes in sich. A heißt diagonalisierbar, wenn die Matrix von A zu einer geeigneten Basis Diagonalgestalt hat. Ist A diagonalisierbar nach Grundkörpererweiterung zum algebraischen Abschluß von k, so heißt A halbeinfach. Sei zunächst k algebraisch abgeschlossen. Aus der jordanschen Normalform von A liest man ab, daß sich A als Summe

A = S + N

darstellen läßt einer mit A vertauschbaren halbeinfachen linearen Abbildung S und einer mit A vertauschbaren nilpotenten linearen Abbildung N. Man kann dann zeigen, daß S und N eindeutig durch diese Eigenschaften bestimmt sind. Man weist nach, daß es ein nicht unbedingt eindeutig bestimmtes Polynom P (x) ∈ k[x] gibt mit

P (A) = S , P (0) = 0

Mit Hilfe von Grundkörpererweiterung und Galoistheorie überträgt man dieses Resultat auf den Fall eines nicht algebraisch abgeschlossenen Grundkörpers.

4.1 Lemma und Definition. Sei V ein Vektorraum endlicher Dimension und sei A : V → V eine lineare Abbildung. Dann läßt sich A zerlegen als Summe

A = S + N

einer halbeinfachen Abbildung S und einer nilpotenten N. Dabei sind S und N mit A und damit auch untereinander vertauschbar. Diese Eigenschaften charakterisieren S, N eindeutig. S ist der halbeinfache, N der nilpotente Anteil von A. Es gibt ein Polynom P (x) ∈ k[x] mit

P (A) = S , P (0) = 0

S, N sind mit allen linearen Abbildungen vertauschbar, die mit A vertauschbar sind.

Wiederholung aus der Gruppentheorie

Gegeben sei eine Darstellung der Liealgebra L

ρ : L ⟶ End(V)

in dem endlichdimensionalen Vektorraum V, oder äquivalent dazu ein endlichdimensionaler L-Modul V: gv = ρ(g)(v) g ∈ L , v ∈ V. Es gibt dann eine Kette von Untermoduln

0 = V_{0} {⊂ \atop ≠} V_{1} {⊂ \atop ≠} … {⊂ \atop ≠} V_{n−1} {⊂ \atop ≠} V_{n} = V

der Länge n, die sich nicht verfeinern läßt; äquivalent dazu: die Faktormoduln V_{ν}/V_{ν−1} sind einfache Moduln. Eine solche Kette von Untermoduln heißt Jordan-Hölder Kette. In der Gruppentheorie zeigt man:

Alle Jordan-Hölder Ketten haben die gleiche Länge. Die Faktormoduln zweier Jordan-Hölder Ketten sind bis auf Permutation isomorph.

Besonders ausgezeichnet ist die adjungierte Darstellung ad = ad_{L} einer Liealgebra L

\begin{matrix}ad : L & → End(L) \\ g & ↦ (x ↦ ad(g)(x) = [g, x]) \end{matrix}

Der Kern der adjungierten Abbildung einer Liealgebra L ist ein Ideal, das Zentrum Z von L. Es ist als Unterliealgebra kommutativ.

Z = {x ∈ L | [x, y] = 0 ∀y ∈ L}

Das abgeleitete Ideal der Liealgebra L ist

BL = [L, L]

Dabei ist [L, L] der von allen Elemente [x, y] x, y ∈ L erzeugte Vektorraum. L/[L, L] ist kommutativ. Das abgeleitete Ideal von L ist im Kern eines jeden Homomorphismus von L in eine kommutative Liealgebra enthalten.

4.2 Definition, Abgeleitete Reihe, absteigende und aufsteigende Zentralreihe einer Liealgebra.

1) Wir erklären induktiv die Ideale B^{p}L durch

B^{0}L = L ; B^{p+1}L = [B^{p}L, B^{p}L]

Wir nennen abgeleitete Reihe von L die absteigende Reihe von Idealen

B^{0}L ⊇ B^{1}L ⊇ B^{2}L ⊇ …

Alle auftretenden Faktoren B^{p}L/B^{p+1}L sind kommutativ.

2) Wir definieren induktiv die Ideale Z^{p}L von L durch: Z^{0}L = L ; Z^{p+1}L = [L, Z^{p}L]. Die absteigende Zentralreihe von L ist die absteigende Kette von Idealen

Z^{0}L ⊇ Z^{1}L ⊇ Z^{2}L ⊇ …

Alle auftretenden Faktoren sind kommutativ; ad_{L}(x) operiert auf allen Faktoren trivial.

3) Wir definieren induktiv die Ideale Z_{p}L von L durch: Z_{0}L = 0; Z_{p+1}L ist das Urbild in L des Zentrums von L/Z_{p}L. Die aufsteigende Zentralreihe von L ist die aufsteigende Kette von Idealen

Z_{0}L ⊆ Z_{1}L ⊆ Z_{2}L ⊆ …

Alle auftretenden Faktoren sind kommutativ; ad_{L}(x) operiert auf allen Faktoren trivial.

4.3 Definition. Eine Liealgebra L heißt nilpotent, wenn es eine Kette von Idealen von L

L = L^{0} ⊇ L^{1} ⊇ … ⊇ L^{p−1} ⊇ L^{p} = 0

gibt mit [L, L^{ν}] ⊆ L^{ν+1}. Die letzte Eigenschaft ist äquivalent zu: Für alle g ∈ L induziert ad_{L}(g) in allen Faktoren L^{ν}/L^{ν+1} die Nullabbildung. Insbesondere sind die Faktoren kommutativ.

Da die Faktoren kommutativ sind, läßt sich die Kette in der Definition verfeinern zu einer Jordan-Hölder Kette, in der alle Faktoren die Dimension 1 haben.

4.4 Satz. Sei L eine Liealgebra. Folgende Eigenschaften sind äquvivalent:

1) L ist nilpotent.

2) Für die absteigende Zentralreihe gibt es eine natürliche Zahl p mit Z^{p}L = 0.

3) Für die aufsteigende Zentralreihe gibt es eine natürliche Zahl p mit Z_{p}L = L.

4) Die Faktoren einer, damit jeder Jordan-Hölder Reihe von Idealen haben die Dimension 1; auf den Faktoren operiert L mit der adjungierten Abbildung trivial.

5) Es gibt eine natürliche Zahl p, so daß für alle Elementfolgen x_{1}, …, x_{p} ∈ L gilt ad_{L}(x_{1}) ∘ ad_{L}(x_{2}) ∘ ⋯ ∘ ad_{L}(x_{p}) = 0.

4.5 Satz. Jede Unterliealgebra und jede Faktorliealgebra einer nilpotenten Liealgebra ist nilpotent.

4.6 Definition und Satz. Eine Liealgebra L heißt auflösbar, wenn eine der folgenden äquivalenten Eigenschaften erfüllt ist.

1) Für die abgeleitete Reihe gibt es eine natürliche Zahl p mit B^{p}L = 0

2) Es gibt eine Kette von Unteralgebren – nicht unbedingt Idealen–

L = L^{0} ⊇ L^{1} ⊇ … ⊇ L^{p−1} ⊇ L^{p} = 0

so daß L^{ν+1} Ideal in der Liealgebra L^{ν} ist und so daß alle Faktoralgebren kommutativ sind.

3) Das gleiche wie in 1) aber mit Faktoralgebren der Dimension 1.

4.7 Satz. Jede Unteralgebra und jede Faktoralgebra einer auflösbaren Liealgebra L ist auflösbar. Ist H ⊂ L ein Ideal, so daß die Liealgebren H und L/H auflösbar sind, so ist L auflösbar. Die Summe zweier auflösbarer Ideale einer Liealgebra ist auflösbar. Insbesondere gibt es ein größtes auflösbares Ideal, das dann alle auflösbaren Ideale enthält.

4.8 Definition. Das größte auflösbare Ideal 𝔯 einer Liealgebra L heißt das Radikal der Liealgebra L. Äquivalente Charakterisierung:

Das Ideal 𝔯 ist auflösbar und L/𝔯 enthält kein kommutatives Ideal, oder gleichwertig dazu, kein auflösbares Ideal, das von 0 verschieden ist.

4.9 Definition. Eine Liealgebra L heißt halbeinfach, wenn eine der folgenden äquivalenten Bedingungen erfüllt ist:

1) Das Radikal von L verschwindet.

2) L enthält kein kommutatives Ideal.

4.10 Bemerkung. Die Faktoralgebra einer Liealgebra nach seinem Radikal ist halbeinfach. Eine einfache Liealgebra L, d.h. eine Liealgebra, die kein Ideal ≠ 0, L besitzt und die außerdem nicht kommutativ ist ist halbeinfach. Beispiel: Die Matrizen aus M_{n}(k) mit Spur 0 bilden eine einfache Liealgebra für n ≥ 2.

Die Untersuchung nilpotenter Liealgebren

4.11 Lemma. Sei V ein Vektorraum und x ∈ End(V) ein nilpotentes Element. Dann ist ad_{End(V)}(x) : End(V) → End(V) nilpotent.

Beweis. (ad_{End(V)}(x))^{n}(y) ist eine Linearkombination von Elementen x^{ν}yx^{μ} ∈

End(V) mit ν + μ = n.  ⊔ ⊓

4.12 Theorem von Engel. Sei L ⊆ End(V) eine Unterliealgebra des Endomorphismenringes End(V) des Vektorraumes V ≠ 0 aufgefaßt als Liealgebra. Alle Elemente x ∈ L seien nilpotente Endomorphismen. Dann gibt es einen von 0 verschiedenen Vektor v ∈ V mit g(v) = 0 ∀g ∈ L.

Beweis. Vollständige Induktion nach der Dimension n = dim(L). Sei n > 0. Sei H {⊂ \atop ≠} L eine Unterliealgebra. Es gilt: h ∈ H ⇒ ad_{L}h(H) ⊆ H. Also induziert für h ∈ H ad_{L}h eine lineare Abbildung σ(h) : L/H → L/H. Nach dem vorhergehenden Lemma ist ad_{L}h, damit auch σ(h) nilpotent. Ferner gilt: σ([x, y]) = σ(x) ∘ σ(y) − σ(y)∘σ(x). Nach Induktionsvoraussetzung gibt es in L/H ein von Null verschiedenes Element das durch σ(H) annulliert wird. Deshalb gibt es ein Element x ∈ L, das nicht in H liegt mit [x, H] ⊆ H. Dann ist H ein Ideal der Liealgebra H̃ = ky + H. Es gilt dim(H̃/H) = 1. Durch Induktion erhält man, daß es in L ein Ideal H der Kodimension 1 gibt. Nach Induktionsvoraussetzung ist der Untervektorraum W = {v ∈ V |Hv = 0} von Null verschieden. Sei g ∈ L , h ∈ H , w ∈ W. Dann gilt: hgw = [h, g]w + ghw = 0 + 0 = 0. Es folgt: LW ⊆ W. Wähle ein Element α ∈ L α ∉ H. Dann ist L = kα + H. α operiert nilpotent auf V. Deshalb besitzt α einen von Null verschiedenen Eigenvektor v ∈ W zum Eigenwert 0. Es gilt dann

Lv = 0.  ⊔ ⊓

4.13 Folgerung. Die Liealgebra L ist genau dann nilpotent, wenn ad_{L}x für alle x ∈ L nilpotent ist.

Beweis.Sei L ≠ 0. Dann ist nach dem Satz von Engel das Zentrum Z = {a ∈ L | ad_{L}x(a) = 0 ∀x ∈ L} von L von Null verschieden. Man führt den gleichen Schluß für L/Z durch. . . .. Es folgt, daß die aufsteigende Zentralreihe mit L endet.

⊔ ⊓

4.14 Folgerung. Sei ρ : L → End(V) eine injektive Darstellung der Liealgebra L. Seien alle Endomorphismen ρ(x) ∈ End(V) , x ∈ L nilpotent. Nach Lemma 4.11 sind dann auch die Elemente ad_{End(V)}(ρ(x)) , x ∈ L nilpotent. Nach der vorhergehenden Folgerung ist also ρ(L) damit auch L nilpotente Liealgebra.

Gegeben sei eine Darstellung

ρ : L ⟶ End(V)

der Liealgebra L im – endlichdimensionalen – Vektorraum V. Wir betrachten Ideale 𝔞 ⊆ L so daß gilt: ρ(x) ist nilpotent ∀x ∈ 𝔞. Ein solches Ideal heißt nilpotent bzgl. ρ. Wie schon bewiesen wurde ist dann ρ(𝔞) eine nilpotente Liealgebra (Folgerung 13 ). Sei zunächst ρ eine einfache Darstellung, also V ein einfacher L-Modul. Sei W = {v ∈ V | ρ(𝔞)v = 0}. Wie schon vorher gezeigt wurde (siehe Beweis von Engels Theorem), gilt: ρ(L)(W) ⊆ W. Weil W nach Engels Theorem nicht verschwindet und weil V einfacher L-Modul ist folgt W = V. Im allgemeinen Fall sei

0 = V_{0} ⊆ V_{1} ⊆ V_{2} ⊆ … ⊆ V_{r−1} ⊆ V_{r} = V

eine Jordan-Hölder Reihe des L-Moduls V. Die Faktoren F_{ν} = V ν/V_{ν−1} sind dann einfache L-Moduln. Für das bzgl. ρ nilpotente Ideal 𝔞 gilt dann 𝔞F_{ν} = 0.

Ist umgekehrt x ∈ L ein Element mit xF_{ν} = 0 für ν = 1, …, r so folgt ρ(x)(V_{ν}) ⊆ V_{ν−1} , ν = 1, …, r damit ρ(x)^{r}(V) ⊂ V_{0} = 0; also ist ρ(x)^{r} = 0. Der Durchschnitt der Kerne der Abbildungen

L → End(F_{ν})

ist ein bzgl. ρ nilpotentes Ideal. Es enthält alle bzgl. ρ nilpotente Ideale.

4.15 Definition und Satz. Sei ρ : L ⟶ End(V) eine Darstellung der Liealgebra L. Dann gibt es ein größtes bzgl. ρ nilpotentes Ideal 𝔫_{ρ}, das alle bzgl. ρ nilpotenten Ideale enthält. Es ist der Durchschnitt der Kerne aller induzierten Darstellungen in den Faktormoduln einer Jordan-Hölder Reihe.

4.16 Definition und Satz. Es gibt ein größtes nilpotentes Ideal 𝔫 in der Liealgebra L. Es is das größte nilpotente Ideal bzgl. der adjungierten Darstellung ad_{L}.

Für den Beweis des folgenden Satzes benötigen wir ein Lemma

4.17 Lemma. Sei V ein einfacher Modul über der Liealgebra L und H ⊆ L ein Ideal. Wir fassen V als Modul über H auf. Sei

0 = V_{0} {⊂ \atop ≠} V_{1} {⊂ \atop ≠} V_{2} {⊂ \atop ≠} … {⊂ \atop ≠} V_{r−1} {⊂ \atop ≠} V_{r} = V

eine Jordan-Hölder Reihe dieses H-Moduls. Dann sind alle Faktormoduln F_{ν} = V_{ν}/V_{ν−1} isomorph.

Denn: Sei V ≠ 0 und W ein von Null verschiedener einfacher H-Untermodul von V. Für g ∈ L, h ∈ H, w ∈ W gilt: hgw = [h, g]w + ghw ≡ ghw modulo W. Die zusammengesetzte Abbildung

W ⟶^{g} gW ⟶ V/W

von W in V/W ist also ein Homomorphismus von H-Moduln. Das Bild von W bei diesem Homomorphismus in V/W sei der H-Modul W_{g} {⊂ \atop ≠} V/W. Weil W einfacher H-Modul ist, ist W_{g} isomorph zum einfachen H-Modul W, oder es gilt W_{g} = 0. Weil V einfacher L-Modul ist gilt

V/W = ∑_{g∈L} W_{g}

Damit besitzt der H-Modul V/W eine Jordan-Hölderreihe, deren Faktoren iso-

morph zu gewissen Moduln W_{g} sind, damit auch zu W.  ⊔ ⊓

4.18 Satz. Sei L eine Liealgebra und 𝔯 das Radikal von L. Dann ist 𝔯 ∩ [L, L] bzgl. jeder Darstellung nilpotent. Insbesondere ist 𝔯 ∩ [L, L] eine nilpotente Liealgebra.

4.19 Bemerkung, Umkehrung. Sei 𝔞 im Kern aller einfachen Darstellungen enthalten, oder, äquvalent dazu, sei nilpotent bzgl. aller Darstellungen. Dann ist 𝔞 nilpotentes, damit auflösbares Ideal, also im Radikal enthalten. Die Faktoralgebra \overline{L} = L/[L, L] ist kommutativ. Eine eindimensionale, damit einfache Darstellung \overline{L} → End(k) = k ist dann jede lineare Abbildung ϕ : \overline{L} → k = End(k). Der Durchschnitt der Kerne dieser Abbildungen ϕ ist 0. Deshalb ist 𝔞 auch in [L, L] enthalten. 𝔯 ∩ [L, L] ist der Durchschnitt der Kerne aller einfachen Darstellungen von L.

Beweis des Satzes. Wir wollen der Einfachheit halber voraussetzen, daß der Grundkörper k algebraisch abgeschlossen ist. Der allgemeine Fall kann leicht auf diesen mittels Grundkörpererweiterung zurückgeführt werden. Sei ρ : L → End(V) eine einfache Darstellung von L in einem Vektorraum V ≠ 0. Sei 𝔞 = 𝔯 ∩ [L, L]. Dann ist 𝔞 eine auflößbare Liealgebra. Wir benutzen an dieser Stelle den erst im nächsten Abschnitt über auflösbare Liealgebren behandelten Satz von Lie. Nach diesem Satz gibt es einen gemeinsammen Eigenvektor v ≠ 0 für 𝔞. Es gibt also eine lineare Abbildung ϕ : 𝔞 → k mit: ρ(x)v = ϕ(x)v ∀x ∈ 𝔞. kv ist ein einfacher nicht verschwindender 𝔞-Modul. Nach dem vorhergehenden Lemma 4.17 sind alle Faktoren einer Jordan-Hölde Reihe des 𝔞-Moduls V isomorph zu kv. Sei n ≥ 1 die Anzahl dieser Faktoren. Dann gilt für die Spur von ρ(x) : V → V x ∈ 𝔞 :

Sp(ρ(x)) = nϕ(x)

Andererseits gilt für ein Element x = [y, z], damit für jedes Element x ∈ [L, L]: Sp(ρ(x)) = 0. Es folgt

ϕ(x) = 0 ∀x ∈ 𝔞

W = {u ∈ V | ρ(𝔞)u = 0} enthält den Vektor v, ist also von Null verschieden. Wie wir im Beweis von Engels Theorem gezeigt haben gilt ρ(L)W ⊆ W. Weil V

einfacher L-Modul ist folgt V = W. 𝔞 operiert trivial auf V.  ⊔ ⊓

4.20 Folgerung. Sei L eine auflösbare Liealgebra. Dann ist [L, L] nilpotentes Ideal.

Sei ρ : L → End(V) eine Darstellung einer auflösbaren Liealgebra L. Sei L ein einfacher L-Modul. Dann ist ρ([L, L]) = 0. Da L/[L, L] kommutativ ist, sind alle Endomorphismen ρ(g) , g ∈ L vertauschbar. Es folgt, daß die Elemente g ∈ L, für die ρ(g) nilpotent ist, ein Ideal in L bilden. Mit Hilfe einer Jordan-Hölder Reihe folgert man, daß auch für eine beliebige Darstellung ρ von L die Menge der g ∈ L, für die ρ(g) nilpotent ist, ein Ideal von L ist.

4.21 Folgerung. Sei 𝔯 das Radikal der Liealgebra L, ρ : L → End(V) eine injektive Darstellung und 𝔫 das größte nilpotente Ideal bzgl. ρ. Dann gilt

x ∈ 𝔫 ⟺ x ∈ 𝔯 und ρ(x) ist nilpotent.

Zum Beweis. 𝔫 ist nilpotentes Ideal, also auflösbar, also im Radikal enthalten. Die Menge H = {x ∈ 𝔯 | ρ(x) nilpotent } ist ein Ideal von 𝔯. Es gilt

[L, H] ⊆ [L, L] ∩ 𝔯 ⊆ 𝔫 ⊆ H
Deshalb ist H ein bzgl. ρ nilpotentes Ideal von L.  ⊔ ⊓

Die Untersuchung auflösbarer Liealgebren

Aus der linearen Algebra

4.22 Lemma. Sei V ein Vektorraum, H ⊆ End(V) ein Untervektoraum der Endomorphismenalgebra. Seien v_{1}, …, v_{n} ∈ V von Null verschiedene gemeinsame Eigenvektoren aller Endomorphismen aus M. D.h.: Es gibt lineare Abbildungen ϕ_{1}, …, ϕ_{n} : H → k mit h(v_{ν}) = ϕ_{ν}(h)v_{ν} ∀h ∈ H. Die linearen Abbildungen ϕ_{1}, …, ϕ_{n} seien paarweise verschieden. Dann sind die Vektoren v_{1}, …, v_{n} ∈ V linear unabhängig.

4.23 Folgerung. Sei ρ : L → End(V) eine Darstellung der Liealgebra L. Es gibt dann nur endlich viele eindimensionale L-Untermoduln von V, die untereinander nicht isomorph sind.

4.24 Theorem von Lie. Der Grundkörper k sei algebraisch abgeschlossen. Sei ρ : L → End(V) eine Darstellung der auflösbaren Liealgebra L in einem von Null verschiedenen – endlichdimensionalen – Vektorraum V. Dann gibt es einen von Null verschiedenen gemeinsamen Eigenvektor v ∈ V für ρ(L) : ρ(L)(v) ⊆ kv.

4.25 Folgerung. Jede einfache Darstellung der auflösbaren Liealgebra L in einem Vektorraum V ≠ 0 ist eindimensional.

4.26 Folgerung. Sei ρ : L → End(V) eine Darstellung der Liealgebra L. Dann besitzt V eine Basis, so daß ρ(L) bzgl. dieser Basis durch Dreiecksmatrizen dargestellt wird.

Beweis des Theorems. Wir nehmen der Einfachheit halber an, daß der Grundkörper k = ℂ der Körper der komplexen Zahlen ist. Der allgemeine Fall kann wieder auf diesen speziellen zurückgeführt werden mit Hilfe von Grundkörpererweiterung und der Technik der endlich vielen Konstanten. Dann ist V = ℂ^{n}. Es gibt die Exponentialfunktion, dargestellt durch die Exponentialreihe

exp : M = End(V) = M_{n}(ℂ) ⟶ Gl(V) = Gl_{n}(ℂ)

Sei A ∈ End(V). Wir betrachten die einparametrige Untergruppe

ϕ : ℝ ⟶ Gl(M)

gegeben durch

ϕ(t)(x) = exp(tA)x · exp(−tA) ; t ∈ ℝ, x ∈ M

Dabei wird M als Vektorraum über ℝ ⊆ ℂ und Gl(M) als relle Liegruppe aufgefaßt. Die Ableitung ϕ′(0) ist dann die lineare Abbildung

\begin{matrix}ad_{M} : & M → M \\ x & ↦ [A, x] = Ax − xA \end{matrix}

Wie wir früher bewiesen haben gilt dann

exp(tA)x · exp(−tA) = exp(t · ad_{M}A)(x)

Wir können annehmen, daß L ⊆ End(V) = M. Wir beweisen den Satz durch Induktion nach der Dimension von L. Sei L ≠ 0. Weil L auflösbar ist gibt es ein Ideal H ⊆ L der Kodimension 1 in L. Sei A = g ∈ L ⊆ M. Dann gilt

ad_{M}g(x) ∈ \{ \begin{matrix}L & x ∈ L \\ H & x ∈ H \end{matrix}

Es folgt

exp(g)xexp(−g) ∈ \{ \begin{matrix}L & x ∈ L \\ H & x ∈ H \end{matrix}

Nach Induktionsvoraussetzung gibt es für H einen gemeinsamen Eigenvektor v ≠ 0 in V. Es gibt also eine lineare Abbildung ψ : H → k mit hv = ψ(h)v. Für g ∈ L , h ∈ H gilt dann

h(exp(g)v) = ψ(exp(−g)h exp(g))exp(g)v

exp(g)v ist also gemeinsamer Eigenvektor aller Elemente des Ideals H zur Linearfunktion ψ(exp(−g)hexp(g)) auf H. Betrachte für alle t ∈ ℝ die Linearfunktionen

\begin{matrix}ψ_{t} : H & → k \\ h & ↦ ψ(exp(−tg)h exp(tg)) \end{matrix}

Sie gehören zu eindimensionalen H-Untermoduln von V. Da es nur endlich viele untereinander nichtisomorphe solche Untermoduln gibt, gibt es nur endlich viele verschiedene ψ_{t}. Aus Stetigkeitsgründen sind deshalb alle ψ_{t} gleich: ψ_{t} = ψ_{0} = ψ. Betrachte

W = {w ∈ V | hw = ψ(h)w}

W enthält v, ist also von Null verschieden. Für alle Elemente g ∈ L und t ∈ ℝ gilt exp(tg)W ⊆ W. Durch differenzieren nach t in t = 0 erhält man

gW ⊆ W ∀g ∈ L

Wähle ein Element g_{0} ∈ L mit L = kg_{0} + H. Da k algebraisch abgeschlossen ist, besitzt g_{0} einen von 0 verschiedenen Eigenvektor w ∈ W. w ist dann gemeinsamer

Eigenvektor für ganz L.  ⊔ ⊓

Invariante Bilinearformen. Cartans Kriterium für Auflösbarkeit

4.27 Definition. Sei L eine Liealgebra und β(x, y) eine symmetrische Bilinearform auf L. Diese Bilinearfunktion β(x, y) heißt invariant, wenn gilt

β([x, y], z) + β(y, [x, z]) = 0 ∀x, y, z ∈ L

Ist M eine Teilmenge von L, so ist das orthogonale Komplement von M:

M^{⊥} = {x ∈ L | β(x, y) = 0 ∀y ∈ M}

4.28 Bemerkung. Aus der Invarianz der Bilinearform β folgt, daß das orthogonale Komplement H^{⊥} eines Ideals H wieder ein Ideal ist. Sei Z das Zentrum von L. Es gilt [L, L]^{⊥} ⊇ Z. Ist β nicht ausgeartet, so gilt [L, L]^{⊥} = Z.

4.29 Beispiel. Sei ρ → End(V) eine Darstellung der Liealgebra L im Vektorraum V. Dann ist β_{ρ}(x, y) = Sp(ρ(x)ρ(y)) eine symmetrische invariante Bilinearform, die zur Darstellung ρ assozierte Bilinearform. Dabei ist Sp die Spur für einen Endomorphismus. Ist ρ = ad_{L} die adjungierte Darstellung, so heißt die zu ad_{L} assozierte Bilinearform die Killingform κ(x, y).

4.30 Lemma. Sei A ∈ End(V) ein Endomorphismus des Vektorraumes V und A = A_{s} + A_{n} die kanonische Zerlegung von A in den halbeinfachen Anteil A_{s} und in den nilpotenten Anteil A_{n}. Betrachte die adjungierte Abbildung zu A

\begin{matrix}ad A : End(V) & → End(V) \\ x & ↦ [A, x] = Ax − xA \end{matrix}

Dann ist ad A = ad A_{s} +ad A_{n} die Zerlegung von ad A in den halbeinfachen Anteil ad A_{s} und den nilpotenten Anteil.

Beweis. Wir können uns auf den Fall eines algebraisch abgeschlossenen Grundkörpers k beschränken. Es gilt

ad A = ad A_{s} + ad A_{n}

Sei e_{1}, …, e_{n} eine Basis von Eigenvektoren von A_{s}. Dann gilt Ae_{ν} = λ_{ν}e_{ν}. Seien E_{ij} die Basiselemente von V mit

E_{ij}e_{ν} = \{ \begin{matrix}e_{i} & ν = j \\ 0 & ν ≠ j \end{matrix}

Dann gilt

ad A_{s} E_{ij} = (λ_{i} − λ_{j})E_{ij}

Also ist ad A_{s} halbeinfach. Wie früher bewiesen (siehe Lemma 4.11) ist ad A_{n}

nilpotent. Es gilt ad A_{s} ∘ ad A_{n} − ad A_{n} ∘ ad A_{s} = ad [A_{s}, A_{n}] = 0.  ⊔ ⊓

4.31 Theorem von Cartan. Sei L ⊆ End(V) eine Unterliealgebra des Endomorphismenringes eines Vektorraumes V. Wir betrachten die symmetrische invariante Bilinearform β(x, y) = Sp(xy) auf L. Äquivalent sind

1) L ist auflösbar.

2)  β(x, [y, z]) = 0 ∀x, y, z ∈ L.

Anders ausgedrückt

L⊥[L, L]

Beweis. Der Einfachheit halber nehmen wir an, daß der Grundkörper k = ℂ der Körper der komplexen Zahlen ist. Der allgemeine Fall läßt sich mit Hilfe von Grundkörpererweiterung auf diesen speziellen zurückführen. Sei λ ∈ ℂ. Wir bezeichnen mit \overline{λ} ∈ ℂ die konjugiert komplexe Zahl.

Sei L auflösbar. Aus Lie’s Theorem folgt, daß bei einer geeigneten Basis von V die Liealgebra L dargestellt wird durch eine Liealgebra von Dreiecksmatrizen. Es folgt, daß für x, y, z ∈ L die Matrix des Endomorphismus x[y, z] ∈ End(V) nur Nullen in der Diagonale enthält. Deshalb ist x[y, z] ∈ End(V) nilpotent und β(x, [y, z]) = Sp(x[y, z]) = 0.

Die umgekehrte Richtung: Sei L⊥[L, L]. Sei x ∈ [L, L] und sei

x = x_{s} + x_{n}

die Zerlegung von x in den halbeinfachen Anteil x_{s} ∈ End(V) und den nilpotenten Anteil x_{n}. Wir wollen zeigen, daß x_{s} verschwindet, daß also x nilpotent ist. Aus Engels Theorem (Folgerung 4.14) folgt dann, daß [L, L] ein nilpotentes Ideal ist, damit auch ein auflösbares Ideal. Da L/[L, L] kommutativ ist, folgt die Auflösbarkeit von L.

Die Zerlegung von ad_{End(V)}x = ad x : End(V) → End(V) in den halbeinfachen und den nilpotenten Anteil ist nach dem vorhergehenden Lemma:

ad x = ad x_{s} + ad x_{n}

Sei v_{1}, …, v_{n} eine Basis des Vektorraumes V aus Eigenvektoren von x_{s} zu den Eigenwerten λ_{ν}. Dann bilden die Endomorrphismen E_{ij}

E_{ij}v_{ν} = \{ \begin{matrix}v_{i} & ν = j \\ 0 & ν ≠ j \end{matrix}

eine Basis von End(V). Es gilt

\begin{aligned}x_{s}(v_{ν}) & = λ_{ν}v_{ν} \\ ad x_{s}(E_{νμ}) & = (λ_{ν} − λ_{μ})E_{νμ} \end{aligned}

Wir betrachten die “konjugiert komplexe” lineare Abbildung \overline{x}_{s} von x_{s}.

\overline{x}_{s}(v_{ν}) = \overline{λ}_{ν}v_{ν}

Dann ist auch ad \overline{x}_{s} halbeinfach und es gilt:

ad \overline{x}_{s}(E_{νμ}) = (\overline{λ}_{ν} − \overline{λ}_{μ})E_{νμ}

Es gibt dann Polynome P (t), Q(t) ∈ k[t] ohne konstanten Term mit

P (ad x) = ad x_{s} , Q(λ_{ν} − λ_{μ}) = \overline{λ}_{ν} − \overline{λ}_{μ}

Sei R(t) = Q(P (t)). Dann gilt

R(ad x) = ad \overline{x}_{s} , R(0) = 0.

Es folgt

[\overline{x}_{s}, L] = ad \overline{x}_{s}(L) = R(ad x)(L) ⊆ [L, L].

Sei x = ∑_{ν}[x_{ν}, y_{ν}] mit x_{ν}, y_{ν} ∈ L. Dann ist [\overline{x}_{s}, x_{ν}] ∈ [L, L], also nach Voraussetzung y_{ν}⊥[\overline{x}_{s}, x_{ν}], und es gilt

β(x, \overline{x}_{s}) = ∑_{ν} β([x_{ν}, y_{ν}], \overline{x}_{s}) = ∑_{ν} β(y_{ν}, [\overline{x}_{s}, x_{ν}]) = 0

Es gibt ein Polynome H(t) mit H(λ_{ν}) = \overline{λ}_{ν}. Dann folgt, daß \overline{x}_{s} = H(x_{s}) mit x_{n} vertauschbar ist, weil x_{s} mit x_{n} vertauschbar ist. Deshalb ist mit x_{n} auch x_{n}\overline{x}_{s} nilpotent, und es gilt Sp(x_{n}\overline{x}_{s}) = 0. Es folgt

\begin{aligned}0 = β(x, \overline{x}_{s}) & = Sp(x\overline{x}_{s}) = Sp(x_{s}\overline{x}_{s}) + Sp(x_{n}\overline{x}_{s}) = Sp(x_{s}\overline{x}_{s}) \\ & = ∑_{ν} λ_{ν} \overline{λ}_{ν} = ∑_{ν} |λ_{ν}|^{2} \end{aligned}

Also verschwinden alle Eigenwerte λ_{ν} und damit auch x_{s}.  ⊔ ⊓

4.32 Folgerung. Sei κ(x, y) die Killingform der Liealgebra L. Dann ist L genau dann auflösbar, wenn gilt

L⊥[L, L]

Beweis. Der Kern der Abbildung ad_{L} : L → End(L) ist das Zentrum von L. Deshalb ist L genau dann auflösbar, wenn das Bild ad_{L}(L) ⊆ End(L) auflösbar ist. Man

wendet Cartans Theorem auf dieses Bild an.  ⊔ ⊓

5. Halbeinfache Liealgebren

Hauptresultat dieses Paragraphen ist Satz 5.7: Jeder Modul über einer halbeinfachen Liealgebra L ist halbeinfach, also direkte Summe einfacher Untermoduln.

Auch in diesem Abschnitt nehmen wir an, daß der Grundkörper k die Charakteristik 0 hat. Alle betrachteten Lieagebren und alle Darstellungen von Liealgebren haben in der Regel endliche Dimension.

Wiederholung Aus dem vorhergehenden Abschnitt über auflösbare und nilpotente Liealgebren: Eine Liealgebra L ist halbeinfach, wenn das Radikal von L verschwindet, oder, äquivalent dazu, wenn L kein von 0 verschiedenes kommutatives Ideal enthält.

Mit Hilfe von Cartans Auflösbarkeitskriterium zeigt man

5.1 Satz. Sei L eine Liealgebra und κ(x, y) = Sp(ad_{L}(x)ad_{L}(y)) die Killingform von L. Dann ist L genau dann halbeinfach, wenn die Killingform nicht ausgeartet ist.

§5. Halbeinfache Liealgebren 71

5.2 Bemerkung. Ebenso zeigt man: Sei L eine halbeinfache Liealgebra, sei ρ : L → End(V) eine treue Darstellung von L und sei β(x, y) = Sp(ρ(x)ρ(y)) die assozierte invariante symmetrische Bilinearform. Dann ist β nicht ausgeartet; d.h. L^{⊥} = 0. Die Umkehrung gilt aber im allgemeinen nicht!

Beweis. Sei L halbeinfach. Da das Zentrum von L verschwindet ist die Darstellung ρ = ad_{L} : L → End(L) treu. Sei 𝔞 = L^{⊥} das orthogonale Komplement von L bei κ. 𝔞 ist ein Ideal und es gilt 𝔞⊥[𝔞, 𝔞]. Aus Cartans Auflösbarkeitskriterium (Theorem 4.31 des vorhergehenden Paragraphen ) folgt dann, daß 𝔞 auflösbar ist. Da L halbeinfach ist verschwindet jedes auflösbare Ideal.

Sei die Killingform nicht ausgeartet. Sei H ⊆ L ein kommutatives Ideal von L. Gegeben seien Elemente x ∈ H ; y, z ∈ L. Dann gilt

\begin{aligned}ad_{L}y(ad_{L}x(ad_{L}y(z))) = [y, [x, [y, z]]] & ∈ H \\ ad_{L}x(H) & = 0 \\ ad_{L}x(ad_{L}y(ad_{L}x(ad_{L}y(z)))) & = 0 \end{aligned}

Es folgt: (ad_{L}x ∘ ad_{L}y)^{2} = 0. ad_{L}x ∘ ad_{L}y ist nilpotent und deshalb verschwindet κ(x, y) = Sp(ad_{L}x ∘ ad_{L}y). Es gilt also

H⊥L

Da κ nicht ausgeartet ist, verschwindet H. L enthält also kein nicht verschwindendes

kommutatives Ideal und ist daher halbeinfach.  ⊔ ⊓

5.3 Folgerung. Sei L eine halbeinfache Liealgebra. Dann gilt

L = [L, L].

Beweis. Die Killingform κ ist nicht ausgeartet; deshalb ist das orthogonale Komplement von [L, L] das Zentrum Z von L: [L, L]^{⊥} = Z = 0. Da κ nicht ausgeartet ist, folgt [L, L] = 0. ⊔ ⊓

Sei H ⊆ L ein Ideal der halbeinfachen Liealgebra L. Dann ist das orthogonale Komplement H^{⊥} bei der Killingform κ wieder ein Ideal. Für das Ideal 𝔞 = H ∩ H^{⊥} gilt dann 𝔞⊥𝔞 also auch 𝔞⊥[𝔞, 𝔞]. Wieder folgt aus Cartans Auflösbarkeitskriterium, daß das Ideal 𝔞 auflösbar ist und deshalb verschwindet. Es gilt also H ∩H^{⊥} = 0 und daher L = H ⊕ H^{⊥}. Außerdem: [H, H] ⊆ H ∩ H^{⊥} = 0. Man folgert, daß L direkte Summe einfacher Ideale ist; wie man leicht sieht, sind diese eindeutig bestimmt.

5.4 Satz. Sei L eine halbeinfache Liealgebra. Dann gibt es zu jedem Ideal A ⊆ L ein Komplementärideal B ⊆ L :

L = A ⊕ B

Es folgt, das A und B vertauschbar sind: [A, B] = 0. Insbesondere ist L direkte Summe der einfachen nicht verschwindenden Ideale 𝔞_{ν}:

L = ⊕_{ν}𝔞_{ν}

Man sagt auch, daß die adjungierte Darstellung halbeinfach ist.

Im folgenden wollen wir dieses Resultat verallgemeinern.

5.5 Definition. Sei L → End(V) eine Darstellung der Liealgebra L im Vektorraum V. Diese Darstellung bzw. der Modul V über L heißt halbeinfach oder vollständig reduzibel, wenn eine der folgenden äquivalenten Bedingungen erfüllt ist:

1) Zu jedem Untemodul W ⊆ V gibt es einen Komplementärmodul W̃ ⊆ V : V = W ⊕ W̃ .

2) V ist direkte Summe einfacher Untermoduln, d.h. von Moduln, die keinen echten von Null verschiedenen Untermodul besitzen.

5.6 Bemerkung. Jeder Untermodul und jeder Faktormodul eines halbeinfachen Moduls ist halbeinfach.

5.7 Satz. Sei L eine halbeinfache Liealgebra. Dann ist jede Darstellung von L in einem ( endlichdimensionalen ) Vektorraum halbeinfach.

Der Satz wird durch folgendes Lemma auf einen Spezialfall zurückgeführt.

5.8 Lemma. Sei L eine Liealgebra. Folgende Aussagen sind äquivalent:

1) Jede Darstellung von L ist halbeinfach.

2) Sei μ : L → End(N) eine Darstellung von L im Vektorraum N und P ⊆ N ein Untermodul der Kodimension 1 mit μ(L)(N) ⊆ P. Dann gibt es einen Komplementärmodul P̃ : N = P ⊕ P̃.

Beweis des Lemmas. (2) ⟹ (1) :

Sei ρ : L → End(V) eine Darstellung von L und W ≠ 0 ein Untermodul von V. Wir definieren eine Darstellung μ von L im Vektorraum M = End(V). Sei x ∈ L

\begin{matrix}μ(x) = ad_{M}ρ(x) : End(V) & → End(V) \\ ϕ & ↦ ρ(x) ∘ ϕ − ϕ ∘ ρ(x) \end{matrix}

Es gilt also für x ∈ L, ϕ ∈ M, v ∈ V

μ(x)(ϕ)(v) = xϕ(v) − ϕ(xv)

Sei N ⊆ M der Untervektorraum derjenigen Endomorphismen ϕ ∈ M = End(V), die V in W abbilden und deren Einschränkung auf W die Multiplikation mit einem Element λ des Grunkörpers ist. Sei P = {ϕ ∈ N | ϕ|W = 0}. Dann ist P ein Untervektorraum von N der Kodimension 1. Es gilt

μ(L)(N) ⊆ P

Denn: Sei ϕ ∈ N und sei ϕ|W = λ · Id. Dann gilt für x ∈ L , v ∈ V , w ∈ W : μ(x)(ϕ)(v) = xϕ(v) − ϕ(xv) ∈ W und μ(x)(ϕ)(w) = xλw − λxw = 0. Also gibt es nach der Voraussetzung in (2) einen L-Komplementärmodul P̃ = kϕ der Dimension 1 von P in N: N = P ⊕ kϕ. Sei x ∈ L. Dann ist μ(x)(P̃) ⊆ P̃; also gibt es eine Konstante c ∈ k mit μ(x)(ϕ) = cϕ. Weil μ(x)(ϕ) in P enthalten ist folgt c = 0. Sei v ∈ V. Es gilt also

0 = μ(x)(ϕ)(v) = xϕ(v) − ϕ(xv)

ϕ : V → V ist deshalb ein Homomorphismus von L-Moduln. Da ϕ ∉ P, gilt ϕ|W = λ · Id mit λ ≠ 0. Für den Homomorphismus ψ = \frac{1}{λ}ϕ : V → V von L-Moduln gilt:

ψ|W = Id , ψ(V) ⊆ W

ψ ist eine Projektions abbildung auf W. Es gilt: W̃ = Kern(ψ) ist ein L-Untermodul mit

V = W ⊕ W̃ .
⊔ ⊓

Der Beweis des Satzes. Sei ρ : L → End(V) eine Darstellung der Liealgebra L und W ⊆ V ein Untermodul der Kodimension 1 mit ρ(L)(V) ⊆ W. ρ induziert eine Darstellung σ : L → End(W) im Untermodul W.

(1) Sei ρ(L)(W) = 0. Dann gilt für x, y ∈ L die Gleichung ρ(x)ρ(y) = 0. Wegen L = [L, L] folgt ρ(L) = ρ([L, L]) = 0. Also operiert L trivial auf V. Jeder Untervektorraum von V ist ein L-Untermodul. Wähle als Komplementärmodul W̃ von W in V einen Komplementärvektorraum.

(2) Sei W ein einfacher L-Modul und sei σ ≠ 0. Sei 𝔫 der Kern von σ. Wegen der Halbeinfachheit der adjungierten Darstellung (Satz 5.4) gibt es ein Ideal 𝔪 ≠ 0 mit L = 𝔪 ⊕ 𝔫. Die Einschränkung von σ auf 𝔪 ist dann eine treue Darstellung. Sei β_{σ}(x, y) = Spur(σ(x) ∘ σ(y)) die assozierte Bilinearform zur Darstellung σ. Dann ist die Einschränkung von β_{σ} auf die halbeinfache Liealgebra 𝔪 nicht ausgeartet (Bemerkung 5.2). Sei C ∈ U(L) das Casimirelement zur Darstellung σ und zum Ideal 𝔪. Da C im Zentrum der universellen Hülle U(L) liegt ist ρ(C) : V → V ein Homomorphismus von L-Moduln und es gilt ρ(C)(W) ⊆ W. Weil W ein einfacher L-Modul ist, ist der Homomorphismus ρ(C)|W : W → W ein Isomorphismus. Wegen ρ(L)(V) ⊆ W gilt auch ρ(C)(V) ⊆ W. Es folgt

V = W ⊕ Kern(ρ(C))

(3) Reduktion auf den Fall: ρ | W ist einfach. Induktion nach der Dimension von V. Sei U ein von 0 und W verschiedener einfacher L-Untermodul von W. Nach Induktionsvoraussetzung besitzt W/U einen Komplementärmodul in V/U. Sei T das Urbild in V. Es gilt dann:

T ∩ W = U , T + W = V , dim(T/U) = 1

Sei W̃ ein Komplementärmodul von U in T. Dann gilt

V = W ⊕ W̃ .
⊔ ⊓

Wir benötigen ein weiteres Lemma aus der linearen Algebra.

5.9 Lemma von Schur. Sei der Grundkörper k algebraisch abgeschlossen, V ≠ 0 ein endlicher Vektorraum über k und M ⊆ End(V) eine Teilmenge von Endomorphismen von V. Ein Unterraum W ⊆ V mit M(W) ⊆ W sei entweder ganz V oder der Nullraum. Sei R ⊆ End(V) der Ring der mit allen Elementen von M vertauschbaren Endomorphismen. k = k · id ⊆ R. Dann gilt

k = R

Beweis. Sei ϕ : V → V ein von Null verschiedener Endomorphismus aus R. Ist W = Kern(ϕ) oder W = Bild(ϕ) so gilt wegen der Vertauschbarkeitseigenschaft von ϕ: (M)(W) ⊆ W. Nach Voraussetzung gilt also Kern(ϕ) = 0 , Bild(ϕ) = V. ϕ ist ein Isomorphismus und es gilt ϕ^{−1} ∈ R. Sei α ∈ R und K = k[α] ⊆ R ⊆ End(V) der kleinste Unterring von R, der k und α enthält. Dann ist K ein Oberkörper endlichen Grades von k. Weil k algebraisch abgschlossen ist folgt K = k und α ∈ k.

5.10 Lemma. Sei L eine halbeinfache Liealgebra und ρ : L ↪ End(V) eine treue Darstellung von L.

1) Sei g ein Element von L und ρ(g) = S + N die Zerlegung von ρ(g) ∈ End(V) in den halbeinfachen Teil S ∈ End(V) und den nilpotenten Anteil N ∈ End(V). Dann sind S, N im Bild ρ(L) von L enthalten. Es gilt also:

g = g_{s} + n_{s} mit ρ(g_{s}) = S , ρ(g_{n}) = N.

Insbesondere sind g_{s} und g_{n} vertauschbar: [g_{s}, g_{n}] = 0.

2) Genau dann ist ρ(g) halbeinfach bzw. nilpotent, wenn ad_{L}(g) halbeinfach bzw. nilpotent ist.

Beweis. Wir können annehmen daß L in End(V) enthalten ist und daß der Grundkörper k algebraisch abgeschlossen ist. Weil jede halbeinfache Liealgebra eine direkte Summe einfacher Ideale ist (Satz 5.4) können wir auch annehmen, daß L eine einfache Liealgebra ist. Sei L ≠ 0, V ≠ 0.

Erster Fall . Sei V ein einfacher L-Modul. Sei {L̃ = x ∈ End(V) | [x, L] ⊆ L , Spur(x) = 0}. Dann ist L̃ ein L-Untermodul von End(V) bei der adjungierten Darstellung in End(V). Wegen [L, L] = L verschwindet die Spur aller Elemente aus L; also ist L ein L-Untermodul von L̃. Weil jeder L-Modul halbeinfach ist (Satz 5.7), gibt es einen L-Untermodul H ⊆ L̃ mit L̃ = L⊕H. Es gilt [L, H] ⊆ L∩H = 0. Daher ist H mit allen Endomorphismen aus L vertauschbar. Weil V ein einfacher L-Modul ist, folgt aus Schurs Lemma: H ⊆ k = k · id. Weil die Spur auf H verschwindet folgt: H = 0 , L̃ = L. Sei g ∈ L und g = S + N die Zerlegung von g in End(V) in den halbeinfachen und in den nilpotenten Anteil. Nach Lemma 4.30 des vorigen Paragraphen gilt ad_{End(V)}g = ad_{End(V)}S + ad_{End(V)}N die Zerlegung von ad_{End(V)}g : End(V) → End(V) in den halbeinfachen und nilpotenten Anteil. Es gibt dann ein Polynom P (t) ∈ k[t] mit P (ad_{End(V)}g) = ad_{End(V)}S , P (0) = 0. Wegen [g, L] = ad_{End(V)}(g)(L) ⊆ L gilt auch [S, L] = ad_{End(V)}S(L) = P (ad_{End(V)}g)(L) ⊆ L. Außerdem ist Sp(S) = Sp(g) = 0. Es folgt S ∈ L̃ = L und N = g − S ∈ L. Genau dann ist g halbeinfach, wenn N verschwindet. Da die adjungierte Darstellung eine treue Darstellung ist, ist das genau dann der Fall, wenn ad_{L}N, der nilpotente Anteil von ad_{L}g, verschwindet.

§6. Der Satz von Ado 75

Der allgemeine Fall. Weil wir vorausgesetzt haben, daß L eine einfache Liealgebra ist, ist die adjungierte Darstellung einfach. Wie wir im ersten Teil bewiesen haben, besitzt dann g eine Darstellung g = g_{s} + g_{n} mit g_{s}, g_{n} ∈ L , [g_{s}, g_{n}] = 0, so daß ad_{L}g = ad_{L}g_{s} + ad_{L}g_{n} die Zerlegung von ad_{L}g in den halbeinfachen und in den nilpotenten Anteil ist. Sei ρ : L → End(W) eine Darstellung. Ist ρ einfach so ist, wie wir bereits bewiesen haben, ρ(g) = ρ(g_{s}) + ρ(g_{n}) die Zerlegung von ρ(g) in den halbeinfachen bzw. in den nilpotenten Anteil in End(W). Da jede Darstellung der Liealgebra L halbeinfach ist, also in die direkte Summe einfacher Darstellungen

zerfällt, gilt diese Aussage auch für eine beliebige Darstellung.  ⊔ ⊓

5.11 Definition. Ein Element g der halbeinfachen Liealgebra L heißt halbeinfach bzw. nilpotent, wenn für jede Darstellung ρ : L → End(V) der Endomorphismus ρ(g) ∈ End(V) halbeinfach bzw. nilpotent ist.

Aus dem vorhergehenden Lemma folgt dann

5.12 Satz, Zerlegung von Elementen in den halbeinfachen und in den nilpotenten Anteil. Sei L eine halbeinfache Liealgebra und g ∈ L. Dann gibt es ein halbeinfaches Element g_{s} ∈ L und ein nilpotentes Element g_{n} ∈ L mit

g = g_{s} + g_{n} , [g_{s}, g_{n}] = 0.

g_{s}, g_{n} sind durch diese Bedingungen eindeutig bestimmt und heißen halbeinfacher bzw. nilpotenter Anteil des Elements g.

Mit Hilfe der Halbeinfachheit aller Darstellungen einer halbeinfachen Liealgebra kann ein weiterer wichtiger Satz über Liealgebren gefolgert werden. ( Siehe Bourbaki, Groupes et algèbres de Lie. Chap I )

5.13 Satz, Die Levizerlegung einer Liealgebra. Sei L eine beliebige Liealgebra L und sei 𝔯 das Radikal von L. Dann gibt es eine halbeinfache Lieunteralgebra H von L mit

L = 𝔯 ⊕ H

Offenbar ist dann H isomorph zu L/𝔯. L heißt eine Levialgebra von L, die angegebene Zerlegung von L die Levizerlegung.

Offenbar ist dann H isomorph zu L/𝔯. L heißt eine Levialgebra von L, die angegebene Zerlegung von L die Levizerlegung.

6. Der Satz von Ado

In diesem Abschitt soll gezeigt werden, daß jede endlichdimensionale Liealgebra über einem Grundkörper k der Charakteristik 0 isomorph ist zu einer Unterliealgebra einer Matrizenalgebra M_{n}(k).

Wir wollen wieder annehmen, daß der Grundkörper k die Charakteristik 0 hat. Alle betrachteten Liealgebren haben endliche Dimension über k.

Zuerst eine Folgerung aus dem Satz von Poincaré-Birkhoff-Witt

6.1 Lemma. Sei L eine Liealgebra und U(L) die universelle Hülle von L. Dann ist U(L) links nöthersch. Das bedeutet, daß jedes Linksideal J ⊆ U(L) von U(L) endlich viele Erzeugende x_{ν} ∈ J besitzt.

J = U(L)x_{1} + … + U(L)x_{n}

Ebenso ist U(L) rechts nöthersch.

Sei 𝔞 ⊆ L eine Unterliealgebra der Liealgebra L und sei ρ : 𝔞 → End(V) eine Darstellung von 𝔞. Wir betrachten Fortsetzungen ρ′ : L → End(W) von ρ in einem Vektorraum W, der V enthalt:

ρ′|𝔞 = ρ.

Ist 𝔞 ein auflösbares Ideal von L, so kann es eine endlichdimensionale Fortsetzung von ρ in einen endlichdimensionalen Vektorraum V nur geben, wenn [L, 𝔞] im größten nilpotenten Ideal der Darstellung ρ liegt. (siehe Auflösbare und nilpotente Liealgebren: Satz 4.18)

6.2 Satz, Fortsetzung einer Darstellung eines Ideals auf die ganze Liealgebra. Sei L eine Liealgebra, 𝔞 ein Ideal von L und H eine Unterliealgebra von L mit

L = 𝔞 ⊕ H

Gegeben sei eine Darstellung ρ : 𝔞 → End(V) in einem endlichdimensionalen Vektorraum V. [H, 𝔞] sei im größten nilpotenten Ideal von ρ enthalten. Dann gilt

1) Es gibt eine Fortsetzung ρ′ : L → W von ρ in einem endlichdimensionalen Vektorraum W ⊇ V, deren größtes nilpotentes Ideal das größte nilpotente Ideal von ρ umfaßt.

2) Sei für alle Elemente h ∈ H die Restriktion ad_{L}h|𝔞 : 𝔞 → 𝔞 von ad_{L}h auf 𝔞 nilpotent. Dann kann ρ′ so gewählt werden, daß H im größten nilpotenten Ideal von ρ′ enthalten ist.

Beweis. Zu einer Darstellung π : L → End(V) einer Liealgebra L im endlichdimensionalen Vektorraum V erklärt man die duale Darstellung im dualen Vektorraum V^{⋆} = Hom_{k}(V, k) :

\begin{aligned}π̂ : L & → End(V^{⋆}) \\ π̂(ϕ)(x) & = −ϕ(π(x)) ; ϕ ∈ Hom_{k}(V, k), x ∈ V \end{aligned}

Es gilt: ˆ π̂ = π.

Mit Hilfe von Dualisierung sieht man, daß es genügt, die duale Aussage des Satzes zu beweisen:

Es gibt eine Darstellung ρ′ : L → End(W) in einem endlichdimensionalen Vektorraum W und einen surjektiven Homomorphismus

W ⟶ V

von Moduln über der Liealgebra 𝔞. Die Darstellung ρ′ hat die in (1) bzw. (2) geforderten Eigenschaften.

Wir betrachten die universellen Hüllen der Liealgebren 𝔞 ⊆ L: U(𝔞) ⊆ U(L). Da 𝔞 ein Ideal von L ist gilt für x ∈ L ⊆ U(L) und y ∈ U(𝔞), daß xy − yx ∈ U(𝔞). Wir konstruieren eine Darstellung π von L = 𝔞 ⊕ H im Vektorraum U(𝔞) : Sei y ∈ U(𝔞).

π(x)(y) = \{ \begin{matrix}xy & x ∈ 𝔞 \\ xy − yx & x ∈ H \\ π(a)(y) + π(h)(y) & x = a + h für a ∈ 𝔞 , h ∈ H \end{matrix}

Man rechnet leicht nach, daß π eine Darstellung der Liealgebra L ist.

Wir betrachten eine Jordan-Hölder Folge des 𝔞-Moduls V.

0 = V_{0} ⊆ V_{1} ⊆ … ⊆ V_{d} = V

Die Darstellung ρ wird fortgesetzt zu einem Ringhomomorphismus ρ : U(𝔞) → End(V). Sei I ⊆ U(𝔞) das Kernideal dieses Homomorphismus. U(𝔞) operiert auch auf allen Faktormoduln der Jordan-Hölder Reihe. Sei

J = ⋂_{ν} Kern(U(𝔞) → End(V_{ν+1}/V_{ν}))

I, J sind zweiseitige Ideale von U(𝔞). Es gilt

J^{d} ⊆ I ⊆ J

J ∩ 𝔞 = 𝔫 ist das größte bzgl. ρ nilpotente Ideal der Liealgebra 𝔞, enthält also nach Voraussetzung [H, 𝔞]. Es folgt:

\begin{aligned}π(H)U(𝔞) & ⊆ J \\ π(H)J^{d} & ⊆ J^{d} \\ π(L)J^{d} & ⊆ J^{d} \end{aligned}

Aus Lemma 6.1 folgt, daß der Linksmodul J^{ν}/J^{ν+1} ein endlicher U(𝔞)-Modul ist. I ist in J enthalten und annuliert deshalb J^{ν}/J^{ν+1}. J^{ν}/J^{ν+1} ist also endlicher U(𝔞)/I-Modul. Da U(𝔞)/I isomorph ist zu einem Untervektorraum von End(V), ist J^{ν}/J^{ν+1} ein endlicher Vektorraum für ν = 0, …, d − 1; also ist auch M = U(𝔞)/J^{d} ein endlicher Vektorraum. π induziert eine Darstellung von L im endlichen Vektorraum M. Sei v ein Vektor aus V. Betrachte die Abbildung

\begin{matrix}U(𝔞) & → V \\ x & ↦ ρ(x)v \end{matrix}

Diese lineare Abbildung ist ein Homomorphismus von Darstellungsmoduln über 𝔞, wenn man die Darstellung π|𝔞 von 𝔞 auf U(𝔞) betrachtet. Wegen J^{d} ⊆ I induziert dieser Homomorphismus einen Homomorphismus von Darstellungsmoduln über der Liealgebra 𝔞 :

ϕ_{v} : M ⟶ V.

v ist im Bild von ϕ_{v} enthalten. Sei v_{1}, …, v_{r} eine Basis von V. Dann ist

ϕ_{v_{1}} + … + ϕ_{v_{r}} : W = M ⊕ … ⊕ M ⟶ V

ein surjektiver Homomorphismus von Moduln über 𝔞. Die Darstellung ρ′ von L in W ist die gesuchte Darstellung der Liealgebra L. Nach Konstruktion ist ρ′(𝔫) nilpotent.

Der Teil (2) des Satzes:

Nach Konstruktion ist ρ′(h) für h ∈ H nilpotent. Betrachte das Bild G = \overline{𝔫} + \overline{H} von 𝔫 + H in End(W). Dann ist nach dem Satz von Engel \overline{𝔫} ein nilpotentes Ideal von G und \overline{H} eine nilpotente Unteralgebra von G. Deshalb ist G auflösbar. Aus Folgerung 19 des Paragraphen über auflösbare und nilpotente Algebren folgt dann, daß G bzgl. der Darstellung in W nilpotent ist. Damit ist 𝔫 + H in End(W) ein nilpotentes Ideal von L bzgl. der Darstellung ρ′, also im größten nilpotenten Ideal

dieser Darstellung enthalten.  ⊔ ⊓

6.3 Folgerung. Sei L eine Liealgebra, 𝔯 ihr Radikal, 𝔫 das größte nilpotente Ideal von L und 𝔞 ein nilpotentes Ideal von L.

ρ : 𝔞 → End(V)

sei eine nilpotente Darstellung von 𝔞 in einem endlichdimensionalen Vektorraum V. Dann gibt es einen endlichdimensionalen Vektorraum W ⊇ V und eine Darstellung von L in W, die die Darsellung ρ von 𝔞 fortsetzt und für die 𝔫 ein nilpotentes Ideal ist.

Beweis. Sei 𝔟 ein Ideal der Kodimension 1 in der Lieunteralgebra 𝔠 von L. Sei c ∈ 𝔠 ein Element, das nicht in 𝔟 enthalten ist. Dann ist H = kc eine Lieunteralgebra von 𝔠 und es gilt 𝔠 = 𝔟 ⊕ H.

Es gilt

𝔞 ⊆ 𝔫 ⊆ 𝔯 ⊆ L.

Durch iterierte Anwendung des Fortsetzungssatzes 6.2 erhält man zuerst eine Fortsetzung der Darstellung ρ zu einer Darstellung ρ′ von 𝔫 in einem größeren Vektorraum, so daß ρ′(𝔫) nilpotent ist. Wegen [L, 𝔯] ⊆ 𝔫 (Satz 4.18) in Auflösbare Liealgebren kann man diese Darstellung schrittweise auf 𝔯 fortsetzen. Sei H ⊆ L eine Leviunteralgebra von L. (Satz 5.13 des vorigen Paragraphen):

L = 𝔯 ⊕ H

Da [L, 𝔯] in 𝔫 enthalten ist und weil ρ(𝔫) nilpotent ist, kann die Darstellung von 𝔯 zu einer endlichdimensionalen Darstelung von L fortgesetzt werden. ⊔ ⊓

Sei Z das Zentrum der Liealgebra L. Weil Z kommutativ ist besitzt Z eine treue nilpotente Darstellung. Wegen Folgerung 3 kann diese Darstellung zu einer endlichdimensionalen Darstellung ρ : L → End(V) fortgesetzt werden. Nach Konstruktion ist ρ auf Z treu. Die adjungierte Darstellung ad_{L} : L → End(L) hat als Kern das Zentrum Z. Dann ist die direkte Summe dieser Darstellungen

L ⟶ End(V ⊕ L)

injektiv.

6.4 Theorem. Sei L eine endlichdimensionale Liealgebra über einem Grundß-kör-per k der Charakteristik 0. Dann besitzt L eine treue Darstellung

Ł ↪ End(V).

Man kann sogar erreichen, daß das größte nilpotente Ideal von L im größten nilpotenten Ideal dieser Darstellung enthalten ist. Insbesondere ist L isomorph zu einer Unterliealgebra eines Matrizenringes M_{n}(k).

Chapter V. Klassifikation halbeinfacher Liealgebren

Bibliographie

[1] N.Bourbaki Groupes et algèbres de Lie. Chap. 4,5,6

[2] N.Bourbaki Groupes et algèbres de Lie. Chap. 7 et 8

[3] J-P.Serre Algèbres de Lie semisimples complexes

1. Wurzeln und Wurzelräume

Der Grundkörper k habe im folgenden die Charakteristik Null. Sei V ein endlichdimensionaler Vektorraum.

1.1 Definition. Sei n = dim(V) und sei s : V ⟶ V eine lineare Abbildung. s heißt Spiegelung, wenn gilt

1) s^{2} = id, s ≠ id (=Identität)

2) dim(Kern(s − 1)) = n − 1

Es folgt, daß 1 Eigenwert der Vielfachheit n − 1 der Spiegelung s ist. Es gibt dann genau einen weiteren Eigenwert, der den Wert -1 hat. Sei W = Kern(s − 1) und α ein Eigenvektor zum Eigenwert -1. Es gilt

\begin{aligned}V & = W ⊕ kα \\ s(w + λα) & = w − λα \end{aligned}

Sei (x, y) eine symmetrische nicht ausgeartete Bilinearform auf V. Diese Bilinearform sei invariant gegen die Spiegelung s, also s orthogonal bzgl. (x, y):

(sx, sy) = (x, y)

Dann ist W = (α)^{⊥}. Es gilt

s(v) = v − 2\frac{(v, α)}{(α, α)}α

Im allgemeinen gibt es genau ein Element α̌ im Dualraum V^{⋆} von V mit α̌(α) =< α̌, α >= 2 und

s(x) = x− < α̌, x > α

1.2 Lemma (Maschke). Die endliche Gruppe H operiere auf dem (endlichdimensionalen ) Vektorraum V. Sei U ⊆ V ein H–invarianter Unterraum. Dann gibt es einen H–invarianter Unterraum W ⊆ V mit V = U ⊕ W, einen zu U komplementären H–Untermodul.

1.3 Satz. Seien s, t Elemente einer endlichen Gruppe G, die auf dem Vektorraum V operiert und sei α ≠ 0 ein Element aus V. s, t seien Spiegelungen und es gelte s(α) = −α = t(α). Dann ist s = t. Wir setzen s_{α} = s = t für diese eindeutig durch α bestimmte Spiegelung aus G (falls sie existiert).

Beweis. Sei H die von s, t erzeugte Untergruppe. Sei W ein Komplementärmodul zum H–Untermodul kα = U :

V = U ⊕ W
Es folgt: s|W = t|W = id, damit s = t  ⊔ ⊓

1.4 Satz. Sei G eine endliche Gruppe von Automorphismen des Vektorraums V. V sei G–irreduzibel. Sei s ∈ G eine Spiegelung.

1) Sei A : V → V ein G–Endomorphismus. Dann gibt es ein Element λ ∈ k mit A = λ · id.

2) Sei β(x, y) eine G–invariante nicht unbedingt symmetrische Bilinearform auf V, die nicht identisch verschwindet. Dann ist β nicht ausgeartet. β ist durch diese Vorgaben eindeutig bis auf einen Faktor bestimmt; insbesondere gilt β(x, y) = λβ(y, x) mit λ^{2} = 1. β ist entweder symmetrisch oder antisymmetrisch.

Beweis. (1): Sei 0 ≠ α ∈ V mit s(α) = −α und sei U = kα. Wegen sA = As gilt A(U) ⊆ U. Deshalb ist α ein Eigenvektor von A zum Eigenwert λ ∈ k. Der Kern K von A−λ·id ist G–invariant und von Null verschieden. Da V irreduzibler G–Modul ist, folgt K = V, damit A = λ · id.

(2) V^{⊥} ist G–invariant. Daher gilt V^{⊥} = 0 oder V^{⊥} = V. Da β ≠ 0 ist gilt V^{⊥} = 0. Sei B eine zweite Bilinearform mit den vorgegebenen Eigenschaften. Da β nicht ausgeartet ist gibt es eine eindeutig bestmmte lineare Abbildung A : V → V mit B(x, y) = β(Ax, y). wegen der G–Invarianz von B und β folgt für σ ∈ G:

β(Ax, y) = β(Aσx, σy) = β(σ^{−1}Aσx, y)

Weil β nicht ausgeartet ist, folgt σ^{−1}Aσ = A ∀σ ∈ G. Nach (1) gilt also A = λ·id, damit B = λβ ⊔ ⊓

Sei V wieder ein vorgegebener (endlichdimensionaler ) Vektorraum und R ⊆ V eine endliche Teilmenge, die V erzeugt. Sei

0 ∉ R

Die Gruppe der Automorphismen σ von V mit σ(R) = R ist endlich. Nach 1.3 gibt es zu einem Element α ∈ R höchstens eine Spiegelung s = s_{α} mit

sα = −α, s(R) = R

1.5 Definition. (V, R) heißt Wurzelraum oder Wurzelsystem, wenn zusätzlich gilt

1) Zu jedem Element α ∈ R gibt es eine Spiegelung s_{α} mit s_{α}^{2}α = −α und mit s_{α}R = R.

2) Wir schreiben

s_{α}(x) = x− < α̌, x > α

mit der eindeutig bestimmten Kowurzel α̌ ∈ V^{⋆} zur Wurzel α Dann ist < α̌, α >= 2. Es gelte: < α̌, β >∈ ℤ, ∀ α, β ∈ R R ist die Menge der Wurzeln des Wurzelraumes.

§1. Wurzeln und Wurzelräume 81

1.6 Definition. Die durch die Spiegelungen s_{α}, α ∈ R erzeugte endliche Gruppe heißt die Weylgruppe des Wurzelsystems. Schreibweise W = W (R).

1.7 Lemma.

\begin{aligned}β ∈ R & ⟹ −β ∈ R \\ β ∈ R, λβ ∈ R & ⟹ λ ∈ \{±1, ±\frac{1}{2}\} \end{aligned}

Denn: Sei β = λα, α, β ∈ R. Dann gilt < α̌, β >= λ < α̌, α >= 2λ ∈ ℤ.

Vertausche α, β Dann folgt 2/λ ∈ ℤ  ⊔ ⊓

1.8 Definition. Das Wurzelsystem (V, R) heißt reduziert, wenn gilt

α ∈ R, λα ∈ R ⟹ λ = ±1

Im folgenden sind alle Wurzelsysteme reduziert.

1.9 Bemerkung. Die Kowurzeln α̌ sind die Wurzeln eines Wurzelsystems (V^{⋆}, Ř) im Dualraum V^{⋆}. Die Weylgruppen der beiden Wurzelsysteme sind isomorph.

\begin{matrix}W (R) ∈ σ & ↦ σ^{⋆} ∈ W (Ř) \\ σ^{⋆}(ψ) & = ψ ∘ σ^{−1} \end{matrix}

Grundkörpererweiterung

Gegeben sei ein Erweiterung k ⊆ K des Grunkörpers und ein Wurzelsytem (V, R) über dem Grundkörper k. Sei V_{K} = V ⊗_{k} K. Wir identifizieren R mit dem Bild von R in V_{K}. R = R_{K} = R ⊗ K. Dann ist (V_{K}, R) = (V, k) ⊗ K ein Wurzelsystem über dem Grundkörper K. Dann ist V_{0} = ∑_{α∈R} ℚα ⊆ V ein endlich erzeugter Vektorraum über ℚ der R enthält. Man sieht leicht, daß (V_{0}, R) ein Wurzelsystem über dem Grundkörper ℚ ist. Wir betrachten die natürliche surjektive Abbildung

V_{0} ⊗_{ℚ} k ⟶ V.

Übungsaufgabe : Seien α_{1}, …, α_{r} ∈ R linear unabhängig über ℚ. Dann sind diese Wurzeln auch linear unabhängig über k. Oder: dim_{ℚ}(V_{0}) = dim_{k}(V). Dazu zeige, daß im Falle des Grundkörpers ℚ oder auch ℝ die Kowurzeln den dualen Vektorraum erzeugen. Man beschreibe die Kowurzeln mit Hilfe einer positiv definiten W –invarianten Bilinearform. Es folgt, daß diese k–lineare Abbildung auch injektiv ist.

1.10 Lemma. (V, R) entsteht aus (V_{0}, R) durch Grundkörpererweiterung:

(V_{0}, R) ⊗_{ℚ} k = (V, R)

Die Vorgabe von (V_{0}, R) ist äquivalent zur Vorgabe von (V, R). Man erweitere (V_{0}, R) mit dem Körper ℝ der reellen Zahlen. Man kann (V_{0}, R) durch diese Erweiterung ersetzen.

(V_{0}, R) ≅ (V′_{0}, R′) ⟺ (V_{0} ⊗_{ℚ} ℝ, R) ≅ (V′_{0} ⊗_{ℚ} ℝ, R′)

Ohne Beschränkung der Allgemeinheit genügt es, Wurzelsysteme über dem Grundkörper ℝ zu untersuchen bzw. zu klassifizieren. Von nun an sei

k = ℝ

der vorgegebene Grundkörper. Sei(V, R) ein Wurzelsystem und W = W (R) die zugehörige Weylgruppe. Dann gibt es auf V eine positiv definite symmetrische Bilinearform (−, −), die W –invariant ist:

(σ(x), σ(y)) = (x, y) ∀σ ∈ W, ∀x, y ∈ V.

Wir wählen ein solches Skalarprodukt aus. Sei α ∈ R eine Wurzel. Dann ist die Spiegelung s = s_{α} eine orthogonale lineare Abbildung. Es gilt:

\begin{aligned}s_{α}(x) & = x − 2\frac{(α, x)}{(α, α)}α \\ < α̌, x > & = 2\frac{(α, x)}{(α, α)} \end{aligned}

Sei U ⊆ V ein W –invarianter Untervektorraum; U ≠ 0 , U ≠ V. Dann ist auch U^{⊥} = Ũ W –invariant. Sei x ∈ U und α ∈ R

⟹ s_{α}x = x − 2\frac{(α, x)}{(α, α)}α ∈ U ⟹ (α, x)α ∈ Ũ

Es gibt zwei Möglichkeiten

1) α⊥U also α ∈ Ũ

2) ∃x_{0} ∈ V mit (x_{0}, α) ≠ 0 Damit gilt α ∈ U.

Man sieht leicht

R = R ∩ U ∪^{·} R ∩ Ũ

Sei R_{U} = R ∩ U, RŨ = R ∩ Ũ. Dann sind (U, R_{U}) und (Ũ, RŨ) Wurzelsysteme.

(U, R) = (U, R_{U}) ⊕ (Ũ, RŨ)

Wenn es keinen solchen Unterraum U ≠ 0, V gibt, es also keine solche orthogonale Zerlegung des Wurzelsystems gibt, so heißt das Wurzelsystem (V, R) irreduzibel. Äquivalent: Es gibt keine Zerlegung

R = R_{1} ∪^{.} R_{2} mit R_{1}⊥R_{2}

Die letzte Bedingung läßt sich ohne invariantes Skalarprodukt beschreiben:

< α̌, β >=< β̌, α >= 0 ∀α ∈ R_{1} , ∀β ∈ R_{2}

Jedes Wurzelsystem kann eindeutig in eine “direkte Summe“ von irreduziblen Wurzelsystemen zerlegt werden. Es genügt solche irreduziblen Wurzelsysteme zu untersuchen. Aus dem früher bewiesenen Satz 1.4 folgt:

1.11 Satz. Sei das Wurzelsystem (V, R) irreduzibel und (x, y) ein vorgegebenes gegen die Weylgruppe W invariantes Skalarprodukt. Sei β(x, y) eine weitere nicht verschwindende symmetrische W –invariante Bilinearform. Dann gibt es ein Element 0 ≠ λ ∈ ℝ mit β(x, y) = λ(x, y). Falls β positiv definit ist, gilt λ > 0. Das invariante Skalarprodukt (−, −) ist also in diesem irreduziblen Falle eindeutig bis auf einen positiven Faktor bestimmt.

Seien α, β ∈ R zwei Wurzeln. Setze

n(β, α) =< α̌, β >= 2\frac{(α, β)}{(α, α)} ∈ ℤ

Sei |α| = \sqrt{(α, α)} die Länge der Wurzel α und sei ϕ der Winkel zwischen den beiden Wurzeln.

\begin{aligned}n(β, α) & = 2\frac{|β|}{|α|}cos(ϕ) \\ n(β, α)n(α, β) & = 4(cos(ϕ))^{2} \end{aligned}

Es folgt: 4(cos(ϕ))^{2} = 0, 1, 2, 3, 4. Der letzte Fall tritt nur ein, wenn α und β proportional sind, d.h. wenn β = ±α—das Wurzelsystem wurde reduziert angenommen! Wenn β ≠ ±α, so gibt es nur die Möglichkeiten:

n(α, β) = 0, ±1, ±2, ±3

Übungsaufgabe : Tabeliere die Verschiedenen Kombinationsmöglichkeiten für die Zahlen n(α, β), n(β, α), |α|, |β|, ϕ.

1.12 Lemma. Sei α, β ∈ R β ≠ ±α Dann folgt aus n(β, α) > 0 daß auch α−β ∈ R eine Wurzel ist. Die Bedingung ist äquivalent zu (α, β) > 0; insbesondere ist sie symmetrisch in α, β.

Beweis. (α, β) = (β, α) > 0 ⟹

1 ≤ n(α, β), n(β, α) ≤ 3 ⟹ n(α, β) = 1 oder n(β, α) = 1,

n(β, α) = 1 ⟹ s_{α}(β) = β − n(β, α)α = β − α ∈ R ⟹ α − β = −(β − α) ∈ R.

Entsprechend:

n(α, β) = 1 ⟹ … ⟹ α − β ∈ R
⊔ ⊓

Wurzelbasen; einfache Wurzelsysteme

1.13 Definition. Sei (V, R) ein Wurzelraum und S ⊆ R eine Teilmenge von Wurzeln. S heißt Wurzelbasis oder einfaches Wurzelsystem, wenn gilt:

(a) Die Elemente aus S sind linear unabhängig.

(b) β ∈ R ⟹ β = ∑_{α∈S} n_{α}α mit n_{α} ∈ ℤ. Entweder n_{α} ≥ 0 ∀α ∈ S oder n_{α} ≤ 0 ∀α ∈ S Da die Wurzeln den Vektorraum V erzeugen ist insbesondere S eine Basis des

Vektorraumes V.

Wenn eine solche Wurzelbasis S vorgegeben ist, heißen die Wurzeln aus S einfache ( positive ) Wurzeln. Eine Wurzel β heißt positiv–bzgl. S–, wenn β = ∑_{α∈S} n_{α}α mit n_{α} ≥ 0 ∀α ∈ S. Schreibweise α > 0. Sonst negativ: β < 0.

\begin{aligned}R^{+} & = \{β > 0\} \\ R^{−} & = \{β < 0\} = −R^{+} \end{aligned}

1.14 Lemma. Sei V ein Vektorraum, (−, −) ein ( positives ) Skalarprodukt auf V und t : V → ℝ eine Linearfunktion. Gegeben seien Vektoren α_{1}, …, α_{n} ∈ V. Es gelte

t(α_{ν}) > 0, (α_{i}, α_{j}) ≤ 0 für i ≠ j

Dann sind die Vektoren α_{1}, …, α_{n} linear unabhängig.

Beweis. Sei

v = λ_{1}α_{ν_{1}} + … + λ_{r}α_{ν_{r}} = c_{1}α_{μ_{1}} + … + c_{s}α_{μ_{s}}

mit λ_{i}, c_{j} ≥ 0 ν_{i} ≠ μ_{j} Es folgt 0 ≤ (v, v) = ∑_{i,j} λ_{i}c_{j}(α_{ν_{i}}, α_{μ_{j}}) ≤ 0 und daraus (v, v) = 0, also v = 0. Sei r ≥ 1. 0 = v = λ_{1}α_{ν_{1}} + … + λ_{r}α_{ν_{r}}. Wir wenden auf diese Gleichung t an und erhalten

0 = t(v) = ∑_{i} λ_{i}t(α_{ν_{i}}) mit t(α_{ν_{i}}) > 0, λ_{i} ≥ 0

Es folgt λ_{1} = … = λ_{r} = 0. Ebenso zeigt man c_{1} = … = c_{s} = 0.  ⊔ ⊓

Sei nun t : V → V eine vorgegebene Linearform mit t(α) ≠ 0 ∀α ∈ R. Sei

\begin{aligned}R_{t}^{+} & = \{α ∈ R|t(α) > 0\} \\ R_{t}^{−} & = \{α ∈ R|t(α) < 0\} = −R_{t}^{+} \end{aligned}

Dann ist R = R_{t}^{+} ∪^{·} R_{t}^{−}. Eine Wurzel α ∈ R_{t}^{+} heißt unzerlegbar , wenn es keine Wurzeln γ, δ ∈ R_{t}^{+} gibt mit α = γ + δ, sonst zerlegbar. Sei

S_{t} = {α ∈ R_{t}^{+}|α unzerlegbar }.

1.15 Satz, Existenz einer Wurzelbasis des Wurzelsystems (V, R).. Unter den gemachten Voraussetzungen für die Linearform t ist S_{t} eine Wurzelbasis.

Bemerkung: Wir beweisen mit: α, β ∈ S_{t} mit α ≠ β ⟹ (α, β) ≤ 0.

1.16 Bemerkung. Man sieht leicht, daß jede Wurzelbasis von der Form S_{t} ist.

Beweis des Satzes. Weil R endlich ist, gibt es keine unendliche Kette

β_{1}, β_{2}, β_{3}, … ∈ R_{t}^{+}

mit t(β_{1} > t(β_{2}) > t(β_{3}) > . . .. Es folgt: Jede positive Wurzel läßt sich darstellen als

α = α_{ν_{1}} + α_{ν_{2}} + … + α_{ν_{m}} α_{ν_{j}} ∈ S_{t}

Dabei sind nicht alle α_{ν_{j}} verschieden. Sei S_{t} = {α_{1}, …, α_{r}}. Dann ist

α = n_{1}α_{1} + n_{2}α_{2} + … + n_{r}α_{r} n_{j} ∈ ℤ, n_{j} ≥ 0

Für jede negative Wurzel β(= −α, α > 0) gilt:

β = n_{1}α_{1} + n_{2}α_{2} + … + n_{r}α_{r} n_{j} ∈ ℤ, n_{j} ≤ 0

Behauptung (α_{i}, α_{j}) ≤ 0. Sonst wäre nach Lemma 1.12 δ = α_{i} − α_{j} ∈ R. Entweder: t(δ) > 0 ⇒ α_{i} = δ + α_{j} δ, α_{j} ∈ R_{t}^{+}; also wäre α_{i} zerlegbar. Oder: t(δ) < 0. Dann wäre α_{j} = −(δ) + α_{i} zerlegbar. Widerspruch! Aus Lemma 1.14 folgt dann, daß die Elemente α_{1}, …, α_{r}linear unabhängig sind. Damit ist S_{t} = {α_{1}, …, α_{r}} eine

Wurzelbasis.  ⊔ ⊓

Eigenschaften von Wurzelbasen

Sei S eine Wurzelbasis des Wurzelraumes (V, R). Wir wählen eine Linearfunktion t → ℝ mit S = S_{t}. Dann ist S die Menge der unzerlegbaren positiven Wurzeln. Sei S = {α_{1}, …, α_{r}}. Dann gilt nach Satz 1.15 (α_{i}, α_{j}) ≤ 0 ∀i ≠ j. Sei α > 0 eine beliebige positive Wurzel. Dann kann nicht für alle j gelten (α, α_{j}) ≤ 0. Sonst wäre α nach Lemma 1.14 linear unhabängig von der Basis S. Sei also α eine nicht in S enthaltene Wurzel und sei (α, α_{i}) > 0 für einen Index i. Es folgt nach Lemma 1.12 β = α−α_{i} ∈ R. Da α sich eindeutig als Linearkombination der α_{j} mit ganzzahligen nichtnegativen Koeffizienten darstellen läßt und da sich β als Linearkombination der α_{j} mit ganzzahligen Koeffizienten darstellen läßt, die enweder alle nicht negativ oder alle nicht positiv sind, folgt β > 0, α = β + α_{i}. Klar ist: t(β) < t(α). Durch Induktion erhält man:

1.17 Satz. Gegeben sei eine Wurzelbasis S = {α_{1}, …, α_{r}} und sei α ∈ S^{+} eine positive Wurzel. Dann gibt es eine Darstellung

α = α_{ν_{1}} + α_{ν_{2}} + … + α_{ν_{m}}

so daß für alle 1 ≤ s ≤ m die Linearombination α_{ν_{1}} + α_{ν_{2}} + … + α_{ν_{s}} eine, damit positive, Wurzel ist.

1.18 Theorem. Sei (V, S) ein Wurzelsystem mit der Wurzelbasis S und W = W (R) die Weylgruppe des Systems. W operiert auch auf dem Dualraum V^{⋆}: σ ∈ W , ψ : V → ℝ ⇒ σ(ψ) = ψ ∘ σ^{−1} : V → ℝ

1) Sei t ∈ V^{⋆}. Dann gibt es ein Element σ ∈ W mit < σ(t), α >≥ 0 ∀α ∈ S

2) Sei S′ eine weitere Wurzelbasis. dann gibt es ein Element σ der Weylgruppe mit σ(S′) = S

3) Sei β eine beliebige Wurzel. Dann gibt es ein Element σ der Weylgruppe mit σ(β) ∈ S.

4) Die Weylgruppe W ist gleich der durch die Spiegelungen {s_{α}|α ∈ S} erzeugten Untergruppe W_{S}.

1.19 Bemerkung zu Punkt (2). Man kann zeigen, daß das Element σ ∈ W eindeutig bestimmt ist.

Zuerst beweisen wir einen Hilfssatz.

1.20 Lemma. Sei S eine Wurzelbasis.

1) α ∈ S ⟹ s_{α}(R^{+} \ {α}) = R^{+} \ {α}

2) (Folge von (1) ) Sei ρ = \frac{1}{2} ∑_{α∈R^{+}} α Dann gilt s_{α}(ρ) = ρ − α

Beweis des Lemmas. Sei β = ∑_{γ∈S} m_{γ}γ ∈ R^{+} \ {α}. Dann sind alle Koeffizienten m_{γ} nicht negativ. Wegen β ≠ α gibt es ein γ ∈ S, verschieden von α mit m_{γ} > 0. Denn aus β = mα folgt m = ±1. s_{α}(β) = β − n(β, α)α ∈ R mit n(β, α) ∈ ℤ. Der Koeffizient von γ in der Darstellung von s_{α}(β) durch die Wurzelbasis ist wieder m_{γ} > 0. Also sind alle Koeffizienten nicht negativ. Es folgt: s_{α}(β) ∈ R^{+} und wegen m_{γ} ≠ 0 auch s_{α}(β) ≠ α ⊔ ⊓

Beweis des Theorems. Zur Bezeichnung: ψ ∈ V^{⋆}, x ∈ V; dann ist < ψ, x >= ψ(x). Beweis von (1) Sei ρ wie im vorigen Lemma 1.20. Wir wählen σ ∈ W, so daß < σ(t), ρ > maximal wird. Sei α ∈ S.

\begin{aligned}< σ(t), ρ >≥< s_{α}(σ(t)), ρ > & =< σ(t), s_{α}(ρ) >=< σ(t), ρ − α > \\ & =< σ(t), ρ) > − < σ(t), α > \end{aligned}

Es folgt < σ(t), α >≥ 0 Sei W_{S} die von den Spiegelungen {s_{α}|α ∈ S} erzeugte Untergruppe von W. Offenbar gelingt der Beweis von (1) auch mit der Gruppe W_{S} an Stelle von W . Beweis von (2). Wähle t so, daß S′ = S_{t}. Insbesondere ist dann t(α) ≠ 0 ∀α ∈ R. Wegen (1) gibt es ein Element σ ∈ W_{S}, so daß < σ(t), α >=< t, σ^{−1}(α) >≥ 0 ∀α ∈ S. Es gilt dann sogar ⋯ > 0. Es folgt:

< σ(t), α >> 0 ∀α > 0 (bzgl. S)
⟹< t, σ^{−1}(α) >> 0 ∀α > 0
⟹ σ^{−1}(R^{+}) ⊆ R_{t}^{+}
⟹ σ^{−1}(R^{+}) = R_{t}^{+} Anzahlschluß; Anzahl(R^{+}) = \frac{1}{2}Anzahl(R) etc.

Weil S bzw. S_{t} die Mengen der unzerlegbaren Elemente von R^{+} bzw. R_{t}^{+} sind, folgt σ^{−1}(S) = S_{t}. Beweis von (3) Für 0 ≠ v ∈ V bezeichnen wir mit H_{v} = {ψ ∈ V^{⋆}| < ψ, v >= 0} die Orthogonalhyperebene. Es gilt: H_{v} = H_{w} ⟺ w = λv. Sei β eine Wurzel und H = H_{β}. Es folgt: H_{α} ≠ H ∀α ∈ R mit α ≠ ±β Weil es nur endlich viele Wurzeln gibt, gilt H ⊂ ∪_{α∈R,α≠±β}H_{α}. Deshalb gibt es eine nicht verschwindende Linearfunktion t_{0} ∈ V^{⋆} mit

< t_{0}, β >= 0 , < t_{0}, α >≠ 0∀α ∈ R, α ≠ ±β

Sei t_{1} irgend eine weitere Linearfunktion mit t_{1}(β) > 0 und δ > 0. Wir ändern t_{0} ab zu t = t_{0} + δt_{1} und können bei genügend kleiner Wahl von δ > 0 erreichen:

ϵ =< t, β >> 0 , | < t, α > | > ϵ ∀α ∈ R, α ≠ ±β

Wir betrachten die Wurzelbasis S_{t} zur Linearfunktion t. Dann liegt β ∈ R_{t}^{+}. Bzgl. dieser Basis ist β unzerlegbar. Denn für β = α_{1} + α_{2} mit α_{1}, α_{2} > 0 bzgl. S_{t} wäre ϵ = t(β) = t(α_{1}) + t(α_{2}) > 2ϵ. Ein Widerspruch! Also ist β ∈ S_{t}. Nach (2) gibt es ein Element σ in W_{S} mit σ(S_{t}) = S. Damit ist σ(β) ∈ S!

Beweis von (4). Sei β eine Wurzel und s_{β} die zugehörige Spiegelung. Wegen (3) gibt es ein Element σ ∈ W_{S} mit σ(β) = α ∈ S. Dann ist s_{α} = s_{σ(β)} = σs_{β}σ^{−1} ∈ W_{S}, damit s_{β} ∈ σ^{−1}W_{S}σ = W_{S}. Die Spiegelungen {s_{β}|β ∈ R} erzeugen aber die

Weylgruppe W.  ⊔ ⊓

Klassifikation von irreduziblen Wurzelsystemen

Sei (V, R) ein Wurzelsystem.

n(β, α) =< α̌, β >= 2\frac{(α, β)}{(α, α)} ∈ ℤ ∀α, β ∈ R

Betrachte eine Wurzelbasis S. Die Cartanmatrix des Wurzelsystems bzgl. S ist die Matrix

C = (n(β, α))_{β,α∈S}

Wie wir früher gezeigt haben, gilt dann für die Matrixelemente

\begin{aligned}n(α, α) & = 2 \\ n(β, α) & ≤ 0 α ≠ β \\ n(β, α) & ∈ \{0, −1, −2, −3\} α ≠ β \end{aligned}

Aus Theorem 15 (2) folgt, daß die Cartanmatrix ( bis auf die Reihenfolge der Wurzeln ) unabhängig von der Wurzelbasis ist. Es ist nicht schwer, folgenden Satz zu beweisen.

1.21 Satz. Seien (V, R) und (V′, R′) zwei Wurzelsysteme und S, S′ zugehörige Wurzelbasen. Sei f : S → S′ eine bijektive Abbildung mit

n(α, β) = n(f(α), f(β)) ∀α, β ∈ S.

Wir bezeichnen mit f : V → V auch die eindeutige lineare Fortsetzung von f von der Basis S auf ganz V zu einem Isomorphismus V ≅ V′ von Vektorräumen. Dann gilt f(R) = R′. f ist also ein Isomorphismus von Wurzelsystemen. Die Cartanmatrix bestimmt eindeutig das Wurzelsystem!

Die Cartanmatrizen von irreduziblen Wurzelsystemen beschreibt man durch gewisse zusammenhängende Graphen, die sogenannten Dynkindiagramme . Diese Dynkindiagramme können vollständig klassifiziert werden, damit auch alle irreduziblen Wurzelsysteme. Siehe [1, Chap. VI,§4 und Planche I–X]

2. Cartanalgebren

Der Grundkörper k ist in diesem und den folgenden Paragraphen algebraisch abgeschlossen; er hat die Charakteristik 0. Es genügt: k = ℂ. Liealgebren sind, wenn nicht ausdrücklich anders gesagt, endlichdimensional. Sei L eine Liealgebra und H ⊆ L eine Unterliealgebra. Der Normalisator von H in L ist

N(H) = {x ∈ L|[x, H] = ad(x)(H) ⊆ H}

N(H) ist die größte Unterliealgebra von L, die H als Ideal enthält.

2.1 Definition. Die Unterliealgebra H der Liealgebra L heißt Cartanalgebra von L, wenn folgende Bedingungen erfüllt sind

1) H ist nilpotente Liealgebra

2) Es gilt N(H) = H

Konstruktion einer Cartanalgebra

Sei L eine von Null verschiedene Liealgebra. Zu einem Element x ∈ L betrachten wir das charakteristische Polynom

P_{x}(T) = det(T · id − ad(x)) ∈ k[T]

der adjungierten linearen Abbildung ad(x) : L → L. Sei n = dim(V). Dann gilt P_{x}(T) = ∑_{ν=0}^{n} a_{ν}(x)T^{ν} mit von x abhängigen Koeffizienten a_{ν}(x) ∈ k. Beschreibt man x mit Hilfe einer Basis e_{1}, …, e_{n} von L

x = x_{1}e_{1} + … + x_{n}e_{n}

so sind die Koeffizientenfunktionen a_{ν}(x) Polynome in den Koeffizienten x_{1}, …, x_{n}, und es gilt a_{n}(x) = 1.

§2. Cartanalgebren 89

2.2 Definition. Der Rang von L ist die kleinste Zahl e, für die die Koeffizientenfunktion a_{e}(x) nicht identisch verschwindet. Offenbar gilt 0 ≤ e ≤ n. Man sieht leicht:

e = n ⟺ L ist nilpotente Liealgebra

Ein Element a ∈ L heißt regulär, wenn gilt a_{e}(a) ≠ 0.

Zu einem nicht identisch verschwindenden Polynom

P (x_{1}, …, x_{n}) ∈ k[x_{1}, …, x_{n}]

gibt es immer unendlich viele Elementfolgen (a_{1}, …, a_{n}) ∈ k^{n} mit

P (a_{1}, …, a_{n}) ≠ 0.

Denn der Grundkörper k enthält nach Voraussetzung unendlich viele Elemente.

2.3 Satz3. Die Liealgebra L(≠ 0) besitzt unendlich viele reguläre Elemente. Genauer: Die Menge der regulären Elemente von L ist eine Zariski offene dichte Teilmenge von L.

Sei A : V → V eine lineare Abbildung und λ ∈ k. Der Nilraum von A−λ(= A−λ·id) ist die Menge der Vektoren v ∈ V, zu denen es eine natürliche Zahl p gibt mit (A − λ)^{p} = 0. Er ist ein Untervektorraum von V und enthält den Eigenraum zum Eigenwert λ von V. Wenn A halbeinfach ist, ist der Nilraum gleich diesem Eigenraum. Der Nilraum ist genau dann von Null verschieden, wenn λ ein Eigenwert von A ist. Sei x ein Element der Liealgebra L. Wir bezeichnen mit L_{x}^{λ} ⊆ L den Nilraum von ad(x) − λ. Man zeigt

(a) L = ⊕_{λ∈k}L_{x}^{λ}

(b) [L_{x}^{λ}, L_{x}^{μ}] ⊆ L_{x}^{λ+μ} λ, μ ∈ k

(c) L_{x}^{0} ist eine Unterliealgebra von L. x ≠ 0 ⇒ x ∈ L_{x}^{0} ≠ 0

(d) Sei x regulär. Dann ist dim(L_{x}^{0}) = e = Rang(L).

2.4 Theorem, Existenz von Cartanalgebren. Sei x ≠ 0 ein reguläres Element der Liealgebra L. Dann ist L_{x}^{0} eine Cartanalgebra von L.

Beweis. Sei H = L_{x}^{0}. Dann gilt dim(H) = e = rang(L). Für ein Element y ∈ H betrachten wir die induzierten Abbildungen der adjungierten Abbildung ad(y) : L → L.

\begin{matrix}ad_{1}(y) : & H → H \\ ad_{2}(y) : & L/H → L/H \end{matrix}

(a) H ist nilpotente Liealgebra. Nach dem Satz von Engel ist zu zeigen, daß für alle y ∈ L die Abbildung ad_{1}(y) nilpotent ist. Wir betrachten die Zariski offenen Teilmengen von H

U = {y ∈ H|ad_{1}(y) ist nicht nilpotent }
V = {y ∈ H|ad_{2}(y) ist invertierbar }

Weil x in V enthalten ist, ist V ≠ ∅. Wir müssen zeigen, daß U = ∅. Angenommen es sei U ≠ ∅. Dann ist auch die Zariski offene Teilmenge U ∩ V nicht leer, enthält also ein Element y. Da y ∈ U, ist die Dimension des Nilraumes von ad_{1}(y) echt kleiner, als der Rang e = dim(H) von L. Da y in V liegt, ist 0 kein Eigenwert von ad_{2}(y). Es folgt, daß die Dimension von L_{y}^{0} echt kleiner als der Rang e ist. Das ist ein Widerspruch zur Definition des Ranges!

(b) Der Normalisator N von H in L ist H. Sei z ∈ N. Es gilt ad(z)(H) = [z, H] ⊆ H, also [z, x] ∈ H. Nach Definition von H gibt es eine natürliche Zahl p mit ad(x)^{p}(H) = 0.

0 = ad(x)^{p}([z, x]) = −ad(x)^{p}([x, z] = −ad(x)^{p+1}(z)
Also ist z in L_{x}^{0} = H enthalten.  ⊔ ⊓

2.5 Theorem, Eindeutigkeit von Cartanalgebren bis auf Isomorphie. Seien H_{1}, H_{2} zwei Cartanalgebren der Liealgebra L. Dann gibt es einen Automorphismus von Liealgebren σ : L →^{≅} L mit σ(H_{1}) = H_{2}. Genauer gilt: Aut(L) ist eine affine algebraische Gruppe. Die Zusammenhangskomponente G des Einselementes operiert transitiv auf der Menge der Cartanalgebren.

Zum Beweis dieses Eindeutigkeitssatzes siehe [3 , Chap.III.4 ]

2.6 Folgerung. Sei H eine Cartanalgebra der Liealgebra L. Dann gibt es ein reguläres Element x ∈ L mit H = L_{x}^{0}. Insbesondere ist die Dimension einer Cartanalgebra gleich dem Rang von L.

Wir wollen im folgenden immer annehmen, daß die betrachteten Cartanalgebren von der Form L_{x}^{0} x ∈ L regulär sind.

Der halbeinfache Fall

2.7 Theorem. Sei L eine halbeinfache Liealgebra und H ⊆ L eine Cartanalgebra. Dann gilt

1) H ist abelsch.

2) Jedes Element von H ist halbeinfach.

3) {x ∈ L, [x, H] = 0} = H. H ist damit eine maximale abelsche Unterliealgebra von L.

4) Die Einschränkung der Killingform κ(x, y) auf H ist nicht ausgeartet. Bemerkung für ein reguläres Element a ∈ L und λ ≠ 0 ist auch κ|L_{a}^{λ} ⊕ L_{a}^{−λ}

nicht ausgeartet.

Beweis. Wir wählen ein reguläres Elemen a ∈ L mit H = L_{a}^{0}. Beweis von (4) : Seien λ, μ ∈ k mit λ + μ ≠ 0. Man rechnet nach: L_{a}^{λ} ⊥ L_{a}^{μ}. Man betrachtet die direkte Zerlegung L = H ⊕ ∑_{λ≠0} L_{a}^{λ} und erhält daraus eine Orthogonalzerlegung bzgl. der Killingform κ

L = H ⊕ ⊕_{λ≠0 mod±}H^{λ} mit H^{λ} = L_{a}^{λ} ⊕ L_{a}^{−λ}

§3. Darstellungen der Liealgebra sl_{2} 91

Weil die Killingform nicht ausgeartet ist, sind auch die Einschränkungen κ|H und κ|L_{a}^{λ} ⊕ L_{a}^{−λ} für λ ≠ 0 nicht ausgeartet. Das beweist (4). Beweis von (1) Man wendet Cartans Kriterium (IV.4.31) auf die treue Darstellung ad : H → End(L) an und erhält wegen der Nilpotenz von H: κ(x, y) = Spur(ad(x) ∘ ad(y)) = 0 für x ∈ H, y ∈ [H, H]. Also ist [H, H] ⊥ H. Da wegen (4) die Einschränkung κ|H nicht ausgeartet ist, folgt [H, H] = 0. Der Beweis von (3) :

{x ∈ L|[x, H] = 0} ⊆ N(H) = H

Dabei ist N(H) der Normalisator der Cartanalgebra H. Der Beweis von (2) : Sei x ∈ H ein Element der Cartanalgebra H. Wir betrachten die Zerlegung

x = s + n [s, n] = 0

in den halbeinfachen Anteil s und den nilpotenten Anteil n. ad(s), ad(n) sind Polynome in ad(x) über dem Grundkörper k. Weil die adjungierte Darstellung treu ist, gilt damit:

y ∈ L , [y, x] = 0 ⟹ [y, s] = [y, n] = 0

Insbesondere gilt also [s, H] = [n, H] = 0; wegen (3) gilt also s, n ∈ H. Die lineare Abbildung ad(n) : L → L ist nilpotent. H ist nach (1) kommutativ. Also ist für y ∈ H die lineare Abbildung ad(y) mit ad(n) vertauschbar, damit ad(y) ∘ ad(n) ebenfalls nilpotent. Es folgt

κ(y, n) = Spur(ad(y) ∘ ad(n)) = 0 ∀y ∈ H

Weil κ|H nach (4) nicht ausgeartet ist, folgt n = 0. x = s ist also halbeinfach.

⊔ ⊓

3. Darstellungen der Liealgebra sl_{2}

Sei sl_{2} die Unterliealgebra der Matrizenalgebra

sl_{2} = \{A = ( \begin{matrix}a & b \\ c & d \end{matrix} ) ∈ M_{2}(k)| mit Spur(A) = a + d = 0\}

sl_{2} ist halbeinfache, sogar einfache Liealgebra. Diese Liealgebra besitzt über k eine Vektorraumbasis

X = ( \begin{matrix}0 & 1 \\ 0 & 0 \end{matrix} ) , Y = ( \begin{matrix}0 & 0 \\ 1 & 0 \end{matrix} ) , H = ( \begin{matrix}1 & 0 \\ 0 & −1 \end{matrix} )

Es gilt

[X, Y] = H, [H, X] = 2X, [H, Y] = −2Y

Das Element H ist halbeinfach, d.h. bei der adjungierten, damit bei jeder endlichdimensionalen Darstellung von sl_{2} eine halbeinfache lineare Abbildung. kH ist eine Cartanalgebra von sl_{2}. Wir betrachten im folgenden endlichdimensionale Darstellungen

sl_{2} ⟶ End(V),

äquivalent dazu endlichdimensionale sl_{2}–Moduln V. Da das Element H ∈ sl_{2} halbeinfach ist, operiert H halbeinfach auf V. Da der Grundkörper k algebraisch abgeschlossen ist, gilt

V = ⊕_{λ∈k}V^{λ} mit V^{λ} = {v ∈ V |Hv = λv}

V^{λ} ist der Untervektorraum der Vektoren vom Gewicht λ. Da sl_{2} halbeinfach ist, ist V ein halbeinfacher sl_{2}–Modul, zerfällt also in eine direkte Summe von irreduziblen sl_{2}–Untermoduln. Es sollen im folgenden alle irreduziblen sl_{2}–Moduln bestimmt werden.

Sei m = 0, 1, 2, 3 … eine natürliche Zahl und W_{m} der Vektorraum der Dimension m + 1 mit der Basis e_{0}, …, e_{m}. Wir definieren folgende äußere Multiplikationen

He_{n} = (m − 2n)e_{n} Y e_{n} = e_{n+1} Xe_{n} = (m − n + 1)e_{n−1}

In diesen Formeln bedeutet e_{−1} = e_{m+1} = 0.Man rechnet leicht nach, daß sich diese Multiplikation von Basiselementen von sl_{2} und Basiselementen von W_{m} distributiv zu einer Modulstruktur fortsetzen läßt und zeigt leicht

3.1 Lemma. W_{m} ist ein irreduzibler sl_{2}–Modul.

3.2 Theorem. Jeder irreduzible endlichdimensionale sl_{2}–Modul ist isomorph zu einem Modul W_{m}. Insbesondere ist jeder irreduzible sl_{2}–Modul eindeutig durch seine Dimension m + 1, m = 0, 1, 2, … bestimmt.

Beweisidee. Sei λ ∈ k. Ein Element 0 ≠ e ∈ V heißt primitiv vom Gewicht λ, wenn gilt:

He = λe also e ∈ V^{λ}, Xe = 0

Man zeigt dann: Jeder von Null verschiedene endlichdimensionale sl_{2}–Modul enthält ein primitives Element. Sei e ∈ V ein primitives Element vom Gewicht λ. Wir setzen

e_{n} = Y^{n} \frac{e}{n!} n ≥ 0, e_{1} = 0

Man rechnet nach

\begin{aligned}He_{n} & = (λ − 2n)e_{n} d.h. e_{n} ∈ V^{λ−2n} \\ Y e_{n} & = (n + 1)e_{n+1} \\ Xe_{n} & = (λ − n + 1)e_{n−1} \end{aligned}

§4. Das Wurzelsystem einer halbeinfachen Liealgebra 93

Wir haben angenommen, daß V endlichdimensional ist. Die Vektoren e_{ν} haben verschiedenes Gewicht; diejenigen, die nicht Null sind, sind also linear unabhängig. Aus den Formeln folgt: Es gibt eine ganze Zahl m ≥ 0, so daß e_{0}, …, e_{m} von Null verschieden, damit linear unabhängig sind, und so daß e_{m+1} = e_{m+2} = … = 0. Es gilt 0 = Xe_{m+1} = (λ − m)e_{m}, also λ = m. Es folgt, daß ke_{0} ⊕ … ⊕ ke_{m} ein Untermodul von V ist, der isomorph zu W_{m} ist. Wenn V irreduzibel ist, folgt also

V ≅ W_{m}.
⊔ ⊓

4. Das Wurzelsystem einer halbeinfachen Liealgebra

Sei L eine halbeinfache Liealgebra und H ⊆ L eine Cartanalgebra von L. Sei λ ∈ H^{⋆}, also λ : H → k eine Linearfunktion. Setze

L^{λ} = {x ∈ L|[h, x] = λ(h).x ∀h ∈ H, x ∈ L}

Weil H Cartanalgebra ist, also aus halbeinfachen Elementen von L besteht und maximal kommutativ ist, gilt

\begin{aligned}L^{0} & = H \\ L & = ⊕_{λ∈H^{⋆}}L^{λ} \end{aligned}

Die Einschränkung der Killingform κ auf die Cartanalgebra H ist nach §3 Th. 7 (4) nicht ausgeartet.

4.1 Definition und Lemma. Die endliche Menge

R = {α : H ⟶ k|L^{α} ≠ 0, α ≠ 0}

heißt Menge der Wurzeln der Liealgebra L bzgl. der Cartanalgebra H. Es gilt

L = H ⊕ (⊕_{α∈R}L^{α})

Wir zeichnen eine nichtausgeartete symmetrische Bilinearform (x, y) auf L aus, die L–invariant ist, z.B. die Killingform (x, y) = κ(x, y).

4.2 Lemma. Seien α, β ∈ H^{⋆}. Dann gilt [L^{α}, L^{β}] ⊆ L^{α+β}.

4.3 Satz.

1) α, β ∈ H^{⋆} α + β ≠ 0 ⟹ L^{α} ⊥ L^{β} Insbesondere besitzt L die Orthogonalzerle-

gung

L = H ⊕ (⊕_{α∈R mod±}(L^{α} + Ł^{−α})).

Sei α ≠ 0. Durch die Bilinearform (−, −) werden L^{α}, L^{−α} in Dualität gesetzt. Die Einschränkung der Bilinearform (−, −) auf L^{α} ⊕ L^{−α} ist nicht ausgeartet.

Insbesondere gilt dimL^{α} = dimL^{−α}, (−, −)|L^{α} = 0 wenn α ≠ 0. Die Einschränkung auf H ist nicht ausgeartet.

2) Sei x ∈ L^{α}, y ∈ L^{−α}, h ∈ H Dann gilt (h, [x, y]) = α(h)(x, y)

3) Weil (−, −)|H nicht ausgeartet ist, stiftet (−, −) einen Isomorphismus

\begin{matrix}H & ⟶^{≅} H^{⋆} \\ h & ⟼ (x ↦ (h, x)) \end{matrix}

Ordne α ∈ R das bei diesem Isomorphismus zugeordnete Element h_{α} ∈ H zu:

α(h) = (h_{α}, h) ∀h ∈ H. Dann gilt für x ∈ L^{α}, y ∈ L^{−α}

[x, y] = (x, y)h_{α}, insbesondere dim[L^{α}, L^{−α}] = 1

Beweis von (1). Aus der Invarianz der Form (x, y) folgt

0 = ([h, x], y) + (x, [h, y]) = (α(h)x, y) + (x, β(h)y) ∀x ∈ L^{α}, y ∈ L^{β}, h ∈ H

Also ist (α(h) + β(h))(x, y) = 0. Für α + β ≠ 0 gibt es ein Element h ∈ H mit α(h) + β(h) ≠ 0. Es folgt (x, y) = 0 ∀x ∈ L^{α}, y ∈ L^{β}. Insbesondere gilt

α ≠ 0 ⟹ L^{α} ⊥ L^{α}, L^{α} ⊥ H

Ferner gilt

L = H ⊕ (⊕_{α≠0 mod±}(L^{α} ⊕ L^{−α})).

L ist also die orthogonale direkte Summe von H und den Unterräumen L^{α} ⊕ L^{−α} α ∈ R , α mod ± . Es folgt, daß die Einschränkungen der nicht ausgearteten Form (x, y) auf H und auf die Räume L^{α} ⊕ L^{−α} α ∈ R nicht ausgeartet sind. Wegen (−, −)|L^{α} = 0 für α ∈ R folgt, daß L^{α} und L^{−α} in Dualität stehen. Beweis von (2): Wegen der Invarianz der Form (x, y) gilt für h ∈ H, x ∈ L^{α}, y ∈ L^{−α}, α ∈ R

(h, [x, y]) = ([h, x], y) = α(h)(x, y)

Beweis von (3) Wegen (2) gilt nach Definition von h_{α} für alle h ∈ H

(h, [x, y] − (x, y)h_{α}) = 0

Weil (−, −)|H nicht ausgeartet ist, folgt [x, y] = (x, y)h_{α}.

4.4 Lemma. R erzeugt den Dualraum H^{⋆} von H.

Beweis. Wir beweisen die duale Aussage. Sei h ∈ H ein Element mit α(h) = 0 ∀α ∈ R. Es folgt ad(h)|L^{α} = 0 ∀α ∈ R. Weil die Cartanalgebra H kommutativ ist gilt auch ad(h)|H = 0. Also ist h im Kern von ad, also im Zentrum der halbeinfachen Liealgebra L enthalten; es gilt h = 0. ⊔ ⊓

Wegen (3) gilt [L^{α}, L^{−α}] = kh_{α}. Wir beweisen:

4.5 Lemma. Sei α ∈ R eine Wurzel von L. Dann gilt α([L^{α}, L^{−α}] = kh_{α}) ≠ 0. Es gibt also genau ein Element H_{α} ∈ [L^{α}, L^{−α}] = kh_{α} mit α(H_{α}) = 2. Es ist dann [L^{α}, L^{−α}] = kH_{α}.

Beweis. Angenommen, es sei α([L^{α}, L^{−α}]) = 0. Wähle x ∈ L^{α}, y ∈ L^{−α} mit 0 ≠ z = [x, y] ∈ [L^{α}, L^{−α}]. Es folgt [z, x] = α(z)x = 0, [z, y] = −α(z)y = 0. G = kz +kx+ ky ⊆ L ist eine Unterliealgebra von L. G ist auflösbar. Also ist [G, G] nilpotent bei jeder Darstellung von G, also auch bei der Darstellung ad : G → End(L). Wegen [G, G] = kz ist ad(z) : L → L nilpotente lineare Abbildung; also ist das Element z ∈ H ein nilpotentes Element von L. Die Cartanalgebra H enthält aber

nur halbeinfache Elemente von L. Widerspruch!  ⊔ ⊓

Sei nun α eine Wurzel und H_{α} ∈ [L^{α}, L^{−α}] das Element mit α(H_{α}) = 2. Dann ist [L^{α}, L^{−α}] = kH_{α}. Es gibt dann Elemente X_{α} ∈ L^{α}, Y_{α} ∈ L^{−α} mit [X_{α}, Y_{α}] = H_{α}. Betrachte die Unterliealgebra

G_{α} = kH_{α} + kX_{α} + kY_{α}

Dann ist

\begin{matrix}G_{α} & ≅ sl_{2} \\ X_{α} & ↦ X = ( \begin{matrix}0 & 1 \\ 0 & 0 \end{matrix} ) \\ Y_{α} & ↦ Y = ( \begin{matrix}0 & 0 \\ 1 & 0 \end{matrix} ) \\ H_{α} & ↦ H = ( \begin{matrix}1 & 0 \\ 0 & −1 \end{matrix} ) \end{matrix}

Wir greifen zurück auf die Theorie der Darstellungen der Liealgebra sl_{2} ( siehe §3 ). Man schränkt die adjungierte Darstellung ad : L → End(L) auf G_{α} = sl_{2} ein und betrachtet L als G = sl_{s}–Modul.

4.6 Lemma. Für alle Wurzeln α ∈ R von L gilt dimL^{α} = 1. Insbesondere gilt

G_{α} = [L^{α}, L^{−α}] + L^{α} + L^{−α}

Y_{α} ∈ L^{−α} ist eindeutig durch X_{α} ∈ L^{α} bestimmt.

Beweis. Angenommen, es sei dimL^{α} > 1. L^{α} und L^{−α} stehen nach Satz 4.3 (1) in Dualität zueinander bei (−, −). Wähle ein von 0 verschiedenes Element y in L^{−α} mit y ⊥ X_{α}. Weil (−, −)|H nicht ausgeartet ist folgt aus Satz 4.3 (2) [X_{α}, y,] = 0. Außerdem ist [H_{α}, y,] = −α(H_{α})y = −2y. y ist also ein primitives Element mit negativem Gewicht in einem endlichdimensionalen sl_{2}–Modul. Das steht im

Widerspruch zur Darstellungstheorie der Liealgebra sl_{2} aus §3.  ⊔ ⊓

4.7 Satz.

1)

α, β ∈ R ⟹ β(H_{α}) ∈ ℤ, β − β(H_{α})α ∈ R

R ⊆ H^{⋆} ist also ein Wurzelsystem; H_{α} ∈ H = (H^{⋆})^{⋆} ist die Kowurzel zu α. Die Spiegelung zu α ist

s_{α}(x) = x − x(H_{α})α = x− < x, H_{α} > α

2) Das Wurzelsystem (H^{⋆}, R) ist reduziert, d.h.

α ∈ R, λα ∈ R ⟹ λ = ±1

Beweis. Beweis von (1): Nach Lemma 4.4 erzeugt R den Dualraum H^{⋆}. Seien α, β ∈ R. Wie vorher:

\begin{aligned}G_{α} & = kH_{α} + kX_{α} + kY_{α} = sl_{2} = G \\ 0 ≠ y ∈ L^{β} & ⟹ [H_{α}, y] = ad(H_{α})y = β(H_{α})y \end{aligned}

β(H_{α}) tritt also als Gewicht im sl_{2}–Modul L auf. Es folgt, daß p = β(H_{α}) eine ganze Zahl ist. Setze im Sinne von Darstellungsmoduln

z = \{ \begin{matrix}Y_{α}^{p}y & für p ≥ 0 \\ X_{α}^{−p}y & für p ≤ 0 \end{matrix}

Da jede endlichdimensionale Darstellung von sl_{2} in eine direkte Summe irreduzibler zerfällt, folgt aus der Struktur der irreduziblen sl_{2}–Moduln, daß z nicht verschwindet. (siehe §3 Th. 2) Andererseits gilt nach Lemma 4.2 z ∈ L^{β−pα}. Also ist L^{β−pα} ≠ 0 und deshalb β − β(H_{α}) auch eine Wurzel. Beweis von (2): Angenommen, R ist nicht reduziert. Dann gibt es eine Wurzel α ∈ R mit 2α ∈ R. Wähle 0 ≠ y ∈ L^{2α}. Dann gilt [H_{α}, y] = 2α(H_{α})y = 4y. 3α ist keine Wurzel, [X_{α}, y] ∈ L^{3α} = 0, also [X_{α}, y] = 0. y ist damit im sl_{2}–Modul L ein primitives Element und hat das Gewicht 4. Aus der Darstelungstheorie von sl_{2} folgt, daß der von y erzeugte sl_{2}(= G_{α})–Modul irreduzibel ist und die Dimension 5 hat. Nun gilt H_{α} = [X_{α}, Y_{α}] und

\begin{aligned}ad(H_{α})(y) & = ad(X_{α})ad(Y_{α})(y) − ad(Y_{α})ad(X_{α})(y) = ad(X_{α})ad(Y_{α})(y) \\ & ∈ ad(X_{α})(L^{α}) = k[X_{α}, X_{α}] = 0 \end{aligned}

Also ist 0 = [H_{α}, y] = 4y. Widerspruch!  ⊔ ⊓

4.8 Satz. Sei wieder L ≠ 0 eine halbeinfache Liealgebra, H ⊆ L eine Cartanalgebra und R die Menge der Wurzeln von L bzgl. H. Wir wissen bereits, daß (H^{⋆}, R) ein reduziertes Wurzelsystem ist. Wie vorher ist H_{α} ∈ H die Kowurzel zur Wurzel α und L^{α} = {x ∈ L|ad(h)x = α(h)x ∀h ∈ H}. Seien α, β ∈ R zwei Wurzeln und sei α ≠ ±β. Es gilt

1) Die Menge der Wurzeln aus R der Form β + nα mit einer ganzen Zahl n heißt die α–Reihe durch β. Sie enthält offensichtlich α. Dann ist {n ∈ ℤ|β + nα ∈ R} ein Intervall. Weil 0 in dieser Menge enthalten ist gibt es also ganze Zahlen p, q ≥ 0 so daß gilt

{n ∈ ℤ|β + nα ∈ R} = [−p, q] = {n ∈ ℤ| − p ≤ n ≤ q}

Es gilt außerdem:

p − q = β(H_{α})

2)

[L^{α}, L^{β}] = L^{α+β}(≠ 0 ⟺ α + β ∈ R).

3) Seien U_{α} ∈ L^{α}, V_{α} ∈ L^{−α} und Z ∈ L^{β}. Es sei [U_{α}, V_{α}] = H_{α}. Dann ist

[V_{α}, [U_{α}, Z]] = q(p + 1)Z(≠ 0 ⟺ Z ≠ 0 , α + β ∈ R)

Dabei sind p, q ≥ 0 die ganzen Zahlen aus (1), die die α–Reihe durch β beschreiben. Bemerkung: Es genügt, die Relation für die vorher fest gewählten Elemente U_{α} = X_{α}, V_{α} = Y_{α} zu zeigen.

Beweis. Wir betrachten wieder L mit Hilfe der adjungierten Darstellung als Modul über der Unterliealgebra G_{α} = kH_{α} + kX_{α} + kY_{α} = sl_{2}. Wir benutzen wieder die Klassifikation der Darstellungen von sl_{2} aus §3. Wir betrachten den Untermodul M = ∑_{n∈ℤ} L^{β+nα}. Es gilt: u ∈ L^{β+nα} ⇒ ad(H_{α})(u) = β(H_{α}) + 2n. Die auftretenden Gewichte, d.h. Eigenwerte von H_{α} in M, sind also die ganzen Zahlen

{β(H_{α}) + 2n|β + nα gehört zur α–Reihe durch β}.

Wir zerlegen M in eine direkte Summe von irreduziblen Moduln. Sei N ein solcher irreduzibler direkter Summand. Er ist nach §3 isomorph zu einem sl_{2}–Modul W_{m}; 0 ≤ m ∈ ℤ. Aus der Struktur von W_{m} folgt daß es ganze Zahlen r ≤ s gibt mit W_{m} ≅ N = ∑_{r≤n≤s} L^{β+nα} und β + nα ∈ R ∀r ≤ n ≤ s. Man folgert mit Hilfe der bekannten Gewichte von W_{m}, daß gilt

β(H_{α}) + 2s = m = dimW_{m} − 1 = s − r ⟹ β(H_{α}) = −(r + s).

r + s ist insbesondere unabhängig von der Wahl des direkten Summanden N von M. Wäre N ≠ M, so gäbe es einen zweiten irreduzible direkten Summanden N′ = ∑_{r′≤n≤s′} L^{β+nα} mit r ≤ s ≤ r′ ≤ s′. Wegen r + s = r′ + s′ erhält man einen Widerspruch! Es gilt also

\begin{matrix}M = N = ∑_{r≤n≤s} L^{β+nα} \\ und β + nα ∈ R ∀r ≤ n ≤ s \end{matrix}

Damit ist Punkt (1) bewiesen. Die Punkte (2) und (3) ergeben sich ebenfalls leicht

wegen der Isomorphie M ≅ W_{m} aus der Struktur von W_{m}.  ⊔ ⊓

5. Der Isomorphiesatz

Es soll gezeigt werden, daß zwei halbeinfache Liealgebren über dem algebraisch abgeschlossenen Grundkörper k der Charakteristik Null genau dann isomorph sind, wenn ihre Wurzelsysteme isomorph sind. Dazu beschreibt man eine solche halbeinfache Liealgebra L ≠ 0 durch Erzeugende und Relationen zwischen diesen Erzeugenden.

Sei H ⊆ L eine Cartanalgebra der Liealgebra L und sei R ⊆ H^{⋆} die Menge der Wurzeln von L bzgl. der Cartanalgebra H. Sei S = (α_{1}, …, α_{n}) ⊆ R eine Wurzelbasis und (H_{1}, …, H_{n}) ⊆ H die zugehörigen Kowurzeln. Es gilt insbesondere α_{i}(H_{i}) = 2. Dann ist (H_{1}, …, H_{n}) eine Wurzelbasis des dualen Wurzelsystems der Kowurzeln. Insbesondere ist H_{1}, …, H_{n} eine Basis von H. Sei für eine Wurzel α

L^{α} = {x ∈ L|[h, x] = α(h)x ∀h ∈ H}

Dann ist dim L^{α} = 1, und es gilt L = H ⊕(⊕_{α∈R}L^{α}). Wir wählen zu jeder einfachen Wurzel α_{i} ∈ S ein Paar X_{i} ∈ L_{i}^{α} = kX_{i} , Y_{i} ∈ L^{−α_{i}} = kY_{i} mit [X_{i}, Y_{i}] = H_{i}. Wie wir gezeigt haben (§4 Satz 4.8 (2)) gilt für zwei Wurzeln α, β ∈ R

[L^{α}, L^{β}] = \{ \begin{matrix}L^{α+β} & α + β ∈ R \\ 0 & α + β ≠ 0 , α + β ∉ R \end{matrix}

Wir benutzen folgende Bezeichnungen. Sei ν = (ν_{1}, …, ν_{m}) ein m–Tupel ganzer Zahlen mit 1 ≤ ν_{i} ≤ n. Dann setzen wir

\begin{aligned}X_{ν} & = [X_{ν_{m}}, [X_{ν_{m−1}}, […, [X_{ν_{2}}, X_{ν_{1}}] …] \\ Y_{ν} & = [Y_{ν_{m}}, [Y_{ν_{m−1}}, […, [Y_{ν_{2}}, Y_{ν_{1}}] …] \end{aligned}

Sei

α(ν) = α_{ν_{1}} + … + α_{ν_{m}}

Dann gilt X_{ν} ∈ L^{α(ν)} , Y_{ν} ∈ L^{−α(ν)}.

X_{ν} ≠ 0 bzw. Y_{ν} ≠ 0 ⟺ α_{1} + … + α_{s} ∈ R ∀1 ≤ s ≤ m

Aus früheren Resultaten aus §4, insbesondere aus Satz 4.8, ergeben sich eine Reihe von Relationen, die allein durch das Wurzelsystem bestimmt sind.

5.1 Relationen.

1) H ist kommutativ: [h, h′] = 0 ∀h, h′ ∈ H

§5. Der Isomorphiesatz 99

2) Sei ein m–Tupel ν = (ν_{1}, …, ν_{m}) gegeben mit α = α(ν) ∈ R^{+}. Dann gilt für h ∈ H : [h, X_{ν}] = α(h)X_{ν} , [h, Y_{ν}] = −α(h)Y_{ν} 3) [X_{i}, Y_{j}] = δ_{ij}H_{i} 1 ≤ i, j ≤ n Dabei ist

δ_{ij} = \{ \begin{matrix}1 & i = j \\ 0 & i ≠ j \end{matrix}

das Kroneckersymbol

4) Gegeben sei ein m–Tupel ν = (ν_{1}, …, ν_{m}) mit α(ν) = α_{ν_{1}} + … + α_{ν_{m}} ∉ R. Dann gilt X_{ν} = Y_{ν} = 0.

5) Gegeben sei ein m–Tupel ν = (ν_{1}, …, ν_{m}) und eine einfache Wurzel α_{i} ∈ S. Sei β = α(ν). Falls β eine Wurzel ist betrachte die Wurzeln der α_{i}– Reihe durch β : ( siehe §4 S. 8 )

{β + sα_{i}| − p ≤ s ≤ q}

Dabei sind p, q ganze Zahlen mit p, q ≥ 0. Es gilt

\begin{aligned}[Y_{i}, [X_{i}, X_{ν}]] & = q(p + 1)X_{ν}(≠ 0 ⟺ β + α_{i} ∈ R , X_{ν} ≠ 0) \\ [Y_{i}, [X_{i}, Y_{ν}]] & = p(q + 1)Y_{ν}(≠ 0 ⟺ β − α_{i} ∈ R , Y_{ν} ≠ 0) \\ [X_{i}, [Y_{i}, X_{ν}]] & = −p(q + 1)X_{ν}(≠ 0 ⟺ β − α_{i} ∈ R , X_{ν} ≠ 0) \\ [X_{i}, [Y_{i}, Y_{ν}]] & = −q(p + 1)Y_{ν}(≠ 0 ⟺ β + α_{i} ∈ R , Y_{ν} ≠ 0) \end{aligned}

Wir betrachten nun Variable U_{1}, …, U_{n} , V_{1}, …, V_{n}. Wir betrachten den Vektorraum

𝔙 = H ⊕ kU_{1} ⊕ … ⊕ kU_{n} ⊕ kV_{1} ⊕ … ⊕ kV_{n}

Sei 𝔉 = 𝔉(S) die “freie” Liealgebra über dem Grundkörper k, erzeugt vom Vektorraum 𝔙. Es gilt dann die universelle Abbildungseigenschaft: Zu jeder linearen Abbildung 𝔙 → G in eine beliebige auch unendlichdimensionale Liealgebra G gibt es einen eindeutig bestimmten Homomorphismus von Liealgebren 𝔉 → G, der die lineare Abbildung 𝔙 → G fortsetzt. Dann erzeugt 𝔙 ⊆ 𝔉 die Liealgebra 𝔉. D.h., daß 𝔉 die kleinste Lieunteralgebra von 𝔉 ist, die 𝔙 enthält. 𝔉 ist im allgemeinen unendlichdimensional. Analog zu X_{ν}, Y_{ν} ∈ L definieren wir in 𝔉 zu jedem m–Tupel ν = (ν_{1}, …, ν_{m}) , 1 ≤ m ≤ n Elemente

\begin{aligned}U_{ν} & = [U_{ν_{m}}, [U_{ν_{m−1}}, […, [U_{ν_{2}}, U_{ν_{1}}] …] \\ V_{ν} & = [V_{ν_{m}}, [V_{ν_{m−1}}, […, [V_{ν_{2}}, V_{ν_{1}}] …] \end{aligned}

in 𝔉. Es gibt dann einen natürlichen Homomorphismus von Liealgebren

𝔉 ⟶ L

der die Identität auf H ⊆ 𝔉 , L ist und der U_{i} in X_{i} und V_{i} in Y_{i}, 1 ≤ i ≤ n abbildet. Für ein m–Tupel ν gilt dann:

U_{ν} ⟼ X_{ν} , V_{ν} ⟼ Y_{ν}

Da H und die Elemente X_{ν}, Y_{ν} den Vektorraum L erzeugen, ist dieser Homomorphismus surjektiv. Wir untersuchen das Kernideal. Die aufgezählten Relationen (1)–(5) definieren auf natürliche Weise Elemente von 𝔉, die in diesem Kernideal liegen. Sei 𝔞 das von diesen Elementen erzeugte Ideal der Liealgebra 𝔉 = 𝔉(S), das dann in dem Kernideal von 𝔉 ⟶ L liegt. Es ist allein durch das Wurzelsystem (H^{⋆}, R) und durch die Wurzelbasis S bestimmt.

5.2 Theorem. Der natürliche surjektive Homomorphismus 𝔉 ⟶ L induziert einen Isomorphismus

𝔉(S)/𝔞 ⟶^{≅} L

Aus diesem Theorem erhalten wir als wichtige Folgerung das Hauptresultat dieses Paragraphen. Sei L′ eine zweite halbeinfache Liealgebra und H′ ⊆ L′ eine Cartanalgebra von L′. Sei (H′^{⋆}, R′) das Wurzelsystem von L^{⋆} bzgl. H^{⋆}. Gegeben sei eine isomorphe lineare Abbildung

ϕ : H^{⋆} ⟶^{≅} H′^{⋆} mit ϕ(R) = R′,

also ein Isomorphismus von Wurzelsystemen. Wir betrachten den dualen Isomorphismus ψ = ϕ^{⋆}^{−1} : H ⟶^{≅} H′. S′ = ϕ(S) ist eine Wurzelbasis von R′. Die Kowurzeln H_{i} ∈ H zu α_{i} ∈ S werden durch ψ auf die Kowurzeln ψ(H_{i}) zu ϕ(α_{i}) ∈ S′ abgebildet. Aus Theorem 2 folgt

5.3 Theorem, Isomorphiesatz. Der Isomorphismus

ψ : H ⟶^{≅} H′.

läßt sich zu einem Isomorphismus von Liealgebren

L ⟶^{≅} L′.

fortsetzen. Insbesondere gilt: Halbeinfache Liealgebren mit isomorphen Wurzelsystemen sind isomorph.

5.4 Bemerkung. Zu jedem reduzierten Wurzelsystem kann eine halbeinfache Liealgebra konstruiert werden, deren Wurzelsystem bzgl. einer Cartanalgebra zum vorgegebenen System isomorph ist. Direkte Summen von Wurzelsystemen entsprechen direkte Produkte der zugehörigen Liealgebren. Da man alle irreduziblen Wurzelsysteme mit Hilfe der Dynkindiagramme klassifizieren kann, sind damit alle halbeinfachen Liealgebren über einem algebraisch abgeschlossenen Grunkörper der Charakteristik Null klassifiziert. Bis auf endlich viele exotische Liealgebren erhält man alle einfachen Liealgebren als Liealgebren der klassischen Gruppen, also der speziellen linearen Gruppen, der speziellen orthogonalen Gruppen und der symplektischen Gruppen.

5.5 Bemerkung. Seien H_{1}, H_{2} zwei Cartanalgebren der halbeinfachen Liealgebra L. Dann gibt es einen Autorphismus von Liealgebren σ : L →^{≅} L mit σ(H_{1}) = H_{2}. Man folgert, daß die Wurzelysteme zu verschiedenen Cartanalgebren von L isomorph sind. Es kann die ”äußere“ Automorphismengruppe einer einfachen Gruppe als Automorphismengruppe des zugehörigen Dynkindiagramms bestimmt werden.

Zu diesen Bemekungen siehe Bourbaki Groupes et algebres de Lie Chap, 4,5,6 und Chap. 7 et 8. Beweis von Theorem 5.5. Zu jeder positiven Wurzel α ∈ R^{+} wählen wir ein Tupel ν(α) = (ν_{1}, …, ν_{m}) mit α_{ν_{1}} + … + α_{ν_{m}} = α und

α_{ν_{1}} + … + α_{ν_{s}} ∈ R ∀1 ≤ s ≤ m

(Satz 1.17) Dann sind X_{ν(α)}, Y_{ν(α)} ≠ 0, und es gilt L^{α} = kX_{ν(α)} , L^{−α} = kY_{ν(α)}. Sei

\begin{aligned}U_{α} & = U_{ν(α)} , V_{α} = V_{ν(α)} \\ 𝔉_{0} & = H ⊕ ⊕_{α∈R^{+}}(kU_{α} ⊕ kV_{α}) ⊆ 𝔉 \end{aligned}

Diese Summe ist eine direkte Summe, weil das Bild in L eine direkte Summe ist. 𝔉_{0} wird also isomorph auf L abgebildet. Zum Beweis des Theorems genügt es zu zeigen:

𝔉 = 𝔉_{0} + 𝔞,

d.h., daß mod 𝔞 jedes Element von 𝔉 in 𝔉_{0} enthalten ist. Wir betrachten in 𝔉 die Kongruenz x ≡ y modulo 𝔞. Teil I des Beweises. Sei ν = (ν_{1}, …, ν_{m}) wieder eines der betrachteten Tupel. Sei α = α(ν) eine–damit positive– Wurzel. Dann sind U_{α}, V_{α} definiert. Behauptung

\begin{aligned}U_{ν} & ≡ cU_{α} \\ V_{ν} & ≡ c′V_{α} \end{aligned}

Dabei hängen die Konstanten c, c′ von ν ab. Wir beschränken uns auf den Beweis der ersten Kongruenz für die U. Der Beweis der zweiten verläuft analog. Wegen Relation (4) können wir annehmen, daß

α_{ν_{1}} + … + α_{ν_{s}} ∈ R ∀1 ≤ s ≤ m

Wir führen den Beweis durch Induktion nach der Länge m des Tupels. Sei m ≥ 2. Weil S eine Wurzelbasis ist, hat das ausgewählte T upel zur Wurzel α die gleiche Länge wie das vorgegebene Tupel ν und unterscheidet sich von ν nur durch eine Permutation der Komponenten. Sei

ν(α) = (i_{1}, …, i_{m}) 1 ≤ i_{j} ≤ n

Wir unterscheiden zwei Fälle. Fall (a) Sei ν_{m} = i_{m} = r. Sei

α_{ν_{1}} + … + α_{ν_{m−1}} = α_{i_{1}} + … + α_{i_{m−1}} = γ ∈ R^{+}

Nach Induktionsvoraussetzung gilt

\begin{aligned}U_{ν} & ≡ c[U_{r}, U_{γ}] \\ U_{α} & ≡ d[U_{r}, U_{γ}] \end{aligned}

Es ist d ≠ 0. Sonst wäre U_{α} ≡ 0; das Bild X_{ν(α)} von U_{α} in L würde verschwinden. Aber X_{ν(α)} ≠ 0 nach Voraussetzung über das Tupel ν(α). Es folgt U_{ν} ≡ \frac{c}{d}U_{α}. Der Fall (b) Sei r = i_{m} ≠ ν_{m}. Dann gibt es eine ganze Zahl 1 ≤ s < m mit ν_{s} = i_{m} = r und ν_{s′} ≠ r ∀s′ > s. Sei

\begin{aligned}ν′ & = (ν_{1}, …, ν_{s−1}, ν_{s+1}, …, ν_{m}) \\ ν̃ & = (ν_{1}, …, ν_{s−1}, ν_{s+1}, …, i_{m}) \end{aligned}

Mit Hilfe von Relation (5) erhält man

[U_{r}, [V_{r}, U_{ν}]] ≡ dU_{ν} mit d ≠ 0

wegen α − α_{r} = α_{i_{1}} + … + α_{i_{m−1}} ∈ R^{+}. Durch mehrfache Anwendung der Produktregel( bzw. der Jakobiidentiät ) und durch Verwendung der Relation (3)erhält man

[V_{r}, U_{ν}] ≡ [U_{ν_{m}}, [U_{ν_{m−1}}, …, [U_{ν_{s+1}}, [V_{r}, [U_{ν_{s}}, [U_{ν_{s−1}}, […] …]

Wir verwenden noch einmal die Relation (5) und erhalten wegen r = ν_{s}

\begin{aligned}[V_{r}, U_{r}] & ≡ cU_{ν′} \\ [U_{r}, [V_{r}, U_{ν}]] & ≡ c[U_{r}, U_{ν′}] = cUν̃ \end{aligned}

Es folgt

U_{ν} ≡ \frac{c}{d}Uν̃

Damit ist der Fall (b) auf den Fall (a) zurückgeführt. Teil II des Beweises. (a): Sei ν = (ν_{1}, …, ν_{m}) mit 1 ≤ ν_{j} ≤ n wieder eines der betrachteten Tupel. Aus Teil I des Beweises erhält man mit Hilfe der Relationen (1)—(5) und durch Anwendung der Jakobiidentität ( oder der Produktregel )

H, U_{ν}, V_{ν} ⊆ 𝔉_{0} mod 𝔞
[H, H] ≡ 0 ∈ 𝔉_{0} mod 𝔞
[h, U_{ν}], [h, V_{ν}] ⊆ 𝔉_{0} mod 𝔞 ∀h ∈ H
[U_{i}, U_{ν}], [V_{i}, V_{ν}] ⊆ 𝔉_{0} mod 𝔞 i = 1, …, n
[U_{i}, V_{ν}], [V_{i}, U_{ν}] ⊆ 𝔉_{0} mod 𝔞 i = 1, …, n

Man beachte, daß auch die Elemente U_{α}, V_{α} von der Form U_{ν}, V_{ν} sind. (b): Man zeigt nun allein mit Hilfe der Jakobiidentität, ohne Verwendung von Kongruenzen, daß folgende Elemente A ein Vektorraumerzeugendensystem von 𝔉 bilden:

A = [B_{1}, [B_{2}, […] …] mit B_{j} von der Form h ∈ H, U_{i}, V_{i} i = 1, …, n

Es bleibt dann für solche Elemente A zu zeigen, daß A ∈ 𝔉_{0} mod 𝔞. Mit Hilfe von Induktion nach der Länge m von A folgert man diese Enthaltenseinsrelation mod 𝔞 aus den Enthaltenseinsrelationen mod 𝔞 für die in (a) behandelten Lieprodukte.

⊔ ⊓

§6. Anhang 103

6. Anhang

Irreduziblen Darstellungen einer halbeinfachen Liealgebra

Ein Resümee

Sei L eine halbeinfache Liealgebra über dem algebraisch abgeschlossenenGrundkörper k der Charakteristik Null. Es genügt anzunehmen, daß k = ℂ ist. Sei H ⊆ L eine vorgegebene Cartanalgebra und R ⊆ H^{⋆} die Menge der Wurzeln von L bzgl. H. Wir betrachten das Wurzelsystem (H^{⋆}, R). Zu jeder Wurzel α ∈ R sei H_{α} ∈ H die zugehörige Kowurzel.(siehe auch III.4 L.5, S.7 ) Sei S = (α_{1}, …, α_{n}) eine vorgegebene Wurzelbasis und R^{+} die Menge der positiven Wurzeln. Wir betrachten das Gitter Γ der “Gewichte” ω ∈ H^{⋆} mit < H_{α}, ω >∈ ℤ ∀α ∈ R. R ist in Γ enthalten. Eine Basis von Γ über ℤ sind die Elemente

ω_{i}| < H_{α_{j}}, ω_{i} >= \{ \begin{matrix}1 & i = j \\ 0 & sonst \end{matrix} i = 1, …, n

In Γ betrachten wir die Teilmenge der positiven “Gewichte”

Γ^{+} = {ω ∈ Γ| < H_{α}, ω >≥ 0 ∀α ∈ R^{+}}

Es gilt

Γ^{+} = {m_{1}ω + … + m_{n}|m_{ν} ∈ ℤ , m_{ν} ≥ 0}

Wir betrachten im folgenden endlichdimensionale Moduln bzw. Darstellungen von L. Jeder solche Modul ist direkte Summe von irreduziblen Moduln.( Kap. III §4 ) Wir wollen in Verallgemeinerung der Resultate aus §3 über die Liealgebra sl_{2} alle irreduziblen Darstellungen einer beliebigen halbeinfachen Liealgebra L klassifizieren.

6.1 Definition. Sei V ein L–Modul. Ein Element 0 ≠ v ∈ V heißt primitiv vom Gewicht ω ∈ H^{⋆}, wenn gilt

\begin{aligned}hv & = ω(h)v ∀h ∈ H \\ xv & = 0 ∀x ∈ L^{α} , ∀α ∈ R^{+} \end{aligned}

6.2 Theorem.

1) Sei V ≠ 0 ein endlichdimensionaler irreduzibler L–Modul. Dann enthält V ein bis auf einen Faktor eindeutig bestimmtes primitives Element. Das durch V eindeutig bestimmte Gewicht dieses Elementes liegt in Γ^{+}. Wir nennen es das Hauptgewicht von V.

2) Zwei irreduzible endlichdimensionale nicht verschwindende L–Moduln sind genau dann isomorph, wenn ihre Hauptgewichte gleich sind.

3) Zu jedem Element ω ∈ Γ^{+} gibt es einen irreduziblen nicht verschwindenden endlichdimensionalen L–Modul mit Hauptgewicht ω.

104 Chapter V. Klassifikation halbeinfacher Liealgebren

Siehe auch

1) Serre Algèbres de Lie semi-simples complexes Chap. VII

1) Bourbaki Groupes et algèbres de Lie Chap. VIII, §6,7